### v1

# Estimation of Out-of-Sample Sharpe Ratio for High Dimensional Portfolio Optimization

Abstract
Portfolio optimization aims at constructing a realistic portfolio with significant out-of-sample performance, which is typically measured by the out-of-sample Sharpe ratio. However, due to in-sample optimism, it is inappropriate to use the in-sample estimated covariance to evaluate the out-of-sample Sharpe, especially in the high dimensional settings. In this paper, we propose a novel method to estimate the out-of-sample Sharpe ratio using only in-sample data, based on random matrix theory. Furthermore, portfolio managers can use the estimated out-of-sample Sharpe as a criterion to decide the best tuning for constructing their portfolios. Specifically, we consider the classical framework of Markowits mean-variance portfolio optimization with known mean vector and the high dimensional regime of $p/n\to c\in(0,\infty)$, where $p$ is the portfolio dimension and $n$ is the number of samples or time points. We propose to correct the sample covariance by a regularization matrix and provide a consistent estimator of its Sharpe ratio. The new estimator works well under either of three conditions: (1) bounded covariance spectrum, (2) arbitrary number of diverging spikes when $c<1$, and (3) fixed number of diverging spikes when $c\geq 1$. We can also extend the results to construct global minimum variance portfolio and correct out-of-sample efficient frontier. We demonstrate the effectiveness of our approach through comprehensive simulations and real data experiments. Our results highlight the potential of this methodology as a powerful tool for portfolio optimization in high dimensional settings.
Abstract
This supplementary material contains additional numerical results and all the proofs for theorems in the main text.
Keywords: Portfolio allocation; Efficient frontier recovery; High dimensionality; Spiked covariance structure; Ridge regularization.
1 Introduction
The mean-variance portfolio optimization, first introduced by Markowitz (1952), is a fundamental cornerstone in modern portfolio allocation theory. Suppose that the excessive returns of $p$ risky assets have population excess mean vector ${\bm{\mu}}\in\mathbb{R}^{p}$ and population covariance matrix ${\bm{\Sigma}}\in\mathbb{R}^{p\times p}$. In mean-variance portfolio optimization, one solves for the allocation vector $\mathbf{w}$ that maximizes $\mathbf{w}^{\top}{\bm{\mu}}-\lambda\mathbf{w}^{\top}{\bm{\Sigma}}\mathbf{w}$, where $\lambda$ is the risk aversion. The solution $\mathbf{w}^{*}$ of mean-variance portfolio optimization is then proportional to ${\bm{\Sigma}}^{-1}{\bm{\mu}}$, that is, $\mathbf{w}^{*}\propto{\bm{\Sigma}}^{-1}{\bm{\mu}}$. It is well known that this solution achieves the largest possible Sharpe ratio.
Despite the clean framework and theory of mean-variance portfolio, there are challenges in its practical applications. In practice, the population mean ${\bm{\mu}}$ and population covariance matrix ${\bm{\Sigma}}$ are unknown and have to be estimated using historical data. This leads to the phenomenon known as the “empirical optimism”, which states if we use the in-sample data both for estimation and for out-of-sample assessment, we are overly optimistic about the true out-of-sample performance of the estimated portfolio (Karoui, 2010). Thus, in order to genuinely evaluate the out-of-sample Sharpe ratio using historical data, we have to develop a valid and accurate estimator, which is the main purpose of the current paper. The estimator will help portfolio managers to achieve a portfolio with superior expected future performance.
Many works in the literature have studied the estimation of ${\bm{\mu}}$ (Green et al., 2013; Fischer & Krauss, 2018; Baba Yara, 2020; Petropoulos et al., 2022; Wasserbacher & Spindler, 2022), which is often quite challenging as the signal level of financial data is weak. In comparison, the estimation of the population covariance matrix ${\bm{\Sigma}}$ is also particularly difficult not because of signal level, but due to the curse of high dimensionality. For instance, even when ${\bm{\Sigma}}$ is an identity matrix, it is widely known that the sample covariance matrix $\widehat{\bm{\Sigma}}$ is a poor estimator of ${\bm{\Sigma}}$ when the dimension of the matrix and the sample size increase to infinity proportionally—in this setting, the spectral distribution of $\widehat{\bm{\Sigma}}$ follows the MP-Law (Marchenko & Pastur, 1967), which is drastically different from that of true ${\bm{\Sigma}}$. Therefore, in high dimensional settings, the sample covariance matrix is not a reliable estimator of the population covariance matrix for mean-variance portfolio optimization (Bai et al., 2009). In this paper, we only focus on the issue caused by high dimensional covariance estimation and assumes the knowledge of ${\bm{\mu}}$ for simplicity. Note that we care about covariance estimation via the specific function of Sharpe ratio, which can be different from the majority of the literature on estimating covariance matrix itself, whose performance is typically measured using various matrix norms.
1.1 Literature review
Estimation errors in ${\bm{\mu}}$ and ${\bm{\Sigma}}$ results in a substantial deterioration in out-of-sample portfolio performance, which is typically measured by Sharpe ratio. Effort invested in improving the out-of-sample Sharpe ratio associated with Markowitz portfolio has a long history (Brown, 1976; Klein & Bawa, 1976; Jobson & Korkie, 1981; Craig MacKinlay & Pástor, 2000). More recently, Kan & Zhou (2007) introduced a three-fund rule which can improve the out-of-sample portfolio performance in the case where a risk-free asset is available. DeMiguel, Garlappi & Uppal (2009) empirically demonstrated that the “$1/N$ rule” consistently outperforms various optimally estimated portfolios across a wide range of datasets. Tu & Zhou (2011) enhanced the three-fund rule by incorporating the “$1/N$ rule”.
A second approach to enhance the out-of-sample Sharpe ratio involves integrating various regularization techniques. Bai et al. (2009); Karoui (2010) found that using in-sample data for both creating and evaluating portfolios leads to an overly optimistic assessment of their true out-of-sample performance, particularly as the number of assets grows. Recognizing the need for regularization, several studies (Jagannathan & Ma, 2003; Fan et al., 2012; Bruder et al., 2013; Ao et al., 2019) have highlighted its significance in preventing the deterioration of portfolio performance with increasing number of assets. Jagannathan & Ma (2003) empirically showed the advantage of no-short-sale constraints when estimating optimal portfolios, with Fan et al. (2012) providing theoretical support for this strategy. Additionally, Bruder et al. (2013) reviewed various regularization methods that help stabilize mean-variance allocations, including weight constraints, resampling methods, and shrinkage techniques. DeMiguel, Garlappi, Nogales & Uppal (2009) considered a generalized approach to portfolio optimization by constraining different portfolio norms. More recently, Li et al. (2022) gave a spectral correction to the sample covariance matrix to more accurately estimate the out-of-sample Sharpe ratio when $p<n$. Additionally, Kan et al. (2024) explored the asymptotic distribution of both in-sample and out-of-sample Sharpe ratios under the same condition with Li et al. (2022). Ao et al. (2019) applied a Lasso penalty in portfolio optimization and estimated the maximum Sharpe ratio considering scenarios where the ratio $p/n$ approaches a constant $c\in(0,1)$.
Another line of research focuses on directly estimating the population covariance matrix. Current methods for high-dimensional population covariance estimation primarily utilize either truncation or shrinkage approaches. Truncation-based methods generally rely on a specific sparsity structure within the population covariance or precision matrix, aiming for more precise estimations by exploiting this sparsity (Bickel & Levina, 2008; Lam & Fan, 2009; Cai & Liu, 2011; Fan et al., 2011, 2016; Cai et al., 2020). These methods tend to achieve faster convergence rates or enhanced empirical performance, particularly when the sparsity assumption is valid. In contrast, shrinkage-based methods adjust the spectrum of the sample covariance matrix to improve the estimation of the population covariance without relying on sparsity (Ledoit & Wolf, 2004, 2012; Bodnar et al., 2022; Bodnar & Parolya, 2024). These techniques are largely driven by the bias-variance trade-off, aiming to minimize estimation variance and overall error by introducing a slight bias. For example, Ledoit & Wolf (2004) and Ledoit & Wolf (2012) developed a linear and nonlinear shrinkage estimator on the empirical eigenvalue spectrum that offers improved conditioning and accuracy compared to the traditional sample covariance matrix. In addition to correcting sample covariance matrix, factor models and its low-rank plus sparse covariance structure are also extensively studied in the literature. Readers may refer to Fan et al. (2021) to see an overview of factor models and their applications.
1.2 Our contributions
Several methodologies have been developed for estimating the out-of-sample Sharpe ratio. This raises an interesting question:
Is it possible to estimate the out-of-sample Sharpe ratio with some regularization using in-sample data? Can we then optimize the estimator over the regularization parameter to enhance the out-of-sample Sharpe ratio?
In this paper, we investigate mean-variance portfolio optimization with ridge regularization. Specifically, under the setting where the unknown population covariance matrix is estimated by the sample covariance, we develop a novel method for estimating the out-of-sample Sharpe ratio for the ridge penalized mean-variance portfolio. Our method can be applied to both the settings with and without the risk-free asset: with the risk-free asset, our approach directly proposes an estimator for the out-of-sample Sharpe ratio of the regularized mean-variance portfolio; without the risk-free asset, our approach is given in the form of estimating the efficient frontier, that is, for any given target return, instead of estimating the Sharpe ratio, we equivalently estimate the variance of the regularized mean-variance portfolio. Remarkably, the estimators given in our paper are consistent – while they can be calculated without the knowledge of the population covariance matrix, they still share the same asymptotic behavior as the portfolio optimization solution calculated with the knowledge of the population covariance matrix. Our approach optimizes portfolios using in-sample data and provides accurate performance evaluations. We summarize our contributions as follows:
  1. 1.
    We introduce an innovative in-sample method that estimates the true out-of-sample variance and thus the true out-of-sample Sharpe ratio of a portfolio, considering both scenarios with and without the risk-free asset. With no risk-free asset, we actually provide a way to accurately estimate the efficient frontier. This method provides an effective framework for evaluating portfolio performance in the future, backed by theoretical guarantees of almost surely convergence.
  2. 2.
    Additionally, our method assists in the selection of optimal regularization parameters from a predetermined set, enhancing the precision of investment decisions. By providing valid estimations for the future performances of portfolios under different regularization levels, our method allows data-driven decision making for the regularization parameters, which can be particularly useful in situations where market conditions change rapidly and allows the portfolio allocation to be updated in a timely fashion.
  3. 3.
    From a technical standpoint, our approach establishes theoretical convergence under relatively mild conditions. For instance, our assumption on the population covariance matrix ${\bm{\Sigma}}$ is less restrictive compared to the classical setting, which typically requires the bound of $\|{\bm{\Sigma}}\|_{{\mathop{\mathrm{op}}}}$. In our method, we allow $\|{\bm{\Sigma}}\|_{{\mathop{\mathrm{op}}}}$ to tend towards infinity, expanding the applicability of our approach. In particular, the factor model implies a covariance structure with diverging $\|{\bm{\Sigma}}\|_{{\mathop{\mathrm{op}}}}$ and our method still works under the factor models, as long as either $c<1$ or the number of diverging spikes is fixed.
Compared to the work with lasso regularization (Ao et al., 2019), our work differs from theirs in a few salient aspects. Firstly, we consider ridge regularization under more general and relaxed conditions allowing diverging spikes in population covariance matrix with arbitrary orders. Moreover, when factors exist, they consider holding factor portfolios besides individual stocks, although factor portfolios may not always be accessible. In contrast, we only allocate portfolio among stocks. Most importantly, they only give the estimation of the maximum value of Sharpe ratio, while we provide valid estimations for any regularization $q$, so that we can select the optimal tuning $q$. It is unclear how the Lasso tuning affects Sharpe ratio in Ao et al. (2019).
1.3 Paper organization and notations
The remainder of the paper is organized as follows. We first introduce some notations below. In Section 2 and Section 3, we give the problem settings for the regularized mean-variance portfolio optimization with and without risk free assets, and propose our methodology to estimate the Sharpe Ratio respectively. We then present simulation results in Section 4 to support our theoretical findings. Section 5 contains our demonstration about the effectiveness of our method for optimizing mean-variance portfolios using real financial return data. Finally, we conclude the paper in Section 6 and discuss potential future research. Additional experiments and proof details are provided in the supplementary material.
We employ lower case letters (e.g. $a$) for scalar values. Boldface letters are used to denote vectors and matrices (e.g. $\mathbf{a}$ and $\mathbf{A}$). For a set $I$ of indices, we define ${\bm{1}}(I)$ such that ${\bm{1}}(i)=1$ if $i\in I$, and ${\bm{1}}(i)=0$ if $i\notin I$. Hence, $\mathbf{a}(i)$ specifically refers to the $i$-th element in the vector $\mathbf{a}$. Moreover, the vector $\text{\bf 1}\in\mathbb{R}^{p}$ ($\text{\bf 1}_{n}\in\mathbb{R}^{n}$) denotes the vector with all coordinates equal to 1. For a square matrix $\mathbf{A}$, we use $\lVert\mathbf{A}\rVert_{{\mathrm{tr}}}$ and $\lVert\mathbf{A}\rVert_{{\mathop{\mathrm{op}}}}$ to denote its trace norm and operator norm, respectively, and use ${\mathrm{tr}}(\mathbf{A})$ to denote its trace. The sets of natural, real and complex numbers are denoted by $\mathbb{N}$, $\mathbb{R}$ and $\mathbb{C}$, respectively. The set of integers from $n_{1}$ to $n_{2}$ is denoted by $[n_{1}:n_{2}]=\{n_{1},\ldots,n_{2}\}$ and $[n]=[1:n]=\{1,\ldots,n\}$.
2 High Dimensional Estimation of Sharpe Ratio
2.1 Mean-variance portfolio and Sharpe ratio
In this section, we introduce the problem setting we consider in this paper. We first review mean-variance portfolio optimization under the existence of the risk-free rate $r_{0}\geq 0$.
Definition 2.1 (Mean-variance portfolio optimization).
Given $p$ risky assets with mean $\mathbf{r}\in\mathbb{R}^{p}$ and covariance matrix ${\bm{\Sigma}}\in\mathbb{R}^{p\times p}$, the mean-variance portfolio optimization is to optimize the allocation vector $\mathbf{w}$, which satisfies
$\mathbf{w}^{*}=\operatorname{argmin}\limits_{\mathbf{w}\in\mathbb{R}^{p}}\mathbf{w}^{\top}{\bm{\Sigma}}\mathbf{w},\quad\operatorname{s.t.}\mathbf{w}^{\top}{\bm{\mu}}=\mu_{0}.$
Here, we denote by ${\bm{\mu}}=\mathbf{r}-r_{0}\text{\bf 1}$ the excess return of the risky assets, and $\mu_{0}>0$ is the targeted excess return of the portfolio.
The motivation behind Definition 2.1 is clear. Each allocation vector $\mathbf{w}$ represents a unique portfolio, and the optimal $\mathbf{w}$ in Definition 2.1 minimizes the variance of the portfolio with an expected excess return. It is worth noting that in the mean-variance optimization, it is not required to have $\mathbf{w}^{\top}{\bm{1}}=1$, because the existence of a risk-free asset makes investing or borrowing risk-less bond possible. To be exact, we could allocate $w_{0}=1-\mathbf{w}^{\top}{\bm{1}}$ to the risk-free asset, so that the total mean excess return is indeed $(\mathbf{w}^{\top}\mathbf{r}+w_{0}r_{0})-r_{0}=\mathbf{w}^{\top}{\bm{\mu}}$.
The solution in Definition 2.1 is $\mathbf{w}^{*}\propto{\bm{\Sigma}}^{-1}{\bm{\mu}}$. Although the theory behind mean-variance portfolio optimization is elegant, in practice, a portfolio manager needs to estimate ${\bm{\Sigma}}$ in order to find out the optimal mean-variance portfolio $\mathbf{w}^{*}$. However, the estimation of ${\bm{\Sigma}}$ is very challenging due to the huge number of unknown parameters. To further investigate the effect of estimation error of ${\bm{\Sigma}}$, we model the sample data as follows:
Assumption 2.2.
Given the mean return vector $\mathbf{r}\in\mathbb{R}^{p}$ and the covariance matrix ${\bm{\Sigma}}\in\mathbb{R}^{p\times p}$, we consider the observed data matrix $\mathbf{R}\in\mathbb{R}^{n\times p}$ with the form
$\mathbf{R}={\bm{1}}_{n}\mathbf{r}^{\top}+\mathbf{X},$
where $\mathbf{X}=\mathbf{Z}{\bm{\Sigma}}^{\frac{1}{2}}\in\mathbb{R}^{n\times p}$. Here, the elements in the matrix $\mathbf{Z}\in\mathbb{R}^{n\times p}$ are i.i.d random variables with zero mean, variance 1 and finite $(8+\varepsilon)$-order moment for some $\varepsilon>0$.
In this work, we assume the knowledge of $r_{0}$, ${\bm{\mu}}$ (or equivalently $\mathbf{r}$), and only focus on the effect of estimation error from estimating ${\bm{\Sigma}}$ using the sample covariance $\widehat{\bm{\Sigma}}=\mathbf{X}^{\top}\mathbf{X}/n$, where $\mathbf{X}$ is defined in Assumption 2.2. It has been widely known that when dimension $p$ is large with $p/n\to c>0$, that is $p$ grows with the sample size linearly, the sample covariance matrix is not a good estimator for the true covariance matrix. To mitigate the bad condition number of the sample covariance, it is natural to consider adding regularization in the portfolio construction. We add a regularization term $\mathbf{Q}$ to $\widehat{\bm{\Sigma}}$ in constructing the mean-variance portfolio. Therefore, we consider the following optimization:
$\mathbf{w}^{*}=\operatorname{argmin}\limits_{\mathbf{w}\in\mathbb{R}^{p}}\mathbf{w}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})\mathbf{w},\quad\operatorname{s.t.}\mathbf{w}^{\top}{\bm{\mu}}=\mu_{0}.$
The optimal $\mathbf{w}^{*}$ satisfies
$\mathbf{w}^{*}\propto(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\mu}}.$
Here, $\mathbf{Q}$ is a positive definite matrix. The Sharpe ratio of $\mathbf{w}^{*}$ is defined as the mean of the excessive return of the portfolio over its standard deviation. Mathematically,
$SR(\mathbf{Q})=\frac{\mathbb{E}_{\widetilde{\mathbf{R}}}[\mathbf{w}^{*\top}(\widetilde{\mathbf{R}}-r_{0}\text{\bf 1})]}{\sqrt{\operatorname{{\rm Var}}_{\widetilde{\mathbf{R}}}[\mathbf{w}^{*\top}(\widetilde{\mathbf{R}}-r_{0}\text{\bf 1})]}}=\frac{{\bm{\mu}}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\mu}}}{\sqrt{{\bm{\mu}}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\mu}}}}.$ (2.1)
Here, $\widetilde{\mathbf{R}}$ is an out-of-sample data point with mean return $\mathbf{r}={\bm{\mu}}+r_{0}\text{\bf 1}$ and covariance ${\bm{\Sigma}}$. Note that the randomness from historical data in $\mathbf{w}^{*}$ is treated as given in the Sharpe computation in (2.1). Define the following quantities:
$T_{n,1}(\mathbf{Q})$ $={\mathrm{tr}}\bigg{[}\bigg{(}\frac{\mathbf{X}^{\top}\mathbf{X}}{n}+\mathbf{Q}\bigg{)}^{-1}\mathbf{A}\bigg{]},$ (2.2)
$T_{n,2}(\mathbf{Q})$ $={\mathrm{tr}}\bigg{[}\bigg{(}\frac{\mathbf{X}^{\top}\mathbf{X}}{n}+\mathbf{Q}\bigg{)}^{-1}{\bm{\Sigma}}\bigg{(}\frac{\mathbf{X}^{\top}\mathbf{X}}{n}+\mathbf{Q}\bigg{)}^{-1}\mathbf{A}\bigg{]},$ (2.3)
for some deterministic matrix $\mathbf{A}\in\mathbb{R}^{p\times p}$. Hence, by setting $\mathbf{A}={\bm{\mu}}{\bm{\mu}}^{\top}$, it is easy to see that the Sharpe Ratio (SR) in (2.1) could be expressed as
$SR(\mathbf{Q})=\frac{T_{n,1}(\mathbf{Q})}{\sqrt{|T_{n,2}(\mathbf{Q})|}}.$ (2.4)
Remark 2.3.
We introduce a deterministic matrix $\mathbf{A}\in\mathbb{R}^{p\times p}$ to obtain results more general than the choice of $\mathbf{A}={\bm{\mu}}{\bm{\mu}}^{\top}$. Here we have the underlying assumption that the mean vector of financial returns keeps unchanged over the time period under consideration. However, if daily excess returns have time-varying means, denoted as ${\bm{\mu}}_{i}$, where $i$ represents the trading date, and if the daily returns are independent across dates, the annual out-of-sample Sharpe ratio of the daily-balanced portfolio $\mathbf{w}_{i}^{*}\propto(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\mu}}_{i}$ can be expressed via setting $\mathbf{A}=\sum_{i}{\bm{\mu}}_{i}{\bm{\mu}}_{i}^{\top}$ over all trading dates.
To obtain the largest out-of-sample Sharpe ratio, one might select the regularization matrix $\mathbf{Q}$ by a cross-validation procedure, but this is on the one hand computationally expensive, on the other hand not straightforward to do for a time series setting in practice. Note that although for theoretical analysis, we assume data are independent from time to time, but in practice, it is not recommended to shuffle time order in the computation of out-of-sample Sharpe ratio. In this work, we hope to directly propose a method to estimate Sharpe ratio using in-sample data, which can be used as a good criterion to select the best regularization $\mathbf{Q}$ to guarantee good out-of-sample Sharpe performance of the regulated mean-variance portfolio above. In the next section, without any information of the population covariance matrix ${\bm{\Sigma}}$, we propose a statistic which allows us to apply in-sample data to effectively estimate the Sharpe ratio $SR(\mathbf{Q})$ in (2.4).
2.2 Out-of-sample estimation of Sharpe ratio
In this section, we present our proposed estimator of $SR(\mathbf{Q})$ and its main theoretical results. We need the following assumptions.
Assumption 2.4.
The portfolio dimension $p$ and the sample size $n$ both tend to infinity, with the regime of $p/n\to c$ where $c>0$ is a constant.
Assumption 2.4 provides a classic framework for high dimensional analysis, commonly used in random matrix theory (Yao et al., 2015). Next, we make assumptions about the matrix $\mathbf{A}$ and the structure of the regularization matrix $\mathbf{Q}$.
Assumption 2.5.
The matrices $\mathbf{A}$ is deterministic. The regularization matrix $\mathbf{Q}$ is selected from some candidate set $\mathcal{Q}$, in which all the elements are independent with $\mathbf{X}$. For any $\mathbf{Q}\in\mathcal{Q}$, there exist constants $c_{\mathbf{Q}},C_{\mathbf{Q}}>0$ such that $c_{\mathbf{Q}}\leq\lambda_{\min}(\mathbf{Q})$ and $\|{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{Q}{\bm{\Sigma}}^{-\frac{1}{2}}\|_{{\mathop{\mathrm{op}}}}\leq C_{\mathbf{Q}}$ for any sequences $(n,p)$.
Assumption 2.5 is reasonable, as it makes no sense to apply an ill-conditioned regularization matrix $\mathbf{Q}$. The condition $\|{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{Q}{\bm{\Sigma}}^{-\frac{1}{2}}\|_{{\mathop{\mathrm{op}}}}\leq C_{\mathbf{Q}}$ can also hold as long as $\lambda_{\min}^{-1}({\bm{\Sigma}})$ and $\|\mathbf{Q}\|_{{\mathop{\mathrm{op}}}}$ are upper bounded by constants, which are quite mild assumptions. We then make assumptions regarding the structure of ${\bm{\Sigma}}$.
Assumption 2.6.
${\bm{\Sigma}}$ is well scaled as $\|{\bm{\Sigma}}/p\|_{{\mathrm{tr}}}\leq C$ for some constant $C>0$. Denote by $\lambda_{1}\geq\cdots\geq\lambda_{p}$ the eigenvalues of ${\bm{\Sigma}}$. One of the following cases must hold:
  1. 1.
    (Bounded spectrum) There exists $C_{1}>$ such that $0<\lambda_{1}\leq C_{1}<+\infty$ for any sequences $(n,p)$.
  2. 2.
    (Arbitrary number of diverging spikes when $c<1$) $p/n\to c<1$ and we allow arbitrary number of top eigenvalues goes to infinity. It is assumed that there exists constant $c_{1}>0$ such that $\lambda_{p}\geq c_{1}$.
  3. 3.
    (Fixed number of diverging spikes when $c\geq 1$) $p/n\to c\geq 1$ and we let the number of diverging spikes be $K$, $K$ is fixed and $\lambda_{1}/\lambda_{K}\leq C_{\lambda}$ for some constant $C_{\lambda}>0$. Moreover, it is assumed that there exists a constant $c_{1}>0$ such that $\lambda_{p}\geq c_{1}$.
Assumption 2.6 includes three cases: the first case is where ${\bm{\Sigma}}$ has bounded operator norm, and the second/third case is where ${\bm{\Sigma}}$ has spikes that tend to infinity. This encompasses a wide range of structures of the covariance matrix ${\bm{\Sigma}}$. Specifically, we allow the number of spikes to tend to infinity when $c<1$, with flexibility in the rate at which they tend to infinity. While the condition becomes slightly stronger for $c\geq 1$, it remains mild in practical settings as we only need to make sure we have a fixed number of diverging eigenvalues. The main results are given in the following theorem.
Theorem 2.7.
Suppose Assumptions 2.2, 2.4, 2.5 and 2.6 hold. For any $\mathbf{Q}\in\mathcal{Q}$, a good estimator $\widehat{SR}(\mathbf{Q})$ for $SR(\mathbf{Q})$ which is defined in (2.4), is given as follows.
$\widehat{SR}(\mathbf{Q})=\frac{T_{n,1}(\mathbf{Q})}{\sqrt{\big{|}\widehat{T}_{n,2}(\mathbf{Q})\big{|}}},\quad\text{where}\quad\widehat{T}_{n,2}(\mathbf{Q})=\frac{{\mathrm{tr}}(\widehat{{\bm{\Sigma}}}+\mathbf{Q})^{-1}\widehat{\bm{\Sigma}}(\widehat{{\bm{\Sigma}}}+\mathbf{Q})^{-1}\mathbf{A}}{\big{(}1-\frac{c}{p}{\mathrm{tr}}\widehat{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\big{)}^{2}}.$
If $\mathbf{A}$ is semi-positive definite, it holds that
$\widehat{SR}(\mathbf{Q})/SR(\mathbf{Q})~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1.$
If additionally $\|\mathbf{A}\|_{{\mathrm{tr}}}$ is bounded [^1] The condition can be further relaxed to ${\mathrm{tr}}\big{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\big{)}^{-1}\mathbf{A}$ is bounded, where $s_{0}>0$ is defined by $s_{0}=\frac{c}{p}{\mathrm{tr}}~{}{\bm{\Sigma}}\big{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\big{)}^{-1}$ in Assumption 3.1 below., then $SR(\mathbf{Q})$ is almost surely bounded and
$\widehat{SR}(\mathbf{Q})-SR(\mathbf{Q})~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0.$
Remark 2.8.
The proof of Theorem 2.7 is given in Appendix D in the supplementary material. In the proof, we present several extensions of Lemma C.1, which is a key result from random matrix theory. The first challenge in the extension arises from the relaxation of bounded $\|{\bm{\Sigma}}\|_{{\mathop{\mathrm{op}}}}$ in Lemma C.1. When spikes exist and $\|{\bm{\Sigma}}\|_{{\mathop{\mathrm{op}}}}$ is no longer bounded, we cannot apply Lemma C.1 and have to extend Lemma C.1 to Lemmas C.6 and C.7, and prove that $z\to 0$ and $n\to+\infty$ can be interchanged. The second challenge arises from obtaining ratio consistency as well as difference consistency. As we see, ratio consistency has the advantage of circumventing additional scaling condition (i.e. bounded $\|\mathbf{A}\|_{{\mathrm{tr}}}$). To obtain ratio consistency, we establish Lemma C.9 even when $\|{\bm{\Sigma}}\|_{{\mathop{\mathrm{op}}}}$ is unbounded.
Theorem 2.7 establishes an estimator for $SR(\mathbf{Q})$ with almost sure convergence guarantee in a wide range of scenarios. More importantly, it provides a tractable framework for portfolio managers to determine the best regularization matrix $\mathbf{Q}$ over some given candidate set $\mathcal{Q}$. Consider a candidate set $\mathcal{Q}$ from which we want to select $\mathbf{Q}\in\mathcal{Q}$ in order to generate the maximum out-of-sample Sharpe ratio. It holds from Theorem 2.7 that $\widehat{SR}(\mathbf{Q})/SR(\mathbf{Q})~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1$, hence we can simply maximize $\widehat{SR}(\mathbf{Q})$ over $\mathcal{Q}$ to obtain the optimal $\mathbf{Q}$. However, it is also important to note that optimizing the entire matrix $\mathbf{Q}$ without the constraint $\mathbf{Q}\in\mathcal{Q}$ to obtain the maximum value of $SR(\mathbf{Q})$ is not feasible. In other words, if we define $\widehat{\mathbf{Q}}=\arg\max_{\mathbf{Q}}\widehat{SR}(\mathbf{Q})$, we cannot guarantee that $SR(\widehat{\mathbf{Q}})$ will also be large, that is, $\widehat{SR}(\widehat{\mathbf{Q}})$ may not align well with $SR(\widehat{\mathbf{Q}})$. In our theorem, $\mathbf{Q}$ is independent of $\mathbf{X}$, but here $\widehat{\mathbf{Q}}$ is derived from the observed data matrix $\mathbf{X}$. Nonetheless, the estimator is practically useful to maximize $SR(\mathbf{Q})$ within a suitable candidate set $\mathcal{Q}$ without access to the true ${\bm{\Sigma}}$. For instance, we can let $\mathcal{Q}$ be the set that exists countably many deterministic matrices. The final important comment is that, with well designed $\mathbf{Q}$, we show in Appendix F that the portfolio optimization can indeed achieve the maximum population SR $=\sqrt{{\bm{\mu}}^{\top}{\bm{\Sigma}}^{-1}{\bm{\mu}}}$.
3 High Dimensional Estimation of Efficient Frontier
3.1 Efficient frontier with no risk-free asset
In this section, we give a brief introduction of the efficient frontier with no risk-free asset. Suppose we have $p$ risky assets with mean $\mathbf{r}\in\mathbb{R}^{p}$ and covariance ${\bm{\Sigma}}\in\mathbb{R}^{p\times p}$, but no risk-free asset is available. Given target return $\mu_{0}>0$, the portfolio optimization is given by
$\mathbf{w}^{*}=\arg\min_{\mathbf{w}}\mathbf{w}^{\top}{\bm{\Sigma}}\mathbf{w},\quad\text{s.t. }\mathbf{w}^{\top}\mathbf{r}=\mu_{0}\text{ and }\mathbf{w}^{\top}{\bm{1}}=1.$ (3.1)
Note that $\mu_{0}$ is the target excess return in Section 2,  while here it is the target raw return, with a slight abuse of notation. The solution $\mathbf{w}^{*}$ from (3.1) is straightforward. However, in practice, ${\bm{\Sigma}}$ is not observable and its sample estimation $\widehat{\bm{\Sigma}}=\mathbf{X}^{\top}\mathbf{X}/n$ can deviate significantly from the true ${\bm{\Sigma}}$ in high dimensions. Similar to Section 2.1, the observed sample data $\mathbf{R}$ following Assumption 2.2 can be expressed as $\mathbf{R}=\text{\bf 1}_{n}\mathbf{r}^{\top}+\mathbf{X}$. Again we consider adding a ridge term $\mathbf{Q}>0$ in the portfolio optimization problem:
$\mathbf{w}^{*}=\arg\min_{\mathbf{w}}\mathbf{w}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})\mathbf{w},\quad\text{s.t. }\mathbf{w}^{\top}\mathbf{r}=\mu_{0}\text{ and }\mathbf{w}^{\top}{\bm{1}}=1.$ (3.2)
From Merton (1972), the optimal $\mathbf{w}^{*}$ in (3.2) is given by $\mathbf{w}^{*}=\mathbf{g}+\mu_{0}\cdot\mathbf{h}$, where
$\hfil~{}&\mathbf{g}=D^{-1}\big{[}B(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{1}}-A(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\mathbf{r}\big{]},~{}\mathbf{h}=D^{-1}\big{[}C(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\mathbf{r}-A(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{1}}\big{]},\hfil~{}\\ \hfil~{}&A=\mathbf{r}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{1}},~{}B=\mathbf{r}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\mathbf{r},~{}C={\bm{1}}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{1}},~{}D=BC-A^{2}.\hfil~{}\end{array}$ (3.5)
Using the solution $\mathbf{w}^{*}$, the efficient frontier can be represented as
$\sigma_{0}^{2}=\mathbf{w}^{*\top}{\bm{\Sigma}}\mathbf{w}^{*}=(\mathbf{g}+\mu_{0}\cdot\mathbf{h})^{\top}{\bm{\Sigma}}(\mathbf{g}+\mu_{0}\cdot\mathbf{h}),$ (3.6)
which should be viewed as the curve of $(\sigma_{0},\mu_{0})$ as we change the target return $\mu_{0}$. Our objective is to estimate this curve represented by (3.6) for a given regularization matrix $\mathbf{Q}$.
3.2 Out-of-sample estimation of frontier variances
In this section, we present the main results on estimating the efficient frontier. We give an additional technical assumption on $\mathbf{r}$, assuming that the vectors $\mathbf{r}$ and ${\bm{1}}$ do not coincide, which makes $\mu_{0}=1$ so that we do not have a well-posed problem.
Assumption 3.1.
Let $s_{0}>0$ to be the unique solution of the equation [^2] The uniqueness and existence can be found in Lemma C.5.
$s_{0}=\frac{c}{p}{\mathrm{tr}}~{}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}.$
Define
$\mathcal{A}_{rr}=\mathbf{r}^{\top}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\mathbf{r},\quad\mathcal{A}_{r1}=\mathbf{r}^{\top}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}{\bm{1}},\quad\mathcal{A}_{11}={\bm{1}}^{\top}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}{\bm{1}}.$
It is assumed that there exists a constant $\rho<1$ such that $\mathcal{A}_{r1}^{2}/(\mathcal{A}_{11}\mathcal{A}_{rr})\leq\rho<1.$
This technical assumption states that the vectors $\mathbf{r}$ and ${\bm{1}}$ do not coincide under the inner product $\langle\mathbf{a},\bm{b}\rangle=\mathbf{a}\big{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\big{)}^{-1}\bm{b}$, which is a mild assumption. In fact, Cauchy-Schwartz inequality gives $\frac{\mathcal{A}_{r1}^{2}}{\mathcal{A}_{rr}\mathcal{A}_{11}}\leq 1$. The assumption $\frac{\mathcal{A}_{r1}^{2}}{\mathcal{A}_{rr}\mathcal{A}_{11}}\leq\rho<1$ is a direct relaxation of the upper bound $1$. Given this additional assumption, we have the following theorem.
Theorem 3.2.
Suppose that Assumptions 2.2, 2.4-2.6 and 3.1 hold. Recall $\sigma_{0}^{2}$ is defined in (3.2). Define
$\widehat{\sigma}^{2}=\frac{(\mathbf{g}+\mu_{0}\mathbf{h})^{\top}\widehat{\bm{\Sigma}}(\mathbf{g}+\mu_{0}\mathbf{h})}{(1-c/p\cdot{\mathrm{tr}}\widehat{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1})^{2}}\,,$
where $\mathbf{g}$ and $\mathbf{h}$ are defined in (3.5), it holds that
$\widehat{\sigma}^{2}/\sigma_{0}^{2}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1.$
Moreover, the following properties hold:
  1. 1.
    If $\mathcal{A}_{rr}$ is bounded, then for any $r_{0}=O(\mu_{0})$ it holds that $\frac{\mu_{0}-r_{0}}{\sigma_{0}}-\frac{\mu_{0}-r_{0}}{\widehat{\sigma}}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0$.
  2. 2.
    If $\mu_{0}\leq C\sqrt{\mathcal{A}_{rr}}$ for some large constant $C>0$, then $\widehat{\sigma}^{2}-\sigma_{0}^{2}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0$.
The proof of Theorem 3.2 can be found in Appendix E in the supplementary material. It is worth noting that when the target return $\mu_{0}$ is given, estimating the volatility naturally leads to an estimator of Sharpe ratio. The optimal true Sharpe ratio $SR(\mathbf{Q})$ without the risk-free asset equals $(\mu_{0}-r_{0})/\sigma_{0}$, and Theorem 3.2 implies that the estimator $\widehat{SR}(\mathbf{Q})=(\mu_{0}-r_{0})/\widehat{\sigma}$ is ratio-consistent without additional scaling condition, and is also difference-consistent when $\mathcal{A}_{rr}$ is bounded. In Theorem 2.7, if we set $\mathbf{A}=\mathbf{r}\mathbf{r}^{\top}$, the relaxed condition that ${\mathrm{tr}}\Big{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\Big{)}^{-1}\mathbf{A}$ is bounded (see the footnote for Theorem 2.7) for the difference-consistency of Sharpe ratio is exactly the same as the condition that $\mathcal{A}_{rr}$ is bounded in Theorem 3.2.
Albeit the conditions are similar, we hope to point out new challenges in proving Theorem 3.2. The first challenge arises from the presence of distant diverging spikes in the covariance matrix ${\bm{\Sigma}}$ in Case 2 and Case 3 of Assumption 2.6. It is possible that the vectors $\mathbf{r}$ and ${\bm{1}}$ lie within the space spanned by the eigenvectors of these spikes. Note that in Section 2, due to the existence of the risk-free asset, we directly work with the exccess return ${\bm{\mu}}$ without necessity to discuss the relationship between $\mathbf{r}$ and ${\bm{1}}$. For instance, we consider the following 4 concrete cases with the assumption $\mathbf{r}=a\cdot{\bm{1}}+\mathbf{r}^{\perp}$ where $\mathbf{r}^{\perp}\perp{\bm{1}}$, and the distant spikes in ${\bm{\Sigma}}/(1+s_{0})+\mathbf{Q}$ are all in the order of $p$.
  • Case 1: ${\bm{1}}$ and $\mathbf{r}^{\perp}$ are both orthogonal to the space of distant spikes, or no spikes.
  • Case 2: ${\bm{1}}$ belongs to, and $\mathbf{r}^{\perp}$ is orthogonal to the space of distant spikes.
  • Case 3: $\mathbf{r}^{\perp}$ belongs to, and ${\bm{1}}$ is orthogonal to the space of distant spikes.
  • Case 4: ${\bm{1}}$ and $\mathbf{r}^{\perp}$ both belong to the space of distant spikes.
Table 1: Different rates of $\mathcal{A}_{rr}$ and $\mathcal{A}_{11}$ for specific cases.
Case 1 Case 2 Case 3 Case 4
$\|\mathbf{r}^{\perp}\|_{2}=1$ $\|\mathbf{r}^{\perp}\|_{2}=\sqrt{p}$ $\|\mathbf{r}^{\perp}\|_{2}=1$ $\|\mathbf{r}^{\perp}\|_{2}=\sqrt{p}$ $\|\mathbf{r}^{\perp}\|_{2}=1$ $\|\mathbf{r}^{\perp}\|_{2}=\sqrt{p}$ $\|\mathbf{r}^{\perp}\|_{2}=1$ $\|\mathbf{r}^{\perp}\|_{2}=\sqrt{p}$
Order of $\mathcal{A}_{rr}$ $\Theta(a^{2}p+1)$ $\Theta(a^{2}p+p)$ $\Theta(a^{2}+1)$ $\Theta(a^{2}+p)$ $\Theta\big{(}a^{2}p+\frac{1}{p})$ $\Theta(a^{2}p+1)$ $\Theta\big{(}a^{2}+\frac{1}{p}\big{)}$ $\Theta(a^{2}+1)$
Order of $\mathcal{A}_{11}$ $\Theta(p)$ $\Theta(p)$ $\Theta(1)$ $\Theta(1)$ $\Theta(p)$ $\Theta(p)$ $\Theta(1)$ $\Theta(1)$
$\frac{\widehat{\sigma}^{2}}{\sigma_{0}^{2}}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$
$\frac{\mu_{0}-r_{0}}{\sigma_{0}}-\frac{\mu_{0}-r_{0}}{\widehat{\sigma}}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0$ $a=0$ $\surd$ - $\surd$ - $\surd$ $\surd$ $\surd$ $\surd$
$a=\frac{1}{\sqrt{p}}$ $\surd$ - $\surd$ - $\surd$ $\surd$ $\surd$ $\surd$
$a=1$ $\surd$ - - - $\surd$ $\surd$
Range of $\mu_{0}$ $O(\sqrt{a^{2}p+1})$ $O(\sqrt{a^{2}p+p})$ $O(\sqrt{a^{2}+1})$ $O(\sqrt{a^{2}+p})$ $O\big{(}\sqrt{a^{2}p+\frac{1}{{p}}}\big{)}$ $O(\sqrt{a^{2}p+1})$ $O\big{(}\sqrt{a^{2}+\frac{1}{{p}}}\big{)}$ $O(\sqrt{a^{2}+1})$
for $\widehat{\sigma}^{2}-\sigma_{0}^{2}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0$
The symbol $\surd$” indicates that the convergence is proven to hold.
Table 1 gives the different rates of $\mathcal{A}_{rr}$ and $\mathcal{A}_{11}$ for the above 4 cases, all of which are covered in Theorem 3.2. From Table 1, it is obvious that the ratio-consistency always holds without additional conditions. The Sharpe difference-consistency holds only when $\mathcal{A}_{rr}$ is bounded, which becomes harder to be satisfied if we bet the market factor ${\bm{1}}$ in the mean return. This partially explains why portfolio managers eagerly pursue to be market neutral, that is $a=0$. Furthermore, the target return $\mu_{0}$ needs to be set within a proper range in order for the corresponding volatility to be estimated accurately in terms of the absolute error. We can usually allow $\mu_{0}$ to be on the constant order, except in Case 3 and Case 4 when $\|\mathbf{r}^{\perp}\|=1$ and $a=0$, which represents a very low signal level in $\mathbf{r}$.
Handling general $\mathcal{A}_{rr},\mathcal{A}_{11},\mathbf{r}$ and $\mu_{0}$ with arbitrary rates poses great challenge in the analysis. In the proof of Theorem 3.2, we derive the orders of all the relevant quantities in different cases in terms of $\mathcal{A}_{rr}$, $\mathcal{A}_{11}$ and $\mu_{0}$, without specifying their specific statistical rates. Additionally, we need to divide the proof into two separate cases where $\mu_{0}\leq C\sqrt{\mathcal{A}_{rr}/\mathcal{A}_{11}}$ and $\mu_{0}\geq C\sqrt{\mathcal{A}_{rr}/\mathcal{A}_{11}}$, where distinct analyses are required.
In addition to the above order-related challenge mentioned, the expression of $\sigma_{0}^{2}$ presents another obstacle. While $\sigma_{0}^{2}$ can be expressed as ${\bm{\xi}}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\xi}}$ for some vector ${\bm{\xi}}$, similar to (2.3) in Section 2, it is important to note that ${\bm{\xi}}$ depends on $\widehat{\bm{\Sigma}}$. To tackle this nontrivial dependency, we need to carefully substitute ${\bm{\xi}}$ with a deterministic counterpart and make sure this does not affect the almost sure convergence. More proof details can be found in Appendix E in the supplementary material.
4 Numerical Experiments
In this section, we perform numerical experiments to validate our theoretical findings. In Section 4.1, we provide the basic settings for SR estimation and verify Theorem 2.7, especially the asymptotic behavior as $n$ increases. In Section 4.2, we verify our estimation of the efficient frontier in Theorem 3.2 as we change target portfolio returns. Some additional simulation settings are presented in Appendix A to provide more comparisons on different choices of the population covariance ${\bm{\Sigma}}$, the mean vector of returns ${\bm{\mu}}$ and the regularization term $\mathbf{Q}$. In all the simulations, we set $r_{0}=0$ for simplicity and in this case, $\mathbf{r}={\bm{\mu}}$.
4.1 Sharpe ratio estimation
We present the simulation results under the basic settings. We first fix the sample size as $n=1500$ and consider two dimensions: $p=750$ (ratio $c=1/2$) and $p=2250$ (ratio $c=3/2$). These cases correspond to values of $c$ less than and greater than 1. To simulate financial returns, we construct the population covariance matrix ${\bm{\Sigma}}_{0}$, mean vector ${\bm{\mu}}_{0}$, and the matrix $\mathbf{Q}_{0}$ as follows:
  1. 1.
    The population covariance ${\bm{\Sigma}}={\bm{\Sigma}}_{0}={\rm diag}(\lambda_{1},\dots,\lambda_{p})+2{\bm{1}}{\bm{1}}^{\top}$, where $\{\lambda_{i}\}_{i=1}^{p}$ are generated from a truncated $\Gamma^{-1}(1,1)$ distribution, truncated with the interval $[0.01,9]$, and then ranked in decreasing order. This particular construction of ${\bm{\Sigma}}$ follows from the factor model, mimicking the well-accepted low-rank plus sparse structure of the covariance matrix of financial returns. Specifically, the covariance consists of the market systematic risk, where the vector ${\bm{1}}$ represents a market factor introducing a spike in covariance spectrum, and the specific risk, where $\lambda_{i}$ represents the individual risk of each asset.
  2. 2.
    Assume the risk-free return is $r_{0}=0$. The mean vector of returns is given by ${\bm{\mu}}_{0}=\sqrt{5/p}\cdot({\bm{1}}(S_{+})-{\bm{1}}(S_{-}))\in\mathbb{R}^{p}$. Here, ${\bm{1}}(S)$ is defined in Subsection 1.3. The subsets $S_{+}$ and $S_{-}$ are randomly selected subsets of $[p]$ with $|S_{+}|=|S_{-}|=p/10$ and $S_{+}\cup S_{-}=\emptyset$. This formulation of ${\bm{\mu}}_{0}$ is motivated by the idea that a portion of the assets may yield low positive or negative returns, known as the “alpha” signal in asset pricing, while the majority assets simply have zero mean, ensuring no arbitrage opportunities.
  3. 3.
    The regularization matrix is $\mathbf{Q}=q\cdot\mathbf{Q}_{0}$. The matrix $\mathbf{Q}_{0}$ is defined as ${\rm diag}(3,...,3,1,...,1)$, where the numbers of $3$ and $1$ entries are both $p/2$. By structuring $\mathbf{Q}_{0}$ in this manner, we categorize stocks into two distinct volatility regimes, high and low, roughly capturing some weak knowledge of the covariance structure. In the simulation, we will vary the values of $q$ in order to better verify Theorem 2.7.
In the simulation, for each pair of $(n,p)$ considered, the parameters ${\bm{\Sigma}}_{0}\in\mathbb{R}^{p\times p}$, ${\bm{\mu}}_{0}\in\mathbb{R}^{p}$, and $\mathbf{Q}_{0}\in\mathbb{R}^{p\times p}$ are generated. Once they are generated, they remain fixed throughout the entire simulation procedure. Then we generate i.i.d. Gaussian random vector $\mathbf{R}_{i}$ ($i\in[n]$) with mean ${\bm{\mu}}={\bm{\mu}}_{0}$ and variance ${\bm{\Sigma}}={\bm{\Sigma}}_{0}$. The observed random matrix is $\mathbf{R}=(\mathbf{R}_{1},\mathbf{R}_{2},...,\mathbf{R}_{n})^{\top}\in\mathbb{R}^{n\times p}$. Here, we consider $(n,p)=(1500,750)$ and $(n,p)=(1500,2250)$. Note that since we assume the mean vector ${\bm{\mu}}$ is known, $\mathbf{X}=\mathbf{R}-{\bm{1}}_{n}{\bm{\mu}}_{0}^{\top}$ is also well observed. Recall that $\widehat{\bm{\Sigma}}=\mathbf{X}^{\top}\mathbf{X}/n$, and
$SR(\mathbf{Q})=\frac{{\bm{\mu}}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\mu}}}{\sqrt{{\bm{\mu}}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\mu}}}}.$
We compare the true $SR(\mathbf{Q})$ and the estimator $\widehat{SR}(\mathbf{Q})$ in Theorem 2.7 which is
$\widehat{SR}(\mathbf{Q})=\bigg{(}1-\frac{c}{p}{\mathrm{tr}}\widehat{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\bigg{)}\cdot\frac{{\bm{\mu}}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\mu}}}{\sqrt{{\bm{\mu}}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\widehat{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\mu}}}}.$
The Sharpe computation is then carried out as follows for each pair of $(n,p)$. We repeat the generation of Gaussian random matrices $\mathbf{R}\in\mathbb{R}^{n\times p}$ for $100$ independent times. For each repetition, we vary the value of $q$ in $\mathbf{Q}$ to compute $SR(q\cdot\mathbf{Q}_{0})$ and $\widehat{SR}(q\cdot\mathbf{Q}_{0})$. If $c<1$, we vary $q$ from (1:30)/5(1:30)/5( 1 : 30 ) / 5, while if $c>1$, we vary $q$ from (1:30)/1.5(1:30)/1.5( 1 : 30 ) / 1.5. Here, we assess the alignment between $SR(\mathbf{Q})$ and the proposed statistics $\widehat{SR}(\mathbf{Q})$ with various $q$ values. For each given $q$, we will obtain 100 different values of $SR(q\cdot\mathbf{Q}_{0})$ and $\widehat{SR}(q\cdot\mathbf{Q}_{0})$.
In Figure 1, the $x$-axis represents the different values of $q$, while the $y$-axis gives the Sharpe ratio. The simulation results are presented as the average of $SR(q\cdot\mathbf{Q}_{0})$ and $\widehat{SR}(q\cdot\mathbf{Q}_{0})$ over $100$ independent trials.
Figure 1: Simulation results in the basic settings. Figure 1(a) shows the case when $c=1/2$ and Figure 1(b) depicts the case when $c=3/2$. The x-axis corresponds to different $q$ values, and the y-axis is the value of $SR$. The black solid line connects the values of $SR(q\cdot\mathbf{Q}_{0})$, while the solid points represent the proposed statistics $\widehat{SR}(q\cdot\mathbf{Q}_{0})$ in Theorem 2.7.
It is evident that the value of $\widehat{SR}(q\cdot\mathbf{Q}_{0})$ matches the value of $SR(q\cdot\mathbf{Q}_{0})$ very well. As $q$ increases, $SR(q\cdot\mathbf{Q}_{0})$ initially rises and then falls, indicating the presence of an optimal $q$ that maximizes $SR$. With our proposed statistics in Theorem 2.7, we can replicate this trend without requiring knowledge of the true population matrix ${\bm{\Sigma}}$. Furthermore, the close match between $\widehat{SR}(q\cdot\mathbf{Q}_{0})$ and $SR(q\cdot\mathbf{Q}_{0})$ suggests that our estimator can be used to determine the best regularization matrix $\mathbf{Q}$ from a deterministic candidate set, such that $SR(\mathbf{Q})$ achieves the maximum value over the candidate set too.
4.1.1 Asymptotics with increasing sample size
In this section, we aim to investigate the effect of varying sample size $n$ and data dimension $p$ on our SR predictions, while keeping all other experimental settings consistent with previous settings. We continue to set the parameter $c=p/n$ at either $1/2$ or $3/2$, and focus on observing changes across different $n$ values, specifically $500,1000,1500$, and $2000$. For different values of $n$ and $p$, we still generate and then fix the population covariance matrix ${\bm{\Sigma}}_{0}\in\mathbb{R}^{p\times p}$, the mean vector ${\bm{\mu}}_{0}\in\mathbb{R}^{p}$ and the regularization matrix $\mathbf{Q}_{0}\in\mathbb{R}^{p\times p}$ in the same procedure as before.
Following the same procedure as described above in Section 4.1, for each $(n,p)$ and each given $q$, we obtain $100$ values of $SR(q\cdot\mathbf{Q}_{0})$ and $\widehat{SR}(q\cdot\mathbf{Q}_{0})$. For clear illustration, we let $SR_{b}(q\cdot\mathbf{Q}_{0};n,p)$ and $\widehat{SR}_{b}(q\cdot\mathbf{Q}_{0};n,p)$ denote the values obtained in the $b$-th independent trial ($b\in[100]$) under the pair $(n,p)$. The outcomes are subsequently illustrated in Figure 2.
Figure 2: Simulation results with increasing $n$. Figures 2(a)-2(d) and Figures 2(e)-2(h) respectively correspond to $c=p/n=1/2$ and $c=p/n=3/2$. In Figures 2(a) and 2(e), $y$-axis shows $\sum_{b=1}^{100}(SR_{b}(q\cdot\mathbf{Q}_{0};n,p)-\widehat{SR}_{b}(q\cdot\mathbf{Q}_{0};n,p))^{2}/100$ for different $q$’s. In Figures 2(b) and 2(f), $y$-axis shows $\sum_{b=1}^{100}(SR_{b}(q\cdot\mathbf{Q}_{0};n,p)/\widehat{SR}_{b}(q\cdot\mathbf{Q}_{0};n,p)-1)^{2}/100$ for different $q$’s. In Figures 2(c) and 2(g), $y$-axis gives the boxplot of $\operatorname{argmax}_{q}SR_{b}(q\cdot\mathbf{Q}_{0};n,p)-\operatorname{argmax}_{q}\widehat{SR}_{b}(q\cdot\mathbf{Q}_{0};n,p)$ across the 100 trials for different $n$’s. In Figures 2(d) and 2(h), $y$-axis displays the boxplot of $\max_{q}SR_{b}(q\cdot\mathbf{Q}_{0};n,p)-\max_{q}\widehat{SR}_{b}(q\cdot\mathbf{Q}_{0};n,p)$ across the 100 trials for different $n$’s.
Figure 2 shows that as we increase the sample size $n$, the mean-squared-error (MSE) of $SR(q\cdot\mathbf{Q}_{0};n,p)-\widehat{SR}(q\cdot\mathbf{Q}_{0};n,p)$ (difference) and $SR(q\cdot\mathbf{Q}_{0};n,p)/\widehat{SR}(q\cdot\mathbf{Q}_{0};n,p)-1$ (ratio) starts to get smaller for each $q$ value. So we verified our theory that when $n$ grows to infinity, the estimated value $\widehat{SR}(q\cdot\mathbf{Q}_{0};n,p)$ converges to the value $SR(q\cdot\mathbf{Q}_{0};n,p)$ almost surely, and they are ratio consistent. Additionally, we examine how the optimal $q$ values and the maximum Sharpe ratios of $SR(q\cdot\mathbf{Q}_{0};n,p)$ and $\widehat{SR}(q\cdot\mathbf{Q}_{0};n,p)$ can differ, and Figure 2 clearly demonstrates that as sample size $n$ grows, this discrepancy diminishes, lending further support to our theory.
4.2 Efficient frontiers
In this section, we conduct simulations to validate Theorem 3.2 for estimating the out-of-sample efficient frontiers. We still apply $(n,p)=(1500,750)$ with $c=1/2$ and $(n,p)=(1500,2250)$ with $c=3/2$. Recall the optimization problem in (3.2):
$\mathbf{w}^{*}=\arg\min_{\mathbf{w}}\mathbf{w}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})\mathbf{w},\quad\text{s.t. }\mathbf{w}^{\top}\mathbf{r}=\mu_{0}\text{ and }\mathbf{w}^{\top}{\bm{1}}=1.$ (4.1)
We will use the regularization matrix $\mathbf{Q}=0.2\,\mathbf{Q}_{0}$, where $\mathbf{Q}_{0}$ is defined in Section 4.1. The population covariance matrix ${\bm{\Sigma}}$ and the mean $\mathbf{r}={\bm{\mu}}$ (since the risk-free rate $r_{0}$ is assumed to be zero) we use here are slightly different from those used in Section 4.1. We first generate a vector ${\bm{\xi}}_{3}\in\mathbb{R}^{p}$ whose elements are independently distributed from $\Gamma(1,1)$. Then we let ${\bm{\Sigma}}={\bm{\Sigma}}_{3}={\rm diag}(\lambda_{1},\dots,\lambda_{p})+2\cdot{\bm{1}}{\bm{1}}^{\top}+{\bm{\xi}}_{3}{\bm{\xi}}_{3}^{\top}$. In addition, the mean vector ${\bm{\mu}}$ here has two choices: ${\bm{\mu}}_{3}={\bm{\mu}}_{0}+2\cdot{\bm{1}}$ and ${\bm{\mu}}_{4}={\bm{\mu}}_{0}+2\cdot{\bm{1}}+{\bm{\xi}}_{3}$, where ${\bm{\mu}}_{0}$ is again defined in Section 4.1. The matrix ${\bm{\Sigma}}_{3}$ simulates a covariance matrix with two factors. In the case of ${\bm{\mu}}={\bm{\mu}}_{3}={\bm{\mu}}_{0}+2\cdot{\bm{1}}$, besides betting on individual stock means ${\bm{\mu}}_{0}$, the mean vector also aligns with the market factor ${\bm{1}}$. In the other case of ${\bm{\mu}}={\bm{\mu}}_{4}={\bm{\mu}}_{0}+2\cdot{\bm{1}}+{\bm{\xi}}_{3}$, the mean vector additionally invests upon the second factor ${\bm{\xi}}_{3}$. Note that we do not need to consider the case $\mathbf{r}={\bm{\mu}}={\bm{\mu}}_{0}$ here because if the target return is $\mu_{0}$, the optimal portfolio can be equivalently obtained by setting $\mathbf{r}={\bm{\mu}}_{3}={\bm{\mu}}_{0}+2\cdot{\bm{1}}$ with the updated target return $\mu_{0}+2$, since $\mathbf{w}^{\top}{\bm{\mu}}_{3}=\mathbf{w}^{\top}{\bm{\mu}}_{0}+2(\mathbf{w}^{\top}{\bm{1}})=\mu_{0}+2$.
The following procedure is similar to Section 4.1. For each $(n,p)$, we (1) specify $\mathbf{Q}=0.2\,\mathbf{Q}_{0}$, ${\bm{\Sigma}}={\bm{\Sigma}}_{3}$, and ${\bm{\mu}}={\bm{\mu}}_{3}$ or ${\bm{\mu}}_{4}$; (2) generate 100 independent random matrices $\mathbf{R}\in\mathbb{R}^{n\times p}$ with i.i.d rows following $\mathcal{N}({\bm{\mu}},{\bm{\Sigma}})$; (3) for each given $\mathbf{R}$, compute the sample covariance matrix $\widehat{\bm{\Sigma}}=\frac{(\mathbf{R}-{\bm{1}}_{n}{\bm{\mu}}^{\top})^{\top}(\mathbf{R}-{\bm{1}}_{n}{\bm{\mu}}^{\top})}{n}$; (4) with each $\widehat{\bm{\Sigma}}$, solve the optimization problem in (4.1), for $\mu_{0}$ ranging from $0.2$ to $6$ with the increment of $0.2$; (5) for each $\mu_{0}$, calculate the corresponding values of $\sigma_{0}$ and $\widehat{\sigma}$, whose formula is provided in Theorem 3.2.
According to Theorem 3.2, the volatility of the optimal portfolio $\sigma_{0}$ is defined as $\sigma_{0}^{2}=\mathbf{w}^{*}{\bm{\Sigma}}\mathbf{w}^{*}$, and the corresponding estimator is $\widehat{\sigma}^{2}=\mathbf{w}^{*}\widehat{\bm{\Sigma}}\mathbf{w}^{*}/(1-c/p\cdot{\mathrm{tr}}\widehat{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1})^{2}$. We conduct simulations to confirm that $\widehat{\sigma}_{0}$ matches $\sigma_{0}$ well for each given $\mu_{0}$. The results displayed in Figure 3 are the average over the 100 values of $\sigma_{0}$ and $\widehat{\sigma}$ for different $\mu_{0}$.
Figure 3: Efficient frontiers of $\mathbf{w}^{*}$ from (4.1). The x-axis is the volatility level, and the y-axis is the target return $\mu_{0}$. The solid line characterizes the curve of $(\mu_{0},\sigma_{0})$, while the solid points represent the points in the curve of $(\mu_{0},\widehat{\sigma})$, where $\widehat{\sigma}$ is given in Theorem 3.2.
As shown in Figure 3, we observe that $\widehat{\sigma}$ estimates $\sigma_{0}$ very accurately for various $\mu_{0}$ and different $c$. As $\mu_{0}$ increases, we notice that the volatility $\sigma_{0}$ initially decreases and then increases, actually forming a hyperbola. Remarkably, our estimation $\widehat{\sigma}$ also closely depicts this curve, even without access to the population covariance matrix.
4.2.1 Asymptotics with increasing sample size
Similar to the previous Subsection 4.1.1, we investigate the effect of varying sample size $n$ and data dimension $p$ on volatility predictions on the efficient frontier. Again we set $c=1/2$ or $c=3/2$, with $n=500,1000,1500,2000$, ${\bm{\Sigma}}={\bm{\Sigma}}_{3}$, and ${\bm{\mu}}={\bm{\mu}}_{3}$ or ${\bm{\mu}}_{4}$. Following the same procedure as described above in Section 4.2, for each $(n,p)$ and each given $\mu_{0}$, we obtain $100$ values of $\sigma_{0}^{2}$ and $\widehat{\sigma}^{2}$, denoted as $\sigma_{0,b}^{2}$ and $\widehat{\sigma}^{2}_{b}$ for $b\in[100]$. The outcomes are subsequently illustrated in Figure 4.
Figure 4: Simulation results with increasing $n$. $x$-axis in all figures shows different values of $\mu_{0}$. In Figures 4(a)-4(d), $y$-axis shows $\sum_{b=1}^{100}(\widehat{\sigma}_{b}^{2}/\sigma_{0,b}^{2}-1)^{2}/100$. Similarly, in Figures 4(e)-4(h), $y$-axis shows $\sum_{b=1}^{100}(\widehat{\sigma}_{b}^{2}-\sigma_{0,b}^{2})^{2}/100$.
The MSE’s of both the difference, $\widehat{\sigma}^{2}-\sigma_{0}^{2}$, and the ratio, $\widehat{\sigma}^{2}/\sigma_{0}^{2}-1$, are illustrated in Figure 4 as we change the sample size $n$. Note that $\mathcal{A}_{rr}$ is bounded in this simulation. Therefore, the MSE gradually decreases for each $\mu_{0}$ as $n$ grows. Figures 4(c) and 4(g) exhibit some differences at $n=1500$ and $n=2000$, but the MSEs at these sample sizes are quite similar and considerably smaller than those at $n=500$ and $n=1000$. Another interesting observation is that as $\mu_{0}$ increases, the volatility absolute error also increases, while the volatility relative error stays roughly the same. Note that Theorem 3.2 states as long as $\mu_{0}$ is within a reasonable range, the absolute and relative error converges to zero. However, when $\mu_{0}$ is large, the volatility becomes large too in the efficient frontier, and the absolute difference between $\widehat{\sigma}^{2}$ and $\sigma^{2}$ may also be large, leading to the U-shape in Figures 4(e)-4(h).
5 Real Data Analysis
In this section, we conduct real data experiments to demonstrate the effectiveness and applicability of our methodology across various scenarios. The mean-variance portfolio in Section 5.1 and the efficient frontier in Section 5.2 are both constructed using daily stock returns from the constituents of S&P500. Since estimating the mean vector is difficult, we also consider the global minimum variance portfolio in Appendix B.1, which has the advantage of no requirement about the knowledge of ${\bm{\mu}}$. In addition, we consider calibrated models in Appendix B.2 to show the effect of estimating ${\bm{\mu}}$ in the mean-variance portfolio in a more straightforward way.
After deleting stocks with missing values, we have $p=365$ stocks in total. To robustly control extreme returns, we apply a truncation at 99% quantile. Portfolios are built using historical data spanning different time length, including one year, two years, and four years, and rebalanced monthly to reflect changing market dynamics. To illustrate the portfolio construction, we give an example of building up the portfolio to be held for January 2017 using one-year data. We utilize historical data from January 2016 to December 2016 to determine the optimized portfolio positions (i.e. the allocation vector $\mathbf{w}^{*}$). These positions are then held for the entire January 2017. Then we can calculate the excess return of the portfolio $\mathbf{w}^{*}$ for each trading day in January 2017. We refer to January 2017 as the current testing month. We repeat the procedure in a rolling fashion for all testing months spanning from January 2013 to June 2023 and record the daily returns for each trading day.
5.1 Mean variance portfolio
In this section, we investigate mean-variance portfolio construction in order to achieve high Sharpe ratio. Recall that mean-variance portfolio allocation vector is $\mathbf{w}\propto(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\mu}}$. We set its scaling by assuming $\sum_{i}|\mathbf{w}_{i}|=1$, meaning that the book of the portfolio is one unit. The weight vector $\mathbf{w}$ is thus expressed as:
$\mathbf{w}=\frac{(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\mu}}}{\|(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\mu}}\|_{1}}.$ (5.1)
In this experiment, we set $\mathbf{Q}=q\mathbf{I}$ and vary the value of $q$ from the candidate sets $\mathcal{Q}$. Based on Theorem 2.7, for each $q\in\mathcal{Q}$, we can utilize $\widehat{SR}$ to optimize the value of $q$ that yields a high $SR$. We present the steps below for our mean-variance portfolio analysis.
  1. 1.
    Consider two candidate sets $q\in\mathcal{Q}_{1}$ and $q\in\mathcal{Q}_{2}$: $\mathcal{Q}_{1}=[1:30]/5$ for $c>1$ and $\mathcal{Q}_{1}=[1:30]/1.5$ for $c<1$; $\mathcal{Q}_{2}=[1:30]/10$ considers a narrower range of values, resulting in smaller gaps thus more refined differences among values of $q$.
  2. 2.
    For each testing month, we observe the daily returns $\mathbf{R}\in\mathbb{R}^{n\times p}$, where $n$ is the total number of trading days with one-, two-, and four-year historical data length and $p=365$ is the number of selected stocks. Calculate the sample covariance matrix $\widehat{\bm{\Sigma}}$ and the portfolio weight $\mathbf{w}$ is given by (5.1).
  3. 3.
    For each testing month, we run experiments for all candidate $q$ values and also consider no regularization, i.e. $q=0$, where we have $\mathbf{w}\propto\widehat{\bm{\Sigma}}^{+}{\bm{\mu}}$ and $\widehat{\bm{\Sigma}}^{+}$ is the pseudo inverse, and the optimized $q^{*}\in\mathcal{Q}$ using the estimation in Theorem 2.7. Here,
    $q^{*}=\operatorname{argmax}_{q\in\mathcal{Q}}\bigg{(}1-\frac{c}{p}{\mathrm{tr}}\widehat{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+q\mathbf{I})^{-1}\bigg{)}\cdot\frac{{\bm{\mu}}^{\top}(\widehat{\bm{\Sigma}}+q\mathbf{I})^{-1}{\bm{\mu}}}{\sqrt{{\bm{\mu}}^{\top}(\widehat{\bm{\Sigma}}+q\mathbf{I})^{-1}\widehat{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+q\mathbf{I})^{-1}{\bm{\mu}}}}.$
  4. 4.
    We roll the procedure above for all testing months. Note that the value of $q^{*}$ changes from month to month. With the weight vector $\mathbf{w}$ using all $q\in\mathcal{Q}$, $q=0$ or $q^{*}$, we can then compute the portfolio returns for each trading day in the testing month.
  5. 5.
    We report the average Sharpe ratio of daily portfolio returns over the future three years.
Our target in the experiment is to verify the effectiveness of our methodology, instead of predicting ${\bm{\mu}}$ in this paper. So we use the mean from each testing month as ${\bm{\mu}}$ to construct portfolio.
Figure 5: SR of mean-variance portfolios. The x-axis labels the rolling period, while the y-axis represents the out-of-sample SR of the portfolio returns every three years in the future. The blue solid line, orange dash-dot line and green dash line correspond to the SR with $q=0$, $q=\min_{q\in\mathcal{Q}}q$ and $q=\max_{q\in\mathcal{Q}}q$ respectively. The boxplot displays the Sharpe ratio for all $q\in\mathcal{Q}$, and the red triangle indicates the SR under our optimized $q\in\mathcal{Q}$.
Figures 5(a)-5(c) present the results obtained with $q\in\mathcal{Q}_{1}$, while Figures 5(d)-5(f) present the results obtained with $q\in\mathcal{Q}_{2}$. Obviously, if we apply $q=0$ (the blue curve), we fail to attain the best possible SR, especially when sample size is low. If we always use the constant $q=\max_{q\in\mathcal{Q}}q$ (the orange curve), it is also a bad choice of regularization. In most of the time periods, $q=\min_{q\in\mathcal{Q}}q$ (the green curve) achieves a higher SR. However, more notably, the optimized $q^{*}\in\mathcal{Q}$ (the red triangles) consistently demonstrates superior performance throughout the experiments, regardless of the candidate sets and the varying lengths of historical data. From the results, we see clear advantage of actively optimized $q^{*}\in\mathcal{Q}$ for each testing month over some ad-hoc constant $q\in\mathcal{Q}$ for all testing months. The improvement indeed comes from our newly proposed out-of-sample SR estimator.
5.2 Efficient frontier
In this section, we test the application of Theorem 3.2 to efficient frontier evaluation of real data. For each specific $\mu_{0}$, we need to estimate the standard deviation of the regularized mean-variance portfolio. We consider the optimization problem presented in (3.2). Here are the steps of our real data experiment.
  1. 1.
    Set $\mathbf{Q}=q\mathbf{I}$ and let $\mathcal{Q}_{1}$ and $\mathcal{Q}_{2}$ the same as in Section 5.1. We vary the values of $\mu_{0}$ from $0.4$ to $6$ with an increment of $0.4$. For each $\mu_{0}$, we repeat Steps 2-5 below.
  2. 2.
    For each testing month, centralize the observed daily returns $\mathbf{R}\in\mathbb{R}^{n\times p}$, where $n$ corresponds to one-, two-, or four-year historical data and $p=365$. Then, calculate the sample covariance matrix $\widehat{\bm{\Sigma}}$ from the observed $\mathbf{R}$.
  3. 3.
    Let $\mathbf{r}$ be the average return vector in the testing month. The optimal portfolio is given by $\mathbf{w}^{*}=\mathbf{g}+\mu_{0}\mathbf{h}$, where $\mathbf{g}$ and $\mathbf{h}$ is given by (3.5), with $\mathbf{Q}$ replaced by $q\mathbf{I}$. We run experiments for all $q$ values in the candidate sets, including the case of $q=0$ and the optimized $q^{*}$ value. By Theorem 3.2, the optimized $q^{*}$ value is obtained by minimizing $\frac{(\mathbf{g}+\mu_{0}\mathbf{h})^{\top}\widehat{\bm{\Sigma}}(\mathbf{g}+\mu_{0}\mathbf{h})}{(1-c/p\cdot{\mathrm{tr}}\widehat{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+q\mathbf{I})^{-1})^{2}}$ over all $q\in\mathcal{Q}$. The case $q=0$ replaces $(\widehat{\bm{\Sigma}}+q\mathbf{I})^{-1}$ with $\widehat{\bm{\Sigma}}^{+}$ in $\mathbf{w}^{*}$.
  4. 4.
    We repeat Step 2 for each testing month from January 2013 to June 2023. This allows us to collect daily portfolio returns for each $q$ value.
  5. 5.
    We calculate the standard deviation of the daily returns for each $q$ value, including $q=0$ and $q=q^{*}$, over the ten-year period. This allows us to generate a boxplot of the standard deviation for each candidate $q$ value in $\mathcal{Q}_{1}$ or $\mathcal{Q}_{2}$ given $\mu_{0}$. We also obtain the standard deviation of the portfolios corresponding to $q=0$ and $q=q^{*}$.
Figure 6: The corrected efficient frontier. The x-axis represents the value of $\sigma$, while the y-axis represents different values of $\mu_{0}$. The blue solid line, orange dash-dot line and green dash line in the figure correspond to the different values of $\sigma$ with $q=0$, $q=\min_{q\in\mathcal{Q}}q$ and $q=\max_{q\in\mathcal{Q}}q$ respectively. The boxplot displays the values of $\sigma$ for all $q\in\mathcal{Q}$, and the red triangle indicates the volatility of the portfolio returns under our optimized $q\in\mathcal{Q}$.
As shown in Figure 6, the standard approach with no regularization ($q=0$) substitutes $(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}$ with $\widehat{\bm{\Sigma}}^{+}$ and exhibits a significantly larger volatility in comparison to the regularized portfolios for any given $\mu_{0}$. Furthermore, as illustrated in Figures 6(d) and 6(f), the minimum standard deviation occurs when $q$ is approximately $\max_{q\in\mathcal{Q}_{2}}q$ and $\min_{q\in\mathcal{Q}_{2}}q$ respectively. However, our corrected estimator consistently identify the optimal $q\in\mathcal{Q}_{2}$ and result in a smaller standard deviation. This again demonstrates the practical usefulness of Theorem 3.2 in real efficient frontier recovery of the regularized portfolios.
6 Conclusion
In this work, we introduce a novel in-sample approach to estimate the out-of-sample Sharpe ratio in high-dimensional portfolio optimization, both with and without the risk-free asset, when the population covariance matrix ${\bm{\Sigma}}$ is unknown. By leveraging random matrix theory, we show that the proposed estimator shares the same asymptotic properties as the true Sharpe ratio, even when the operator norm of ${\bm{\Sigma}}$ is unbounded. Specifically, we only need the trace norm of ${\bm{\Sigma}}/p$ to be bounded, which makes our methodology more applicable to the real world financial data. We have also validated our theory through simulations and demonstrated its practical application through real data experiments.
In future investigations, it would be interesting to explore the convergence rate of the proposed estimator and determine if there are alternative estimators that offer faster convergence. This would involve studying the rate at which our estimator approaches the true Sharpe ratio as the sample size increases, and comparing it to other potential estimators. Understanding the convergence properties of different estimators can help identify optimal approaches for estimating the Sharpe ratio in different scenarios.
Supplementary Material for “Estimation of Out-of-Sample Sharpe Ratio for High Dimensional Portfolio Optimization”
Appendix A Alternative Settings for Numeric Experiments
In this section, we change the values of the population covariance matrix ${\bm{\Sigma}}$, the regularization matrix $\mathbf{Q}$ and the mean vector ${\bm{\mu}}$ separately to assess the accuracy of Theorem 2.7, while keeping the values of the other quantities the same as those in Section 4.1. This provides comparison as we evaluate the effects of ${\bm{\Sigma}}$, $\mathbf{Q}$ and ${\bm{\mu}}$ separately. We investigate these changes for $(n,p)=(1500,750)$ and $(n,p)=(1500,2250)$ as in the basic settings. We still generate $\mathbf{R}\in\mathbb{R}^{n\times p}$ for 100 times and follow the procedure in Section 4.1 to get $SR(\mathbf{Q})$ and $\widehat{SR}(\mathbf{Q})$.
Results with different ${\bm{\Sigma}}$’s. Here we fix ${\bm{\mu}}={\bm{\mu}}_{0}$ and $\mathbf{Q}=q\cdot\mathbf{Q}_{0}$. To validate our high dimensional statistical correction, we select two other covariance structures, denoted as ${\bm{\Sigma}}_{1}$ and ${\bm{\Sigma}}_{2}$. We construct ${\bm{\Sigma}}_{1}$ as the diagonal matrix ${\bm{\Sigma}}_{1}={\rm diag}(\lambda_{1},\dots,\lambda_{p})$, where $\{\lambda_{i}\}_{i=1}^{p}$ is the sequence described in Section 4.1. This choice of ${\bm{\Sigma}}_{1}$ represents the situation where we have no factor structure but well-conditioned spectrum. We construct ${\bm{\Sigma}}_{2}$ as ${\bm{\Sigma}}_{2}={\bm{\Sigma}}_{0}+{\bm{\xi}}_{1}{\bm{\xi}}_{1}^{\top}+{\bm{\xi}}_{2}{\bm{\xi}}_{2}^{\top}$, where ${\bm{\Sigma}}_{0}$ is defined in Section 4.1, and ${\bm{\xi}}_{1}$ and ${\bm{\xi}}_{2}$ are two random generated vectors that are orthogonal to each other and also to the vector ${\bm{1}}$. The norm of ${\bm{\xi}}_{1}$ and ${\bm{\xi}}_{2}$ is chosen such that $\|{\bm{\xi}}_{1}\|_{2}^{2}=\|{\bm{\xi}}_{2}\|_{2}^{2}=p$. This choice of ${\bm{\Sigma}}_{2}$ allows more factors besides the market factor, e.g. the Size and Value factors in Fama-French 3 factor model (Fama & French 1993). From a mathematical perspective, the assumptions on ${\bm{\Sigma}}_{1}$ and ${\bm{\Sigma}}_{2}$ cover a broader range of cases to verify our theory. The simulation procedure is similar to Section 4.1, except that we now replace the population covariance ${\bm{\Sigma}}_{0}$ with ${\bm{\Sigma}}={\bm{\Sigma}}_{1}$ or ${\bm{\Sigma}}={\bm{\Sigma}}_{2}$.
Figure 7: Simulation results with different ${\bm{\Sigma}}$’s. The x-axis and the y-axis follow exactly Figure 1.
The results presented in Figure 7 show that our estimated $\widehat{SR}(\mathbf{Q})$, obtained without access to the population covariance matrix, again closely aligns with the true $SR(\mathbf{Q})$ for both ${\bm{\Sigma}}_{1}$ and ${\bm{\Sigma}}_{2}$ in both the cases of $c=1/2$ and $c=3/2$. This observation demonstrates that our high dimensional in-sample correction effectively captures the out-of-sample Sharpe ratio in all 3 cases in Assumption 2.6, no matter if we have bounded eigenvalues or a few diverging spikes.
Results with different $\mathbf{Q}$’s. Here we fix ${\bm{\mu}}={\bm{\mu}}_{0}$ and ${\bm{\Sigma}}={\bm{\Sigma}}_{0}$. In the basic settings, we define $\mathbf{Q}=q\cdot\mathbf{Q}_{0}=q\cdot{\rm diag}(3,...,3,1,...,1)$, where the number of $3$’s and $1$’s are both $p/2$. To provide a comparison, we introduce two additional matrices, $\mathbf{Q}_{1}=0.5\mathbf{I}_{p}+q\cdot\mathbf{Q}_{0}$ and $\mathbf{Q}_{2}=2\mathbf{I}_{p}+q\cdot\mathbf{Q}_{0}$, when $(n,p)=(1500,750)$ thus $c=1/2$. We can treat $\mathbf{Q}_{1},\mathbf{Q}_{2}$ as selection of another parameter $\widetilde{q}\in\{0,0.5,2\}$ in $\mathbf{Q}=\widetilde{q}\mathbf{I}_{p}+q\mathbf{Q}_{0}$. Similarly, we consider $\mathbf{Q}_{1}=2\mathbf{I}_{p}+q\cdot\mathbf{Q}_{0}$ and $\mathbf{Q}_{2}=4\mathbf{I}_{p}+q\cdot\mathbf{Q}_{0}$ when $(n,p)=(1500,2250)$ thus $c=3/2$. The inclusion of the term $\mathbf{I}_{p}$ with different coefficients serves as a universal variance penalty, which makes the covariance matrix more homogeneous across diagonal entries. The simulation procedure is similar to Section 4.1, except that we now replace the regularization $q\cdot\mathbf{Q}_{0}$ with $\mathbf{Q}=\mathbf{Q}_{1}$ or $\mathbf{Q}=\mathbf{Q}_{2}$.
Figure 8: Simulation results with different $\mathbf{Q}$’s. The x-axis and the y-axis follow exactly Figure 1.
In Figure 8, it can be observed that $\widehat{SR}(\mathbf{Q})$ matches with $SR(\mathbf{Q})$ for both $\mathbf{Q}=\mathbf{Q}_{1}$ and $\mathbf{Q}=\mathbf{Q}_{2}$ when $c=1/2$ and $c=3/2$. However, with different structures of $\mathbf{Q}$, the maximum value of SR can vary significantly. For instance, when $c=1/2$, the maximum value of SR in the basic settings is approximately $0.94$. However, in Figures 8(a) and 8(b), the maximum SR exceeds $1$. On the other hand, when $c=3/2$, the maximum value of SR in the basic settings is more than $0.97$. But in Figures 8(c) and 8(d), the maximum value of SR only approaches $0.92$. These results demonstrate the significance of determining the structure of the matrix $\mathbf{Q}$ in order to achieve a higher SR.
Results with different $\bm{\mu}$’s. Here we fix ${\bm{\Sigma}}={\bm{\Sigma}}_{0}$ and $\mathbf{Q}=q\cdot\mathbf{Q}_{0}$ and introduce different settings for the mean vectors, ${\bm{\mu}}_{1}$ and ${\bm{\mu}}_{2}$. For ${\bm{\mu}}_{1}$, we assume that each element follows an independent uniform distribution, ${\mathop{\mathrm{Unif}}}(-\sqrt{2/p},\sqrt{2/p})$. This assumption ensures that each individual asset has a nonzero expected return. For ${\bm{\mu}}_{2}$, we add the vector $2\cdot\text{\bf 1}_{p}$ to ${\bm{\mu}}_{1}$, that is ${\bm{\mu}}_{2}={\bm{\mu}}_{1}+2\cdot\text{\bf 1}_{p}$. Recall that the vector ${\bm{1}}$ represents the market factor, so ${\bm{\mu}}_{2}$ induces a portfolio that not only bets on “alpha” signals but also holds a long position in the market factor. The simulation procedure is similar to Section 4.1, except that we now replace the mean vector ${\bm{\mu}}_{0}$ with ${\bm{\mu}}={\bm{\mu}}_{1}$ or ${\bm{\mu}}={\bm{\mu}}_{2}$.
Figure 9: Simulation results with different ${\bm{\mu}}$’s. The x-axis and the y-axis follow exactly Figure 1.
The results presented in Figure 9 demonstrate that our estimated $\widehat{SR}(\mathbf{Q})$ closely aligns with the true $SR(\mathbf{Q})$ for both ${\bm{\mu}}_{1}$ and ${\bm{\mu}}_{2}$. This comparison across different values of ${\bm{\mu}}$ further highlights the robustness and accuracy of our approach in estimating $SR(\mathbf{Q})$. Furthermore, when comparing the cases of ${\bm{\mu}}={\bm{\mu}}_{1}$ and ${\bm{\mu}}={\bm{\mu}}_{2}$, we see a significant difference in $SR(\mathbf{Q})$. This observation indicates that for different mean returns, Sharpe ratio can vary dramatically, but our method can accurately evaluate the out-of-sample Sharpe ratio regardless of the structure of the mean returns.
Appendix B Additional Real Data Experiments
B.1 Global minimum variance portfolio
In this section, we consider the global minimum variance portfolio, which has the advantage of no requirement about the knowledge of ${\bm{\mu}}$. Our construction of the global minimum variance portfolio is defined as $\mathbf{w}\propto(\widehat{\bm{\Sigma}}+q\mathbf{I})^{-1}{\bm{1}}$, where we simply apply the regularization $\mathbf{Q}=q\mathbf{I}$. Since the total weights sum up to 1, the exact weight vector $\mathbf{w}$ takes the form:
$\mathbf{w}=\frac{(\widehat{\bm{\Sigma}}+q\mathbf{I})^{-1}{\bm{1}}}{{\bm{1}}^{\top}(\widehat{\bm{\Sigma}}+q\mathbf{I})^{-1}{\bm{1}}}.$ (B.1)
Note that the scale is determined slightly differently from the mean variance portfolio, for the reason that ${\bm{1}}^{\top}(\widehat{\bm{\Sigma}}+q\mathbf{I})^{-1}{\bm{1}}$ is always positive in the global minimum variance portfolio, but ${\bm{1}}^{\top}(\widehat{\bm{\Sigma}}+q\mathbf{I})^{-1}{\bm{\mu}}$ may become negative and unstable in the mean-variance portfolio. Consequently, the variance of the portfolio can be expressed as:
$\operatorname{{\rm Var}}_{\widetilde{\mathbf{R}}}(\mathbf{w}^{\top}\widetilde{\mathbf{R}})=\frac{{\bm{1}}^{\top}(\widehat{\bm{\Sigma}}+q\mathbf{I})^{-1}{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+q\mathbf{I})^{-1}{\bm{1}}}{({\bm{1}}^{\top}(\widehat{\bm{\Sigma}}+q\mathbf{I})^{-1}{\bm{1}})^{2}},$
where $\widetilde{\mathbf{R}}$ is an out-of-sample return vector with the distribution $\mathcal{N}({\bm{\mu}},{\bm{\Sigma}})$. Theorem 2.7 provides an in-sample method to estimate
$1/\sqrt{\operatorname{{\rm Var}}_{\widetilde{\mathbf{R}}}(\mathbf{w}^{\top}\mathbf{x})}=\frac{{\bm{1}}^{\top}(\widehat{\bm{\Sigma}}+q\mathbf{I})^{-1}{\bm{1}}}{\sqrt{{\bm{1}}^{\top}(\widehat{\bm{\Sigma}}+q\mathbf{I})^{-1}{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+q\mathbf{I})^{-1}{\bm{1}}}},$
which exactly has the form of (2.1), replacing ${\bm{\mu}}$ with ${\bm{1}}$. To minimize the out-of-sample $\operatorname{{\rm Var}}_{\widetilde{\mathbf{R}}}(\mathbf{w}^{\top}\mathbf{x})$, Theorem 2.7 can help us determine the best $q$ from a finite candidate set. Below we give the steps of our real data analysis for the global minimum portfolio.
  1. 1.
    $\mathbf{Q}=q\mathbf{I}$. Consider two candidate sets $\mathcal{Q}_{1}$ and $\mathcal{Q}_{2}$: $\mathcal{Q}_{1}=[1:30]/5$ for $c<1$ and $\mathcal{Q}_{1}=[1:30]/1.5$ for $c>1$; $\mathcal{Q}_{2}=[1:30]/10$ considers a narrower range of values, resulting in smaller gaps thus more refined differences among values of $q$.
  2. 2.
    For each testing month, we observe the daily returns $\mathbf{R}\in\mathbb{R}^{n\times p}$, where $n$ is the total number of trading days with one-, two-, and four-year historical data length and $p=365$ is the number of selected stocks. Calculate the sample covariance matrix $\widehat{\bm{\Sigma}}$ and the portfolio weight $\mathbf{w}$ is given by (B.1).
  3. 3.
    For each testing month, we run experiments for all candidate $q$ values and also consider no regularization, i.e. $q=0$, where we have $\mathbf{w}\propto\widehat{\bm{\Sigma}}^{+}{\bm{1}}$ and $\widehat{\bm{\Sigma}}^{+}$ is the pseudo inverse, and the optimized $q^{*}\in\mathcal{Q}$ using the estimation in Theorem 2.7. Here,
    $q^{*}=\operatorname{argmax}_{q\in\mathcal{Q}}\bigg{(}1-\frac{c}{p}{\mathrm{tr}}\widehat{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+q\mathbf{I})^{-1}\bigg{)}\cdot\frac{{\bm{1}}^{\top}(\widehat{\bm{\Sigma}}+q\mathbf{I})^{-1}{\bm{1}}}{\sqrt{{\bm{1}}^{\top}(\widehat{\bm{\Sigma}}+q\mathbf{I})^{-1}\widehat{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+q\mathbf{I})^{-1}{\bm{1}}}}.$
  4. 4.
    We roll the procedure above for all testing months. Note that the value of $q^{*}$ changes from month to month. With the weight vector $\mathbf{w}$ using all $q\in\mathcal{Q}$, $q=0$ or $q^{*}$, we can then compute the portfolio returns for each trading day in the testing month.
  5. 5.
    We report the standard deviation of the portfolio returns every three years, as shown in Figure 10. For instance, the results labeled on the x-axis as 202001 in Figure 10 represent the standard deviation of out-of-sample portfolio returns from January 2020 to December 2022 with 36 testing months. When the x-axis label updates to 202007, the results reflect the standard deviation of returns from July 2020 to June 2023.
Figure 10: Standard deviation (volatility) of global minimum variance portfolios. The x-axis represents the rolling time, while the y-axis represents the standard deviation of the portfolio returns every three years. The blue solid line, orange dash-dot line and green dash line in the figure correspond to the standard deviation with $q=0$, $q=\min_{q\in\mathcal{Q}}q$ and $q=\max_{q\in\mathcal{Q}}q$ respectively. The boxplot displays the standard deviation for all $q\in\mathcal{Q}$, and the red triangle indicates the standard deviation under optimized $q=q^{*}\in\mathcal{Q}$.
Figure 10(a)-10(c) present the results obtained using historical data of varying lengths with $q\in\mathcal{Q}_{1}$, and Figure 10(d)-10(f) use $q\in\mathcal{Q}_{2}$. From Figure 10, we can make a few conclusions. Firstly, with no regularization ($q=0$, the blue curve), the out-of-sample volatility is very high. So adding proper regularization to protect the condition of the sample covariance is absolutely necessary. Secondly, as shown in Figure 10, for each of the 6 cases, if we fix one choice $q\in\mathcal{Q}$ for all testing months, for most months $q=\min_{q\in\mathcal{Q}}q$ (the orange curve) achieves the smallest out-of-sample volatility, but occasionally $q=\max_{q\in\mathcal{Q}}q$ (the green curve) outperforms $q=\min_{q\in\mathcal{Q}}q$. If the volatility happens to be monotone in the range of $q$ values in $\mathcal{Q}$, the boxplot will be in the middle between the orange and green curves, but if the best $q$ is neither the minimum $q$ nor the maximum $q$, we also see cases that the boxplot is beneath both the orange and green curves. This indicates that applying a constant level of regularization for all months such as $q=\min_{q\in\mathcal{Q}}q$ to the portfolio may not be suitable for all time periods. So if we want to update $q$ values from month to month, it is important to have a in-sample method to estimate the out-of-sample volatility. Thirdly, from all 6 cases in Figure 10, we see that the portfolios with optimized $q^{*}$ (the red triangles), automatically updated every month, consistently exhibit a smaller standard deviation regardless of the candidate sets and the varying lengths of historical data. So we can conclude that our method can consistently and flexibly select an appropriate $q\in\mathcal{Q}$ based on the data itself, leading to smaller out-of-sample variances of the global minimum variance portfolio and demonstrating significant advantages over naively choosing a fixed $q\in\mathcal{Q}$ for all time points. The last side observation is that the volatility level jumps to a higher level starting from the label 201707, which actually represents July 2017 to June 2020, the first time period covering the COVID outbreak.
B.2 Calibrated Models with Real Data
In the previous experiments for mean-variance portfolios, we made the assumption that the mean vector ${\bm{\mu}}$ in S&P500 is known. However, in reality, it is typically unknown. In this section, we hope to look into how our methods are affected when ${\bm{\mu}}$ is unobserved. To this end, we use real data to calibrate a model for ${\bm{\mu}}$ and then use the calibrated model to simulate synthetic data. Here are the steps of this experiment.
  1. 1.
    Using solely the returns in the testing month $j$, we have the out-of-sample average return vector ${\bm{\mu}}_{j}$ and the out-of-sample estimated covariance matrix $\widetilde{\bm{\Sigma}}_{j}=\widetilde{\mathbf{X}}_{j}^{\top}\widetilde{\mathbf{X}}_{j}/\widetilde{n}_{j}$. Here, $\widetilde{\mathbf{X}}_{j}$ is the return matrix for the testing month $j$, and $\widetilde{n}_{j}$ is the number of trading days in that month. Additionally, we define the in-sample empirical covariance matrix $\widehat{\bm{\Sigma}}_{j}=\mathbf{X}_{j}^{\top}\mathbf{X}_{j}/n_{j}$, where $\mathbf{X}_{j}$ is the return matrix consisting of five years prior to the testing month $j$, and $n_{j}$ is the total number of trading days in those five years.
  2. 2.
    Using the well-defined ${\bm{\mu}}_{j}$, $\widetilde{\bm{\Sigma}}_{j}$, and $\widehat{\bm{\Sigma}}_{j}$, we generate returns of $21$ trading days for each trading month. The returns for the trading days in the month $j$ are drawn from the distribution $\mathcal{N}(\check{{\bm{\mu}}}_{j},\check{{\bm{\Sigma}}}_{j})$, where:
    $\begin{split}&\check{{\bm{\mu}}}_{j}=\rho\cdot{\bm{\mu}}_{j}+\sqrt{1-\rho^{2}}\cdot\mathcal{N}(0,\sigma^{2}_{{\bm{\mu}}_{j}}\mathbf{I}),\\ &\check{{\bm{\Sigma}}}_{j}=0.95\cdot\widehat{\bm{\Sigma}}_{j}+\sqrt{1-0.95^{2}}\cdot\widetilde{\bm{\Sigma}}_{j}.\end{split}$ (B.2)
    Here, $\rho>0$ is a tuning parameter that controls the level of perturbation to ${\bm{\mu}}_{j}$, and $\sigma_{{\bm{\mu}}_{j}}$ is the standard deviation of the elements in ${\bm{\mu}}_{j}$.
    Figure 11: Sharpe ratio of mean-variance portfolio with calibrated models. The three rows of plots correspond to $\rho=0.05,0.2,0.4$ respectively. The x-axis represents the rolling time, while the y-axis represents the Sharpe ratio of portfolio returns every three years.
  3. 3.
    To use the calibrated model, we assume that the mean vector ${\bm{\mu}}_{j}$ is known and used in portfolio construction, but the actual true mean of the returns for the month $j$ is $\check{{\bm{\mu}}}_{j}$ which is unknown. We use different values of $\rho$ to generate different calibrated models, where a larger $\rho$ reflects a higher signal level of predicting $\check{{\bm{\mu}}}_{j}$ using ${\bm{\mu}}_{j}$. The allocation vector for the testing month $j$ is then constructed as $\mathbf{w}\propto(\widehat{\bm{\Sigma}}+q\mathbf{I})^{-1}{\bm{\mu}}_{j}$, and the remaining steps follow the same procedure in Section 5.1.
We present experimental results for $\rho=0.05$, $0.2$, and $0.4$ in three rows of plots respectively in Figure 11. As indicated in equation (B.2), smaller values of $\rho$ result in larger perturbations on the mean vector ${\bm{\mu}}_{j}$. $\rho=0.05$, $0.2$, and $0.4$, the mean cosine values of the angle between ${\bm{\mu}}_{j}$ and $\check{{\bm{\mu}}}_{j}$ are $0.0428$, $0.1368$, and $0.2605$, respectively, with corresponding standard deviations of $0.2676$, $0.2649$, and $0.2562$. We observed that as $\rho$ increases, the Sharpe ratio for the portfolio gradually increases. This makes intuitive sense as higher prediction power leads to better portfolio out-of-sample performance. Specifically, for $\rho=0.2$ and $0.4$, the optimized $q^{*}$ based on our in-sample SR estimator always displays the best performance among the candidate $q$ values in $\mathcal{Q}=\mathcal{Q}_{1}$ or $\mathcal{Q}=\mathcal{Q}_{2}$. For $\rho=0.05$, we may not always achieve the best out-of-sample SR due to the low signal level ($\check{{\bm{\mu}}}_{j}$ becomes nearly orthogonal to ${\bm{\mu}}_{j}$). However, we still see comparable performances.
Appendix C Supporting Lemmas
In this section, we will present several lemmas that play a crucial role in determining the asymptotic behavior of $T_{n,1}(\mathbf{Q})$ and $T_{n,2}(\mathbf{Q})$. These lemmas are instrumental in our analysis and provide valuable insights into the problem at hand.
Lemma C.1 (Theorem 1 in Rubio & Mestre (2011)).
Given $\mathbf{X}\in\mathbb{R}^{n\times p}$ with $\mathbf{X}=\mathbf{Z}{\bm{\Sigma}}^{\frac{1}{2}}$ where the elements of $\mathbf{Z}\in\mathbb{R}^{n\times p}$ are i.i.d with zero mean, variance 1 and finite $8+\varepsilon$ order moment for some $\varepsilon>0$, ${\bm{\Sigma}}^{\frac{1}{2}}\in\mathbb{R}^{p\times p}\geq 0$ is a non negative Hermitian matrix with $\|{\bm{\Sigma}}\|_{{\mathop{\mathrm{op}}}}$ bounded. For any deterministic $\mathbf{A}\in\mathbb{R}^{p\times p}$ with bounded trace norm (${\mathrm{tr}}(\sqrt{\mathbf{A}\mathbf{A}^{\top}})\leq C$ with respect to $n$) and Hermitian positive matrix $\mathbf{S}\in\mathbb{R}^{p\times p}$, define
$m_{n}(z)={\mathrm{tr}}\bigg{(}\frac{\mathbf{X}^{\top}\mathbf{X}}{n}+\mathbf{S}+z\mathbf{I}_{p}\bigg{)}^{-1}\mathbf{A},$
and assume that $p,n$ tends to infinity with proportion $p/n\to c>0$. Then for $z\in\mathbb{C}-(-\infty,0]$,
$m_{n}(z)-{\mathrm{tr}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}(z)}+\mathbf{S}+z\mathbf{I}_{p}\bigg{)}^{-1}\mathbf{A}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0.$ (C.1)
Here, $s_{0}(z)$ is the Stieltjes transform of a positive finite measure and uniquely solves the following equation:
$s_{0}(z)=\frac{c}{p}{\mathrm{tr}}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}(z)}+\mathbf{S}+z\mathbf{I}_{p}\bigg{)}^{-1}.$
Note that Corollary 1.11 in Widder (1938) gives that the convergence of such Stieltjes transform in (C.1) is indeed internally closed uniform convergence, combined with the fact that the functions above are all analytic, we have the following corollary:
Corollary C.2.
In the condition of Lemma C.1, it still holds the convergence that
$\frac{{\mathop{\mathrm{d}}}m_{n}(z)}{{\mathop{\mathrm{d}}}z}-\frac{{\mathop{\mathrm{d}}}~{}{\mathrm{tr}}\Big{(}\frac{{\bm{\Sigma}}}{1+s_{0}(z)}+\mathbf{S}+z\mathbf{I}_{p}\Big{)}^{-1}\mathbf{A}}{{\mathop{\mathrm{d}}}z}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0.$ (C.2)
Lemma C.1 and Corollary C.2 are limited to the constraint $\|{\bm{\Sigma}}\|_{{\mathop{\mathrm{op}}}}$ is bounded. We first extend the results above to the case where $\|{\bm{\Sigma}}\|_{{\mathop{\mathrm{op}}}}$ is unbounded. We give the following definition first.
Definition C.3.
Define $s_{0}(z)$ by the equation
$s_{0}(z)=\frac{c}{p}{\mathrm{tr}}~{}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}(z)}+\mathbf{Q}+z{\bm{\Sigma}}\bigg{)}^{-1}.$
Then let
$f_{n}(z)={\mathrm{tr}}\bigg{(}\frac{\mathbf{Z}^{\top}\mathbf{Z}}{n}+{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{Q}{\bm{\Sigma}}^{-\frac{1}{2}}+z\mathbf{I}\bigg{)}^{-1}{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{A}{\bm{\Sigma}}^{-\frac{1}{2}},$
$\overline{f}_{n}(z)={\mathrm{tr}}\bigg{(}\frac{\mathbf{I}}{1+s_{0}}+{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{Q}{\bm{\Sigma}}^{-\frac{1}{2}}+z\mathbf{I}\bigg{)}^{-1}{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{A}{\bm{\Sigma}}^{-\frac{1}{2}}.$
The existence and uniqueness of $s_{0}(z)$ can be found in Rubio & Mestre (2011). By Definition C.3 above, we have the following lemma.
Lemma C.4.
Suppose that Assumption 2.4, 2.5 and 2.6 hold, let $f_{n}(z)$ and $\overline{f}_{n}(z)$ be defined in Definition C.3 with $\|\mathbf{A}\|_{{\mathrm{tr}}}\leq C<+\infty$, then for any fixed $\0z\in\mathbb{C}\backslash(-\infty,0]$, it holds that
$f_{n}(z)-\overline{f}_{n}(z)~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0,\quad f^{\prime}_{n}(z)-\overline{f}^{\prime}_{n}(z)~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0.$
Proof of Lemma C.4.
Lemma C.1 and Corollary C.2 directly supports Lemma C.4. ∎
From Lemma C.4 above, we directly conclude that for fixed $\0z\in\mathbb{C}\backslash(-\infty,0]$,
${\mathrm{tr}}\bigg{(}\frac{\mathbf{X}^{\top}\mathbf{X}}{n}+\mathbf{Q}+z{\bm{\Sigma}}\bigg{)}^{-1}\mathbf{A}-{\mathrm{tr}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}(z)}+\mathbf{Q}+z{\bm{\Sigma}}\bigg{)}^{-1}\mathbf{A}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0,$
and the convergence also holds for the derivatives of $z$. However, this conclusion is not sufficient for us to prove Theorem 2.7, as it is crucial for us to consider the case when $z=0$. For the Case 1 in Assumption 2.6, by $\lambda_{\min}(\mathbf{Q})\geq c_{\mathbf{Q}}$, we can easily let the matrix $\mathbf{S}$ in Lemma C.1 by $\mathbf{S}=\mathbf{Q}-c_{\mathbf{Q}}/2\mathbf{I}$, and $z=c_{\mathbf{Q}}/2\mathbf{I}$, then we can extend the convergence to $z=0$ when the Case 1 in Assumption 2.6 holds. In the following lemmas (Lemma C.5-C.7), we aim to extend the convergence to $z=0$ in the case 2 and 3 in Assumption 2.6. To prove so, we limit the range of $z$ into $[0,+\infty)$. We have the following lemma.
Lemma C.5.
Suppose that Assumption 2.4, 2.5 and 2.6 hold, when $z\in[0,+\infty)$, $s_{0}(z)$ defined in Definition C.3 exists and is unique. It holds that $0<s_{0}(z)\leq{c\|{\bm{\Sigma}}/p\|_{{\mathrm{tr}}}}/{\lambda_{\min}(\mathbf{Q})}.$ For $z\in[0,+\infty)$, we further define $s_{1,{\bm{\Sigma}}}(z)$ and $s_{1,\mathbf{Q}}(z)$ by
$s_{1,{\bm{\Sigma}}}(z)=\bigg{(}\frac{s_{1,{\bm{\Sigma}}}(z)}{(1+s_{0}(z))^{2}}-1\bigg{)}\frac{c}{p}{\mathrm{tr}}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}(z)}+\mathbf{Q}+z{\bm{\Sigma}}\bigg{)}^{-1}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}(z)}+\mathbf{Q}+z{\bm{\Sigma}}\bigg{)}^{-1},$
$s_{1,\mathbf{Q}}(z)=\frac{c}{p}{\mathrm{tr}}~{}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}(z)}+\mathbf{Q}+z{\bm{\Sigma}}\bigg{)}^{-1}\bigg{(}\frac{s_{1,\mathbf{Q}}(z)}{(1+s_{0}(z))^{2}}{\bm{\Sigma}}-\mathbf{Q}\bigg{)}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}(z)}+\mathbf{Q}+z{\bm{\Sigma}}\bigg{)}^{-1},$
then $s_{1,{\bm{\Sigma}}}(z)$ and $s_{1,\mathbf{Q}}(z)$ exist, and it holds that
$-s_{0}(z)(1+s_{0}(z))^{2}\leq s_{1,{\bm{\Sigma}}}(z)<0,\quad-s_{0}(z)\leq s_{1,\mathbf{Q}}(z)<0.$
Lemma C.5 shows that the values of $s_{0}(z)$, $s_{1,{\bm{\Sigma}}}(z)$ and $s_{1,\mathbf{Q}}(z)$ range in a constant level. The proof can be found in Section C.1. With Lemma C.5, we give the following lemmas which play a key role in the analysis under Case 2 and 3.
Lemma C.6.
Suppose that Assumption 2.4, 2.5 and 2.6 Case 2 hold ($p/n\to c<1$), let $f_{n}(z)$ and $\overline{f}_{n}(z)$ be defined in Definition C.3, where $\mathbf{A}$ is a deterministic matrix with $\|\mathbf{A}\|_{{\mathrm{tr}}}\leq C<+\infty$. For $z\in[0,+\infty)$, it holds that
$\lim\limits_{n\to+\infty}\lim\limits_{z\to 0^{+}}f_{n}(z)=\lim\limits_{z\to 0^{+}}\lim\limits_{n\to+\infty}f_{n}(z),\quad\lim\limits_{n\to+\infty}\lim\limits_{z\to 0^{+}}\overline{f}_{n}(z)=\lim\limits_{z\to 0^{+}}\lim\limits_{n\to+\infty}\overline{f}_{n}(z).$
Therefore, for any fixed $\0z\in\mathbb{C}\backslash(-\infty,0)$, it holds that
$f_{n}(z)-\overline{f}_{n}(z)~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0,\quad f^{\prime}_{n}(z)-\overline{f}^{\prime}_{n}(z)~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0.$
Lemma C.6 established the key asymptotic behavior in Case 2 in Assumption 2.6. However, a notable distinction in Case 3 in Assumption 2.6 is the presence of $p/n\to c\geq 1$. Consequently, the proof technique utilized in Lemma C.6 is no longer applicable due to the inability to uniformly bound $|f^{\prime}_{n}(z)|$ by a trivial constant. Nevertheless, we can still address the case $z=0$ if we make the additional assumption that the number of eigenvalues $K$ tending to infinity remains fixed and $\lambda_{1}/\lambda_{K}\leq C_{\lambda}$:
Lemma C.7.
Suppose that Assumption 2.4, 2.5 and 2.6 Case 3 hold ($p/n\to c\geq 1$), let $f_{n}(z)$ and $\overline{f}_{n}(z)$ be defined in Definition C.3, where $\mathbf{A}$ is a deterministic matrix with $\|\mathbf{A}\|_{{\mathrm{tr}}}\leq C<+\infty$. For $z\in[0,+\infty)$, it holds that
$\lim\limits_{n\to+\infty}\lim\limits_{z\to 0^{+}}f_{n}(z)=\lim\limits_{z\to 0^{+}}\lim\limits_{n\to+\infty}f_{n}(z),\quad\lim\limits_{n\to+\infty}\lim\limits_{z\to 0^{+}}\overline{f}_{n}(z)=\lim\limits_{z\to 0^{+}}\lim\limits_{n\to+\infty}\overline{f}_{n}(z).$
Therefore, for any fixed $\0z\in\mathbb{C}\backslash(-\infty,0)$, it holds that
$f_{n}(z)-\overline{f}_{n}(z)~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0,\quad f^{\prime}_{n}(z)-\overline{f}^{\prime}_{n}(z)~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0.$
Combined with Corollary C.2, Lemma C.6 and C.7, we have the following Proposition.
Proposition C.8.
Suppose that Assumption 2.4, 2.5 and 2.6 hold, let $f_{n}(z)$ and $\overline{f}_{n}(z)$ be defined in Definition C.3, where $\mathbf{A}$ is a deterministic matrix with $\|\mathbf{A}\|_{{\mathrm{tr}}}\leq C<+\infty$. For any fixed $\0z\in\mathbb{C}\backslash(-\infty,0)$, it holds that
$f_{n}(z)-\overline{f}_{n}(z)~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0,\quad f^{\prime}_{n}(z)-\overline{f}^{\prime}_{n}(z)~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0.$
Proof of Proposition C.8.
For Case 1 in Assumption 2.6, we have $\lambda_{\min}(\mathbf{Q})\geq c_{\mathbf{Q}}>0$ in Lemma C.1 and Corollary C.2 above. We may separate $\mathbf{Q}+z\mathbf{I}$ into $\mathbf{Q}-c_{\mathbf{Q}}/2\mathbf{I}+(z+c_{\mathbf{Q}}/2)\mathbf{I}$, therefore the equation still holds when $z=0$. For Case 2 and Case 3 in Assumption 2.6, the results directly hold from Lemma C.6 and C.7. ∎
The difference between Proposition C.8 and Lemma C.4 is the range of $z$. Proposition C.8 covers the case of the convergence when $z=0$. Note that when diverging spikes exist in the matrix ${\bm{\Sigma}}$, $|f_{n}(z)|$ and $|\overline{f}_{n}(z)|$ may tends to 00 as $n$ tends to infinity, we also give the ratio convergence on $f_{n}(z)$ and $\overline{f}_{n}(z)$. We have the following lemma.
Lemma C.9.
Suppose that Assumption 2.4, 2.5 and 2.6 hold. Recall the definitions of $f_{n}(z)$ and $\overline{f}_{n}(z)$ from Definition C.3 where we choose $\mathbf{A}$ to be a semi-positive definite matrix. Then it holds that
$f_{n}(z)/\overline{f}_{n}(z)~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1,\quad f^{\prime}_{n}(z)/\overline{f}^{\prime}_{n}(z)~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1.$
The proof of Lemma C.9 is given in Section C.4. Under these lemmas, we are able to prove Theorem 2.7. We give the proof of Theorem 2.7 in Section D.
C.1 Proof of Lemma C.5
We write $s_{0}(z)$, $s_{1,{\bm{\Sigma}}}(z)$ and $s_{1,\mathbf{Q}}(z)$ as $s_{0}$, $s_{1,{\bm{\Sigma}}}$ and $s_{1,\mathbf{Q}}$ for simplicity. We first prove the existence and uniqueness of $s_{0}$ under the certain condition. Readers may refer to the section 4.1 and 4.2 in Rubio & Mestre (2011) for the existence and uniqueness of $s_{0}$ for $z\in\mathbb{C}_{+}$. With similar proof, the existence and uniqueness of $s_{0}$ when $z\in\mathbb{C}_{-}$ still hold. We prove that when $z\in[0,+\infty)$, $s_{0}>0$ has the unique solution. To prove so, denote $\mathbf{u}=\{u_{i}>0\}$ by the eigenvalue of ${\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{Q}{\bm{\Sigma}}^{-\frac{1}{2}}+z\mathbf{I}$, and $g(s;\mathbf{u})$ by $\frac{c}{p}\sum_{i=1}^{p}\frac{1}{\frac{1}{1+s}+u_{i}}$, hence $s_{0}$ satisfies
$s_{0}=g(s_{0};\mathbf{u})$ $=\frac{c}{p}\sum_{i=1}^{p}\frac{1}{\frac{1}{1+s_{0}}+u_{i}}=\frac{c}{p}\sum_{i=1}^{p}\frac{1}{u_{i}}\cdot\frac{(1+s_{0})u_{i}}{1+u_{i}(1+s_{0})}$
$=\frac{c}{p}\sum_{i=1}^{p}\frac{1}{u_{i}}\cdot\bigg{(}1-\frac{1}{1+u_{i}(1+s_{0})}\bigg{)}>0,$
when $s=0$ we have $g(s;\mathbf{u})-s>0$, when $s\to+\infty$, we have $g(s;\mathbf{u})-s<0$, the existence of $s_{0}$ then comes from Bolzano-Cauchy theorem. As for the uniqueness, assume $g(s;\mathbf{u})-s$ exists two roots $s_{1}$ , $s_{2}$, and $s_{1}<s_{2}$. Note that $g(s;\mathbf{u})-s$ is a concave function, we conclude that $g(s;\mathbf{u})-s$ increases then decreases or just decreases. Note that $g(0;\mathbf{u})>0$ and $g(s;\mathbf{u})-s$ exists roots, we conclude that $g(s;\mathbf{u})-s$ decreases in the neighbourhood of $s_{1}$ and decreases in $[s_{1},+\infty)$. Hence, in the interval $[s_{1},s_{2}]$, $g(s;\mathbf{u})-s\equiv 0$, which violates to the condition $g(s;\mathbf{u})-s$ is concave. Therefore we prove the uniqueness of $s_{0}$. For the bound of $s_{0}$, it holds that
$s_{0}=\frac{c}{p}{\mathrm{tr}}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}+z{\bm{\Sigma}}\bigg{)}^{-1}\leq c\|{\bm{\Sigma}}/p\|_{{\mathrm{tr}}}/\lambda_{\min}(\mathbf{Q}).$
Here, the inequality comes from ${\mathrm{tr}}(\mathbf{A}\mathbf{B})\leq\|\mathbf{A}\|_{{\mathrm{tr}}}\|\mathbf{B}\|_{{\mathop{\mathrm{op}}}}$ and $\lambda_{\min}(\mathbf{A}+\mathbf{B})\geq\lambda_{\min}(\mathbf{A})+\lambda_{\min}(\mathbf{B})$ for non negative matrix $\mathbf{A}$ and $\mathbf{B}$.
We now give the analysis of $s_{1,{\bm{\Sigma}}}$. Denote $u_{i}>0$ by the eigenvalues of ${\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{Q}{\bm{\Sigma}}^{-\frac{1}{2}}+z\mathbf{I}$, then we have
$s_{0}$ $=\frac{c}{p}\sum_{i=1}^{p}\frac{(1+s_{0})}{(1+s_{0})\cdot u_{i}+1},$
$s_{1,{\bm{\Sigma}}}$ $=\left(\frac{s_{1,{\bm{\Sigma}}}}{\left(1+s_{0}\right)^{2}}-1\right)\cdot\frac{c}{p}\cdot\sum_{i=1}^{p}\frac{\left(1+s_{0}\right)^{2}}{\left(1+\left(1+s_{0}\right)\cdot u_{i}\right)^{2}}.$
Define $m=\frac{c}{p}{\mathrm{tr}}\big{(}\frac{\mathbf{I}}{1+s_{0}}+{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{Q}{\bm{\Sigma}}^{-\frac{1}{2}}+z\mathbf{I}\big{)}^{-2}/(1+s_{0})^{2}>0$, then we have
$m$ $=\frac{c}{p}\cdot\sum_{i=1}^{p}\frac{1}{(1+\left(1+s_{0})\cdot u_{i}\right)^{2}}\leq\frac{c}{p(1+s_{0})}\cdot\sum_{i=1}^{p}\frac{\left(1+s_{0}\right)}{1+\left(1+s_{0}\right)\cdot u_{i}}=\frac{s_{0}}{1+s_{0}},$
$s_{1,{\bm{\Sigma}}}$ $=\big{(}s_{1,{\bm{\Sigma}}}-(1+s_{0})^{2}\big{)}\cdot m=\frac{m(1+s_{0})^{2}}{m-1},$
where the first equation is by the definition of $m$ and $s_{0}$, and the second equation is by (C.5). The existence of $s_{1,{\bm{\Sigma}}}$ comes from the second equality. Moreover from $0<m\leq s_{0}/(1+s_{0})$, we conclude that $s_{1,{\bm{\Sigma}}}$ exists, $s_{1,{\bm{\Sigma}}}<0$ and $-s_{1,{\bm{\Sigma}}}\leq s_{0}\cdot(1+s_{0})^{2}$. This completes bound for $s_{1,{\bm{\Sigma}}}$.
We next bound for $s_{1,\mathbf{Q}}$. Denote by $u_{i}$ ($u_{i}>0$) the eigenvalue of the matrix ${\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{Q}{\bm{\Sigma}}^{-\frac{1}{2}}+z\mathbf{I}$, by the definition of $s_{0}$, we can easily see that
$s_{0}=\frac{c}{p}\sum_{i=1}^{p}\frac{(1+s_{0})}{(1+s_{0})\cdot u_{i}+1}.$
By the definition of $s_{1,\mathbf{Q}}$, we can further express $s_{1,\mathbf{Q}}$ as
$s_{1,\mathbf{Q}}$ $=\frac{c}{p}{\mathrm{tr}}~{}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\bigg{(}\frac{s_{1,\mathbf{Q}}}{(1+s_{0})^{2}}{\bm{\Sigma}}-\mathbf{Q}\bigg{)}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}$
$=s_{1,\mathbf{Q}}\cdot\frac{c}{p}\sum_{i=1}^{p}\frac{1}{(1+(1+s_{0})u_{i})^{2}}-\frac{c}{p}{\mathrm{tr}}~{}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\mathbf{Q}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}$
$=s_{1,\mathbf{Q}}\cdot\frac{c}{p}\sum_{i=1}^{p}\frac{1}{(1+(1+s_{0})u_{i})^{2}}-s_{0}+\frac{c}{p}{\mathrm{tr}}~{}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\frac{{\bm{\Sigma}}}{1+s_{0}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}$
$=s_{1,\mathbf{Q}}\cdot\frac{c}{p}\sum_{i=1}^{p}\frac{1}{(1+(1+s_{0})u_{i})^{2}}-s_{0}+\frac{c}{p}\sum_{i=1}^{p}\frac{1+s_{0}}{(1+(1+s_{0})u_{i})^{2}}.$ (C.3)
Here, the second equality comes from the decomposition of the matrix $\mathbf{I}/(1+s_{0})+{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{Q}{\bm{\Sigma}}^{-\frac{1}{2}}$, the third equality comes from $\mathbf{Q}=\mathbf{Q}+{\bm{\Sigma}}/(1+s_{0})-{\bm{\Sigma}}/(1+s_{0})$ and the definition of $s_{0}$, and the last equality still comes from the decomposition of the matrix $\mathbf{I}/(1+s_{0})+{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{Q}{\bm{\Sigma}}^{-\frac{1}{2}}$.
Define $m=\frac{c}{p}{\mathrm{tr}}\big{(}\frac{\mathbf{I}}{1+s_{0}}+{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{Q}{\bm{\Sigma}}^{-\frac{1}{2}}\big{)}^{-2}/(1+s_{0})^{2}$, then we have
$m$ $=\frac{c}{p}\cdot\sum_{i=1}^{p}\frac{1}{(1+(1+s_{0})\cdot u_{i})^{2}}\leq\frac{c}{p(1+s_{0})}\cdot\sum_{i=1}^{p}\frac{1+s_{0}}{1+(1+s_{0})\cdot u_{i}}=\frac{s_{0}}{1+s_{0}}.$
By the inequality above, the coefficient of $s_{1,\mathbf{Q}}$ in the right side in (C.3) is defined as $m$ and is smaller than $1$, we conclude the existence of $s_{1,\mathbf{Q}}$. We first prove that $s_{1,\mathbf{Q}}<0$. By (C.3), if $s_{1,\mathbf{Q}}>0$, then we have
$s_{1,\mathbf{Q}}$ $=m\cdot s_{1,\mathbf{Q}}-s_{0}+\frac{c}{p}\sum_{i=1}^{p}\frac{1+s_{0}}{(1+(1+s_{0})u_{i})^{2}}$
$\leq s_{1,\mathbf{Q}}\cdot\frac{s_{0}}{1+s_{0}}-s_{0}+\frac{c}{p}\sum_{i=1}^{p}\frac{1+s_{0}}{1+(1+s_{0})u_{i}}$
$=s_{1,\mathbf{Q}}\cdot\frac{s_{0}}{1+s_{0}}<s_{1,\mathbf{Q}},$
where the first inequality is by $m\leq s_{0}/(1+s_{0})$ and $1+(1+s_{0})u_{i}\geq 1$, the second equality is by the definition of $s_{0}$ and the last inequality is by $s_{0}/(1+s_{0})<1$. We have a contradictory, therefore we conclude that $s_{1,\mathbf{Q}}<0$. To prove $s_{1,\mathbf{Q}}\geq-s_{0}$, from (C.3) we have
$1+s_{0}+s_{1,\mathbf{Q}}=1+\frac{c}{p}\sum_{i=1}^{p}\frac{1+s_{0}+s_{1,\mathbf{Q}}}{(1+(1+s_{0})u_{i})^{2}}=1+m\cdot(1+s_{0}+s_{1,\mathbf{Q}}).$
From $10<m<$ 1 we conclude that $01+s_{0}+s_{1,\mathbf{Q}}>$ 0. Therefore we have that
$1+s_{0}+s_{1,\mathbf{Q}}=1+m\cdot(1+s_{0}+s_{1,\mathbf{Q}})\geq 1,$
from which we conclude that $s_{1,\mathbf{Q}}\geq-s_{0}$.
C.2 Proof of Lemma C.6
We need to prove that the limit $n\to+\infty$ and $z\to 0^{+}$ can be exchanged in $f_{n}(z)$. We note that
$|f_{n}(z)|$ $=\bigg{|}{\mathrm{tr}}\bigg{(}\frac{{\bm{\Sigma}}^{\frac{1}{2}}\mathbf{Z}^{\top}\mathbf{Z}{\bm{\Sigma}}^{\frac{1}{2}}}{n}+\mathbf{Q}+z{\bm{\Sigma}}\bigg{)}^{-1}\mathbf{A}\bigg{|}\leq\frac{\big{\|}\mathbf{A}\big{\|}_{{\mathrm{tr}}}}{\lambda_{\min}(\mathbf{Q})+z\lambda_{\min}({\bm{\Sigma}})}\leq\frac{\big{\|}\mathbf{A}\big{\|}_{{\mathrm{tr}}}}{\lambda_{\min}(\mathbf{Q})},$
where $\lambda_{\min}(\cdot)$ denote the smallest eigenvalues. We have $|f_{n}(z)|$ is uniformly bounded by $\|\mathbf{A}\|_{{\mathrm{tr}}}/\lambda_{\min}(\mathbf{Q})$ when $z\in[0,+\infty)$. Hence we easily get that $|f_{n}(z)-f_{n}(0)|\leq 2\|\mathbf{A}\|_{{\mathrm{tr}}}/\lambda_{\min}(\mathbf{Q})$. Moreover, when $p/n\to c<1$, we have
$|f_{n}^{\prime}(z)|={\mathrm{tr}}\bigg{[}\bigg{(}\frac{\mathbf{Z}^{\top}\mathbf{Z}}{n}+{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{Q}{\bm{\Sigma}}^{-\frac{1}{2}}+z\mathbf{I}\bigg{)}^{-2}\bigg{]}{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{A}{\bm{\Sigma}}^{-\frac{1}{2}}\leq\frac{\|{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{A}{\bm{\Sigma}}^{-\frac{1}{2}}\|_{{\mathrm{tr}}}}{\lambda_{\min}(\mathbf{Z}^{\top}\mathbf{Z}/n)}\leq\frac{4(\|{\bm{\Sigma}}^{-1}\|_{{\mathop{\mathrm{op}}}}\|\mathbf{A}\|_{{\mathrm{tr}}})}{(1-\sqrt{c})^{4}}.$
The second inequality holds almost surely for large enough $n$, by the Bai-Yin theorem (Bai and Yin, 1993). As its derivatives are bounded, the sequence $f_{n},n=1,2,3,\ldots$ is equicontinuous, and by the Arzela-Ascoli theorem, we deduce that $f_{n}$ converges uniformly to its limit. By the Moore-Osgood theorem, we can exchange limits (as $n,p\rightarrow\infty$ and $z\rightarrow 0^{+}$).
We next prove that the limit $n\to+\infty$ and $z\to 0^{+}$ can be exchanged in $\overline{f}_{n}(z)$. Note that
$\overline{f}_{n}(z)$ $={\mathrm{tr}}\bigg{(}\frac{\mathbf{I}}{1+s_{0}}+{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{Q}{\bm{\Sigma}}^{-\frac{1}{2}}+z\mathbf{I}\bigg{)}^{-1}{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{A}{\bm{\Sigma}}^{-\frac{1}{2}}$
$={\mathrm{tr}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}+z{\bm{\Sigma}}\bigg{)}^{-1}\mathbf{A}\leq\|\mathbf{A}\|_{{\mathrm{tr}}}/\lambda_{\min}(\mathbf{Q}),$
we have $\overline{f}_{n}(z)$ is uniformly bounded by $\|\mathbf{A}\|_{{\mathrm{tr}}}/\lambda_{\min}(\mathbf{Q})$ when $z\in[0,+\infty)$. Recall that
$s_{0}(z)=\frac{c}{p}{\mathrm{tr}}\bigg{(}\frac{\mathbf{I}}{1+s_{0}(z)}+{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{Q}{\bm{\Sigma}}^{-\frac{1}{2}}+z\mathbf{I}\bigg{)}^{-1},$
when $z\geq 0$, Lemma C.5 gives that
$0<s_{0}(z)\leq c\|{\bm{\Sigma}}/p\|_{{\mathrm{tr}}}/\lambda_{\min}(\mathbf{Q}).$
Combined the truth $s_{0}(z)$ is analytic from Rubio & Mestre (2011), we have
$\overline{f}_{n}^{\prime}(z)$ $={\mathrm{tr}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}(z)}+\mathbf{Q}+z{\bm{\Sigma}}\bigg{)}^{-1}\bigg{(}\frac{s_{1}(z)}{(1+s_{0}(z))^{2}}{\bm{\Sigma}}-{\bm{\Sigma}}\bigg{)}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}(z)}+\mathbf{Q}+z{\bm{\Sigma}}\bigg{)}^{-1}\mathbf{A}$
$=\bigg{(}\frac{s_{1}(z)}{(1+s_{0}(z))^{2}}-1\bigg{)}{\mathrm{tr}}\bigg{(}\frac{\mathbf{I}}{1+s_{0}(z)}+{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{Q}{\bm{\Sigma}}^{-\frac{1}{2}}+z\mathbf{I}\bigg{)}^{-2}{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{A}{\bm{\Sigma}}^{-\frac{1}{2}},$
where $s_{1}(z)=s_{0}^{\prime}(z)$ satisfies
$s_{1}(z)$ $=\frac{c}{p}{\mathrm{tr}}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}(z)}+\mathbf{Q}+z{\bm{\Sigma}}\bigg{)}^{-1}\bigg{(}\frac{s_{1}(z)}{(1+s_{0}(z))^{2}}{\bm{\Sigma}}-{\bm{\Sigma}}\bigg{)}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}(z)}+\mathbf{Q}+z{\bm{\Sigma}}\bigg{)}^{-1}.$ (C.4)
Lemma C.5 further gives that $s_{1}(z)$ is uniformly bounded, we can see
$|\overline{f}^{\prime}_{n}(z)|\leq\bigg{|}\frac{s_{1}(z)}{(1+s_{0}(z))^{2}}-1\bigg{|}\cdot c\cdot(1+s_{0}(z))^{2}$
is uniformly bounded. Apply the similar argument above which proved the exchange in $f_{n}(z)$, we obtain that the limit $n\to+\infty$ and $z\to 0^{+}$ can be exchanged in $\overline{f}_{n}(z)$. We complete the proof that $n\to+\infty$ and $z\to 0^{+}$ can be exchanged in $f_{n}(z)$ and $\overline{f}_{n}(z)$.
Therefore, we have
00 0 $=\lim_{z\to 0^{+}}\lim_{n\to+\infty}\overline{f}_{n}(z)-f_{n}(z)=\lim_{n\to+\infty}\lim_{z\to 0^{+}}\overline{f}_{n}(z)-f_{n}(z)$
$=\lim_{n\to+\infty}\overline{f}_{n}(0)-f_{n}(0).$
Here, the first equality comes from Lemma C.4. The internally closed uniform convergence of $\overline{f}_{n}(z)-f_{n}(z)$ is hence extended to the area \$\0\mathbb{C}\backslash(-\infty,0)$, which also indicates that for fixed $\0z\in\mathbb{C}\backslash(-\infty,0)$, $\overline{f}^{\prime}_{n}(z)-f^{\prime}_{n}(z)~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0$.
C.3 Proof of Lemma C.7
When $z\in[0,+\infty)$, from Lemma C.5 we can obtain the existence and uniqueness of $s_{0}>0$. We also have $|f_{n}(z)|$ and $|\overline{f}_{n}(z)|$ are uniformly bounded by $\|\mathbf{A}\|_{{\mathrm{tr}}}/\lambda_{\min}(\mathbf{Q})$, and $|s_{0}(z)|$ is bounded by $c\|{\bm{\Sigma}}/p\|_{{\mathrm{tr}}}/\lambda_{\min}(\mathbf{Q})$. We still define $s_{1}(z)=s_{0}^{\prime}(z)$ which satisfies
$s_{1}(z)$ $=\frac{c}{p}{\mathrm{tr}}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}(z)}+\mathbf{Q}+z{\bm{\Sigma}}\bigg{)}^{-1}\bigg{(}\frac{s_{1}(z)}{(1+s_{0}(z))^{2}}{\bm{\Sigma}}-{\bm{\Sigma}}\bigg{)}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}(z)}+\mathbf{Q}+z{\bm{\Sigma}}\bigg{)}^{-1}$
$=\bigg{(}\frac{s_{1}(z)}{(1+s_{0}(z))^{2}}-1\bigg{)}\cdot\frac{c}{p}{\mathrm{tr}}\bigg{(}\frac{\mathbf{I}}{1+s_{0}(z)}+{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{Q}{\bm{\Sigma}}^{-\frac{1}{2}}+z\mathbf{I}\bigg{)}^{-2},$ (C.5)
then
$\overline{f}_{n}^{\prime}(z)$ $={\mathrm{tr}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}(z)}+\mathbf{Q}+z{\bm{\Sigma}}\bigg{)}^{-1}\bigg{(}\frac{s_{1}(z)}{(1+s_{0}(z))^{2}}{\bm{\Sigma}}-{\bm{\Sigma}}\bigg{)}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}(z)}+\mathbf{Q}+z{\bm{\Sigma}}\bigg{)}^{-1}\mathbf{A}$
$=\bigg{(}\frac{s_{1}(z)}{(1+s_{0}(z))^{2}}-1\bigg{)}{\mathrm{tr}}\bigg{(}\frac{\mathbf{I}}{1+s_{0}(z)}+{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{Q}{\bm{\Sigma}}^{-\frac{1}{2}}+z\mathbf{I}\bigg{)}^{-2}{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{A}{\bm{\Sigma}}^{-\frac{1}{2}}.$
To complete the proof, we only need to show $z\to 0^{+}$ and $n\to+\infty$ can be exchanged in $f_{n}(z)$ and $\overline{f}_{n}(z)$. Hence it remains to prove that $|f_{n}(z)|$ and $|f^{\prime}_{n}(z)|$ are uniformly bounded.
We first prove that $|\overline{f}^{\prime}_{n}(z)|$ is uniformly bounded. To prove so, we have
$|\overline{f}^{\prime}_{n}(z)|\leq|s_{1}(z)-(1+s_{0}(z))^{2}|\cdot\|{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{A}{\bm{\Sigma}}^{-\frac{1}{2}}\|_{{\mathrm{tr}}}.$
It is equal to prove that $|s_{1}(z)|$ is uniformly bounded. By Lemma C.5 we complete the proof that $|\overline{f}^{\prime}_{n}(z)|$ is uniformly bounded.
We next prove that $|f^{\prime}_{n}(z)|$ is uniformly bounded almost surely for large enough $n$. Recall that
$f_{n}^{\prime}(z)$ $={\mathrm{tr}}\bigg{[}\bigg{(}\frac{{\bm{\Sigma}}^{\frac{1}{2}}\mathbf{Z}^{\top}\mathbf{Z}{\bm{\Sigma}}^{\frac{1}{2}}}{n}+\mathbf{Q}+z{\bm{\Sigma}}\bigg{)}^{-1}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}^{\frac{1}{2}}\mathbf{Z}^{\top}\mathbf{Z}{\bm{\Sigma}}^{\frac{1}{2}}}{n}+\mathbf{Q}+z{\bm{\Sigma}}\bigg{)}^{-1}\bigg{]}\mathbf{A},$
if we can prove that for any fixed vector $\mathbf{u}$ with $\|\mathbf{u}\|=1$,
$\mathbf{u}^{\top}\bigg{(}\frac{{\bm{\Sigma}}^{\frac{1}{2}}\mathbf{Z}^{\top}\mathbf{Z}{\bm{\Sigma}}^{\frac{1}{2}}}{n}+\mathbf{Q}+z{\bm{\Sigma}}\bigg{)}^{-1}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}^{\frac{1}{2}}\mathbf{Z}^{\top}\mathbf{Z}{\bm{\Sigma}}^{\frac{1}{2}}}{n}+\mathbf{Q}+z{\bm{\Sigma}}\bigg{)}^{-1}\mathbf{u}$ (C.6)
is uniformly bounded almost surely by constant $C$ for large enough $n$, we will get that
$|f_{n}^{\prime}(z)|$ $=\sum_{i=1}^{p}\lambda_{\mathbf{A},i}{\bm{\xi}}_{\mathbf{A},i}^{\top}\bigg{(}\frac{{\bm{\Sigma}}^{\frac{1}{2}}\mathbf{Z}^{\top}\mathbf{Z}{\bm{\Sigma}}^{\frac{1}{2}}}{n}+\mathbf{Q}+z{\bm{\Sigma}}\bigg{)}^{-1}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}^{\frac{1}{2}}\mathbf{Z}^{\top}\mathbf{Z}{\bm{\Sigma}}^{\frac{1}{2}}}{n}+\mathbf{Q}+z{\bm{\Sigma}}\bigg{)}^{-1}{\bm{\xi}}_{\mathbf{A},i}$
$\leq C\sum_{i=1}^{p}\lambda_{\mathbf{A},i}=C\|\mathbf{A}\|_{{\mathrm{tr}}}<+\infty$
almost surely for large enough $n$. Here, $\lambda_{\mathbf{A},i}$ is the eigenvalue of matrix $\mathbf{A}$ and ${\bm{\xi}}_{\mathbf{A},i}$ is the corresponding eigenvector. Hence, the main target now is to deduce the uniform bound of (C.6). We separate the proof of the uniform bound of (C.6) into the case $z\geq\sqrt{\frac{\lambda_{\min}(\mathbf{Q})}{18cC_{\lambda}K^{2}}}$ and $z\leq\sqrt{\frac{\lambda_{\min}(\mathbf{Q})}{18cC_{\lambda}K^{2}}}$
When $z\geq\sqrt{\frac{\lambda_{\min}(\mathbf{Q})}{18cC_{\lambda}K^{2}}}$, it can be seen that
(C.6)$\leq\frac{\|{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{u}\|_{2}^{2}}{z^{2}}<+\infty,$
which completes the proof that (C.6) is uniformly bounded when $z\geq\sqrt{\frac{\lambda_{\min}(\mathbf{Q})}{18cC_{\lambda}K^{2}}}$.
When $z\leq\sqrt{\frac{\lambda_{\min}(\mathbf{Q})}{18cC_{\lambda}K^{2}}}$, by Assumption 2.6, the number of eigenvalues tending to infinity remains fixed, we can separate ${\bm{\Sigma}}$ by
${\bm{\Sigma}}={\bm{\Gamma}}_{s}{\bm{\Lambda}}_{s}{\bm{\Gamma}}_{s}^{\top}+{\bm{\Gamma}}_{b}{\bm{\Lambda}}_{b}{\bm{\Gamma}}_{b}^{\top},$
here ${\bm{\Gamma}}_{s}\in\mathbb{R}^{p\times K}$ is the eigenvectors corresponding to the diagnal matrix ${\bm{\Lambda}}_{s}\in\mathbb{R}^{K\times K}$ with $\lambda_{\min}({\bm{\Lambda}}_{s})\to+\infty$ as $n\to+\infty$, ${\bm{\Gamma}}_{b}\in\mathbb{R}^{p\times(p-K)}$ is the eigenvectors corresponding to the diagnal matrix ${\bm{\Lambda}}_{b}\in\mathbb{R}^{(p-K)\times(p-K)}$ with $\|{\bm{\Lambda}}_{b}\|_{{\mathop{\mathrm{op}}}}$ uniformly bounded. To give the uniform bound of (C.6), we can easily see that it is equal to give the uniform bound of
$\mathbf{u}^{\top}\bigg{(}\frac{{\bm{\Sigma}}^{\frac{1}{2}}\mathbf{Z}^{\top}\mathbf{Z}{\bm{\Sigma}}^{\frac{1}{2}}}{n}+\mathbf{Q}+z{\bm{\Sigma}}\bigg{)}^{-1}{\bm{\Gamma}}_{s}{\bm{\Lambda}}_{s}{\bm{\Gamma}}_{s}^{\top}\bigg{(}\frac{{\bm{\Sigma}}^{\frac{1}{2}}\mathbf{Z}^{\top}\mathbf{Z}{\bm{\Sigma}}^{\frac{1}{2}}}{n}+\mathbf{Q}+z{\bm{\Sigma}}\bigg{)}^{-1}\mathbf{u}$ (C.7)
due to
$\mathbf{u}^{\top}\bigg{(}\frac{{\bm{\Sigma}}^{\frac{1}{2}}\mathbf{Z}^{\top}\mathbf{Z}{\bm{\Sigma}}^{\frac{1}{2}}}{n}+\mathbf{Q}+z{\bm{\Sigma}}\bigg{)}^{-1}{\bm{\Gamma}}_{b}{\bm{\Lambda}}_{b}{\bm{\Gamma}}_{b}^{\top}\bigg{(}\frac{{\bm{\Sigma}}^{\frac{1}{2}}\mathbf{Z}^{\top}\mathbf{Z}{\bm{\Sigma}}^{\frac{1}{2}}}{n}+\mathbf{Q}+z{\bm{\Sigma}}\bigg{)}^{-1}\mathbf{u}\leq\|{\bm{\Lambda}}_{b}\|_{{\mathop{\mathrm{op}}}}/\lambda_{\min}^{2}(\mathbf{Q})$
is uniformly bounded. Now, for the reason that $K$ is fixed, from (C.7), we only need to prove that for any deterministic $\mathbf{u}$ with $\|\mathbf{u}\|=1$,
$\sqrt{\lambda_{j}}\bigg{|}{\bm{\xi}}_{j}^{\top}\bigg{(}\frac{{\bm{\Sigma}}^{\frac{1}{2}}\mathbf{Z}^{\top}\mathbf{Z}{\bm{\Sigma}}^{\frac{1}{2}}}{n}+\mathbf{Q}+z{\bm{\Sigma}}\bigg{)}^{-1}\mathbf{u}\bigg{|}$ (C.8)
is uniformly bounded for $n$ large enough. Here, $\lambda_{j}\in{\bm{\Lambda}}_{s}$, and ${\bm{\xi}}_{j}$ is the eigenvector corresponding to $\lambda_{j}$. Similar to the decomposition of ${\bm{\Sigma}}$, we decompose ${\bm{\Sigma}}^{\frac{1}{2}}\mathbf{Z}^{\top}\mathbf{Z}{\bm{\Sigma}}^{\frac{1}{2}}/n$ into:
$\frac{{\bm{\Sigma}}^{\frac{1}{2}}\mathbf{Z}^{\top}\mathbf{Z}{\bm{\Sigma}}^{\frac{1}{2}}}{n}=\widehat{\bm{\Gamma}}_{s}\widehat{\bm{\Lambda}}_{s}\widehat{\bm{\Gamma}}_{s}^{\top}+\widehat{\bm{\Gamma}}_{b}\widehat{\bm{\Lambda}}_{b}\widehat{\bm{\Gamma}}_{b}^{\top},$
and rewrite $\widetilde{\mathbf{Q}}=\mathbf{Q}+z{\bm{\Sigma}}$, then (C.8) can be written as
$\sqrt{\lambda_{j}}\big{|}{\bm{\xi}}_{j}^{\top}\big{(}\widehat{\bm{\Gamma}}_{s}\widehat{\bm{\Lambda}}_{s}\widehat{\bm{\Gamma}}_{s}^{\top}+\widehat{\bm{\Gamma}}_{b}\widehat{\bm{\Lambda}}_{b}\widehat{\bm{\Gamma}}_{b}^{\top}+\widetilde{\mathbf{Q}}\big{)}^{-1}\mathbf{u}\big{|}$
$\quad=\sqrt{\lambda_{j}}\left|{\bm{\xi}}_{j}^{\top}\left(\big{[}\widehat{\bm{\Gamma}}_{s},\widehat{\bm{\Gamma}}_{b}\big{]}\begin{bmatrix}\widehat{\bm{\Lambda}}_{s}+\widehat{\bm{\Gamma}}_{s}^{\top}\widetilde{\mathbf{Q}}\widehat{\bm{\Gamma}}_{s}&\widehat{\bm{\Gamma}}_{s}^{\top}\widetilde{\mathbf{Q}}\widehat{\bm{\Gamma}}_{b}\\ \widehat{\bm{\Gamma}}_{b}^{\top}\widetilde{\mathbf{Q}}\widehat{\bm{\Gamma}}_{s}&\widehat{\bm{\Lambda}}_{b}+\widehat{\bm{\Gamma}}_{b}^{\top}\widetilde{\mathbf{Q}}\widehat{\bm{\Gamma}}_{b}\end{bmatrix}\begin{bmatrix}\widehat{\bm{\Gamma}}_{s}^{\top}\\ \widehat{\bm{\Gamma}}_{b}^{\top}\end{bmatrix}\right)^{-1}\mathbf{u}\right|$
$\quad=\sqrt{\lambda_{j}}\left|{\bm{\xi}}_{j}^{\top}\big{[}\widehat{\bm{\Gamma}}_{s},\widehat{\bm{\Gamma}}_{b}\big{]}\begin{bmatrix}\widehat{\bm{\Lambda}}_{s}+\widehat{\bm{\Gamma}}_{s}^{\top}\widetilde{\mathbf{Q}}\widehat{\bm{\Gamma}}_{s}&\widehat{\bm{\Gamma}}_{s}^{\top}\widetilde{\mathbf{Q}}\widehat{\bm{\Gamma}}_{b}\\ \widehat{\bm{\Gamma}}_{b}^{\top}\widetilde{\mathbf{Q}}\widehat{\bm{\Gamma}}_{s}&\widehat{\bm{\Lambda}}_{b}+\widehat{\bm{\Gamma}}_{b}^{\top}\widetilde{\mathbf{Q}}\widehat{\bm{\Gamma}}_{b}\end{bmatrix}^{-1}\begin{bmatrix}\widehat{\bm{\Gamma}}_{s}^{\top}\\ \widehat{\bm{\Gamma}}_{b}^{\top}\end{bmatrix}\mathbf{u}\right|.$
We further define $\mathbf{A}=\widehat{\bm{\Lambda}}_{s}+\widehat{\bm{\Gamma}}_{s}^{\top}\widetilde{\mathbf{Q}}\widehat{\bm{\Gamma}}_{s}$, $\mathbf{B}=\widehat{\bm{\Gamma}}_{s}^{\top}\widetilde{\mathbf{Q}}\widehat{\bm{\Gamma}}_{b}$ and $\mathbf{D}=\widehat{\bm{\Lambda}}_{b}+\widehat{\bm{\Gamma}}_{b}^{\top}\widetilde{\mathbf{Q}}\widehat{\bm{\Gamma}}_{b}$, by the inverse of block matrices, we have
$\begin{bmatrix}\mathbf{A}&\mathbf{B}\\ \mathbf{B}^{\top}&\mathbf{D}\end{bmatrix}^{-1}=\begin{bmatrix}(\mathbf{A}-\mathbf{B}\mathbf{D}^{-1}\mathbf{B}^{\top})^{-1}&-(\mathbf{A}-\mathbf{B}\mathbf{D}^{-1}\mathbf{B}^{\top})^{-1}\mathbf{B}\mathbf{D}^{-1}\\ -\mathbf{D}^{-1}\mathbf{B}^{\top}(\mathbf{A}-\mathbf{B}\mathbf{D}^{-1}\mathbf{B}^{\top})^{-1}&\mathbf{D}^{-1}+\mathbf{D}^{-1}\mathbf{B}^{\top}(\mathbf{A}-\mathbf{B}\mathbf{D}^{-1}\mathbf{B}^{\top})^{-1}\mathbf{B}\mathbf{D}^{-1}\end{bmatrix},$
hence (C.8) can be written as
(C.8)$=\sqrt{\lambda_{j}}\big{|}{\bm{\xi}}_{j}^{\top}\widehat{\bm{\Gamma}}_{s}{\bm{\Delta}}_{ss}\widehat{\bm{\Gamma}}_{s}^{\top}\mathbf{u}+{\bm{\xi}}_{j}^{\top}\widehat{\bm{\Gamma}}_{b}{\bm{\Delta}}_{bs}\widehat{\bm{\Gamma}}_{s}^{\top}\mathbf{u}+{\bm{\xi}}_{j}^{\top}\widehat{\bm{\Gamma}}_{s}{\bm{\Delta}}_{sb}\widehat{\bm{\Gamma}}_{b}^{\top}\mathbf{u}+{\bm{\xi}}_{j}^{\top}\widehat{\bm{\Gamma}}_{b}{\bm{\Delta}}_{bb}\widehat{\bm{\Gamma}}_{b}^{\top}\mathbf{u}\big{|},$ (C.9)
where
${\bm{\Delta}}_{ss}=(\mathbf{A}-\mathbf{B}\mathbf{D}^{-1}\mathbf{B}^{\top})^{-1},$
${\bm{\Delta}}_{bs}={\bm{\Delta}}_{sb}^{\top}=-(\mathbf{A}-\mathbf{B}\mathbf{D}^{-1}\mathbf{B}^{\top})^{-1}\mathbf{B}\mathbf{D}^{-1},$
${\bm{\Delta}}_{bb}=\mathbf{D}^{-1}+\mathbf{D}^{-1}\mathbf{B}^{\top}(\mathbf{A}-\mathbf{B}\mathbf{D}^{-1}\mathbf{B}^{\top})^{-1}\mathbf{B}\mathbf{D}^{-1}.$
We will bound the operator norm of the three terms above. We first note that $\lambda_{\min}(\mathbf{D})\geq\lambda_{\min}(\widehat{\bm{\Gamma}}_{b}^{\top}\widetilde{\mathbf{Q}}\widehat{\bm{\Gamma}}_{b})\geq\lambda_{\min}(\widetilde{\mathbf{Q}})\geq\lambda_{\min}(\mathbf{Q})$, we have $\|\mathbf{D}^{-1}\|_{{\mathop{\mathrm{op}}}}\leq 1/\lambda_{\min}(\mathbf{Q})$ is uniformly bounded. We next bound $\|\mathbf{B}\|_{{\mathop{\mathrm{op}}}}$. Indeed, we have
$\|\mathbf{B}\|_{{\mathop{\mathrm{op}}}}$ $=\big{\|}\widehat{\bm{\Gamma}}_{s}^{\top}(\mathbf{Q}+z{\bm{\Sigma}})\widehat{\bm{\Gamma}}_{b}\big{\|}_{{\mathop{\mathrm{op}}}}\leq\|\mathbf{Q}\|_{{\mathop{\mathrm{op}}}}+z\big{\|}\widehat{\bm{\Gamma}}_{s}^{\top}({\bm{\Gamma}}_{s}{\bm{\Lambda}}_{s}{\bm{\Gamma}}_{s}^{\top}+{\bm{\Gamma}}_{b}{\bm{\Lambda}}_{b}{\bm{\Gamma}}_{b}^{\top})\widehat{\bm{\Gamma}}_{b}\big{\|}_{{\mathop{\mathrm{op}}}}$
$\leq\|\mathbf{Q}\|_{{\mathop{\mathrm{op}}}}+z\|{\bm{\Lambda}}_{b}\|_{{\mathop{\mathrm{op}}}}+z\sum_{j=1}^{K}\lambda_{j}\big{\|}{\bm{\xi}}_{j}^{\top}\widehat{\bm{\Gamma}}_{b}\big{\|},$
where the first inequality is by $\|\mathbf{A}+\mathbf{B}\|_{{\mathop{\mathrm{op}}}}\leq\|\mathbf{A}\|_{{\mathop{\mathrm{op}}}}+\|\mathbf{B}\|_{{\mathop{\mathrm{op}}}}$, and the second inequality is by $\|\mathbf{A}\mathbf{B}\|_{{\mathop{\mathrm{op}}}}\leq\|\mathbf{A}\|_{{\mathop{\mathrm{op}}}}\|\mathbf{B}\|_{{\mathop{\mathrm{op}}}}$ and $\|\widehat{\bm{\Gamma}}_{s}\|_{{\mathop{\mathrm{op}}}},\|{\bm{\Gamma}}_{s}\|_{{\mathop{\mathrm{op}}}},\|\widehat{\bm{\Gamma}}_{b}\|_{{\mathop{\mathrm{op}}}},\|{\bm{\Gamma}}_{b}\|_{{\mathop{\mathrm{op}}}}\leq 1$. We remind readers here ${\bm{\xi}}_{j}\in{\bm{\Gamma}}_{s}$. As demonstrated in Bao et al. (2022)’s Theorem 2.5, for $n$ large enough, almost surely we have
$\lambda_{j}\|{\bm{\xi}}_{j}\widehat{\bm{\Gamma}}_{b}\|^{2}\leq 2\lambda_{j}\bigg{(}1-\frac{\lambda_{j}^{2}-c}{\lambda_{j}(\lambda_{j}+c)}\bigg{)}=\frac{2c(\lambda_{j}+1)}{\lambda_{j}+c}\leq 2c.$ (C.10)
Therefore, when $z\leq\sqrt{\frac{\lambda_{\min}(\mathbf{Q})}{18cC_{\lambda}K^{2}}}$, we have almost surely
$\|\mathbf{B}\|_{{\mathop{\mathrm{op}}}}^{2}/\lambda_{K}\leq(z\cdot K\lambda_{j}\cdot\sqrt{3c}/\sqrt{\lambda_{j}})^{2}/\lambda_{K}\leq z^{2}\cdot K^{2}C_{\lambda}\cdot 3c\leq\frac{\lambda_{\min}(\mathbf{Q})}{6}$
for $n$ large enough. Hence when $z\leq\sqrt{\frac{\lambda_{\min}(\mathbf{Q})}{18cC_{\lambda}K^{2}}}$,
$\frac{\lambda_{\min}(\mathbf{A}-\mathbf{B}\mathbf{D}^{-1}\mathbf{B}^{\top})}{\lambda_{K}}\geq\frac{2\lambda_{K}/3-\|\mathbf{B}\|_{{\mathop{\mathrm{op}}}}^{2}/\lambda_{\min}(\mathbf{Q})}{\lambda_{K}}\geq\frac{1}{2}$
almost surely for all $n$ large enough. Here, the first inequality utilizes ${\widehat{\lambda}}_{K}\geq 2\lambda_{K}/3$ almost surely for large enough $n$. We prove that when $z\leq\sqrt{\frac{\lambda_{\min}(\mathbf{Q})}{18cC_{\lambda}K^{2}}}$,
$\|{\bm{\Delta}}_{ss}\|_{{\mathop{\mathrm{op}}}}\cdot\lambda_{K}=\lambda_{K}/\lambda_{\min}(\mathbf{A}-\mathbf{B}\mathbf{D}^{-1}\mathbf{B}^{\top})\leq 2$ (C.11)
almost surely for all $n$ large enough. Note that $\|\mathbf{B}\|_{{\mathop{\mathrm{op}}}}^{2}/\lambda_{K}\leq\lambda_{\min}(\mathbf{Q})/6$ as long as $z\leq\sqrt{\frac{\lambda_{\min}(\mathbf{Q})}{18cC_{\lambda}K^{2}}}$ for $n$ large enough, we conclude that
$\|{\bm{\Delta}}_{bs}\|_{{\mathop{\mathrm{op}}}}\cdot\sqrt{\lambda_{K}}\leq\lambda_{K}\|{\bm{\Delta}}_{ss}\|_{{\mathop{\mathrm{op}}}}\|\mathbf{B}\|_{{\mathop{\mathrm{op}}}}/(\lambda_{\min}(\mathbf{Q})\cdot\sqrt{\lambda_{K}})\leq\sqrt{\frac{2}{3\lambda_{\min}(\mathbf{Q})}}$ (C.12)
almost surely for all $n$ large enough when $z\leq\sqrt{\frac{\lambda_{\min}(\mathbf{Q})}{18cC_{\lambda}K^{2}}}$. As for $\|{\bm{\Delta}}_{bs}\|_{{\mathop{\mathrm{op}}}}$, we have
$\|{\bm{\Delta}}_{bb}\|_{{\mathop{\mathrm{op}}}}\leq\frac{1}{\lambda_{\min}(\mathbf{Q})}+\frac{1}{\lambda_{\min}^{2}(\mathbf{Q})}\cdot\|\mathbf{B}\|_{{\mathop{\mathrm{op}}}}^{2}/\lambda_{K}\cdot\|{\bm{\Delta}}_{ss}\|_{{\mathop{\mathrm{op}}}}\cdot\lambda_{K}\leq\frac{4}{3\lambda_{\min}(\mathbf{Q})}$ (C.13)
almost surely for all $n$ large enough when $z\leq\sqrt{\frac{\lambda_{\min}(\mathbf{Q})}{18cC_{\lambda}K^{2}}}$. Combined (C.9) with (C.10)-(C.13), we conclude that when $z\leq\sqrt{\frac{\lambda_{\min}(\mathbf{Q})}{18cC_{\lambda}K^{2}}}$,
[…]
$\leq 2\sqrt{\lambda_{j}}\cdot\|{\bm{\xi}}_{j}\widehat{\bm{\Gamma}}_{b}\|\cdot\|{\bm{\Delta}}_{bb}\|_{{\mathop{\mathrm{op}}}}\leq\frac{4\sqrt{2c}}{\lambda_{\min}(\mathbf{Q})}$
almost surely for all $n$ large enough. We hence conclude that when $z\leq\sqrt{\frac{\lambda_{\min}(\mathbf{Q})}{18cC_{\lambda}K^{2}}}$, $|f^{\prime}_{n}(z)|$ is uniformly bounded. Combining with the case $z\geq\sqrt{\frac{\lambda_{\min}(\mathbf{Q})}{18cC_{\lambda}K^{2}}}$, we have that $|f^{\prime}_{n}(z)|$ is uniformly bounded when $z\geq 0$. After similar argument in the proof of Lemma C.6 completes the proof of Lemma C.7.
C.4 Proof of Lemma C.9
Before we give the proof, we prove the following lemma.
Lemma C.10.
For each $n\in N_{+}$ and $i\in[n]$, define sequence $a_{i,n},b_{i,n}\geq 0$ satisfy
$\lim_{n\to+\infty}a_{i,n}/b_{i,n}\to 1$
for any $i\leq n$. Then it holds that
$\lim_{n\to+\infty}\sum_{i=1}^{n}a_{i,n}/\sum_{i=1}^{n}b_{i,n}=1.$
Proof of Lemma C.10.
For any $\varepsilon>0$, there exists $N$, when $n>N$, $|a_{i,n}-b_{i,n}|\leq\varepsilon b_{i,n}$ for any $i\leq n$, therefore for $n>N$ we have
$\sum_{i=1}^{n}a_{i,n}/\sum_{i=1}^{n}b_{i,n}\leq\sum_{i=1}^{n}(1+\varepsilon)b_{i,n}/\sum_{i=1}^{n}b_{i,n}=1+\varepsilon.$
Here, the first inequality is by $a_{i,n}\leq(1+\varepsilon)b_{i,n}$. Similarly we have $\sum_{i=1}^{n}a_{i,n}/\sum_{i=1}^{n}b_{i,n}\geq 1-\varepsilon$. This completes the proof of Lemma C.10. ∎
Proof of Lemma C.9.
Without loss of generality, we can assume $\|\mathbf{A}\|_{{\mathrm{tr}}}$ is bounded. Otherwise, we can over $\|\mathbf{A}\|_{{\mathrm{tr}}}$ in both numerator and denominator.
By Proposition C.8, we have that for any given $z\in\mathbb{C}\backslash(-\infty,0)\bigcup\{0\}$,
$|f_{n}(z)-\overline{f}_{n}(z)|~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0,\quad|f^{\prime}_{n}(z)-\overline{f}^{\prime}_{n}(z)|~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0.$
It is also clear that $|f_{n}(z)|$, $|\overline{f}_{n}(z)|$, $|f^{\prime}_{n}(z)|$ and $|\overline{f}^{\prime}_{n}(z)|$ are all bounded.
When Case 1 in Assumption 2.6 holds ($\|{\bm{\Sigma}}\|_{{\mathop{\mathrm{op}}}}$ is bounded), we decompose $\mathbf{A}$ by $\mathbf{A}=\sum_{i=1}^{p}\lambda_{i,\mathbf{A}}{\bm{\xi}}_{i,\mathbf{A}}{\bm{\xi}}_{i,\mathbf{A}}^{\top}$, where $\lambda_{i,\mathbf{A}}$ is the eigenvalue and ${\bm{\xi}}_{i,\mathbf{A}}$ is the corresponding eigenvector. Then $f_{n}(z)$ and $\overline{f}_{n}(z)$ can be expressed as
$f_{n}(z)$ $=\sum_{i=1}^{p}\lambda_{i,\mathbf{A}}{\bm{\xi}}_{i,\mathbf{A}}^{\top}\bigg{(}\frac{{\bm{\Sigma}}^{\frac{1}{2}}\mathbf{Z}^{\top}\mathbf{Z}{\bm{\Sigma}}^{\frac{1}{2}}}{n}+\mathbf{Q}+z{\bm{\Sigma}}\bigg{)}^{-1}{\bm{\xi}}_{i,\mathbf{A}},$
$\overline{f}_{n}(z)$ $=\sum_{i=1}^{p}\lambda_{i,\mathbf{A}}{\bm{\xi}}_{i,\mathbf{A}}^{\top}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}+z{\bm{\Sigma}}\bigg{)}^{-1}{\bm{\xi}}_{i,\mathbf{A}}.$
Let $a_{i,n}=\lambda_{i,\mathbf{A}}{\bm{\xi}}_{i,\mathbf{A}}^{\top}\big{(}\frac{{\bm{\Sigma}}^{\frac{1}{2}}\mathbf{Z}^{\top}\mathbf{Z}{\bm{\Sigma}}^{\frac{1}{2}}}{n}+\mathbf{Q}+z{\bm{\Sigma}}\big{)}^{-1}{\bm{\xi}}_{i,\mathbf{A}}$ and $b_{i,n}=\lambda_{i,\mathbf{A}}{\bm{\xi}}_{i,\mathbf{A}}^{\top}\big{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}+z{\bm{\Sigma}}\big{)}^{-1}{\bm{\xi}}_{i,\mathbf{A}}$, we have that $a_{i,n}/\lambda_{i,\mathbf{A}}-b_{i,n}/\lambda_{i,\mathbf{A}}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0$ and $b_{i,n}/\lambda_{i,\mathbf{A}}$ has the constant lower bound, hence $a_{i,n}/b_{i,n}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1$ for any $i\leq p$. From Lemma C.10, we have $f_{n}(z)/\overline{f}_{n}(z)~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1$.
When Case 2 and 3 in Assumption 2.6 hold, we have
${f}_{n}(z)$ $={\mathrm{tr}}\bigg{(}\frac{\mathbf{Z}^{\top}\mathbf{Z}}{n}+{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{Q}{\bm{\Sigma}}^{-\frac{1}{2}}+z\mathbf{I}\bigg{)}^{-1}{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{A}{\bm{\Sigma}}^{-\frac{1}{2}},$
$\overline{f}_{n}(z)$ $={\mathrm{tr}}\bigg{(}\frac{\mathbf{I}}{1+s_{0}}+{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{Q}{\bm{\Sigma}}^{-\frac{1}{2}}+z\mathbf{I}\bigg{)}^{-1}{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{A}{\bm{\Sigma}}^{-\frac{1}{2}}.$
Decompose ${\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{A}{\bm{\Sigma}}^{-\frac{1}{2}}$ by ${\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{A}{\bm{\Sigma}}^{-\frac{1}{2}}=\sum_{i=1}^{p}\lambda_{i}{\bm{\xi}}_{i}{\bm{\xi}}_{i}^{\top}$, where $\lambda_{i}$ is the eigenvalue and ${\bm{\xi}}_{i}$ is the corresponding eigenvector. Hence $f_{n}(z)$ and $\overline{f}_{n}(z)$ can be expressed as
$f_{n}(z)$ $=\sum_{i=1}^{p}\lambda_{i}{\bm{\xi}}_{i}^{\top}\bigg{(}\frac{\mathbf{Z}^{\top}\mathbf{Z}}{n}+{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{Q}{\bm{\Sigma}}^{-\frac{1}{2}}+z\mathbf{I}\bigg{)}^{-1}{\bm{\xi}}_{i},$
$\overline{f}_{n}(z)$ $=\sum_{i=1}^{p}\lambda_{i}{\bm{\xi}}_{i}^{\top}\bigg{(}\frac{\mathbf{I}}{1+s_{0}}+{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{Q}{\bm{\Sigma}}^{-\frac{1}{2}}+z\mathbf{I}\bigg{)}^{-1}{\bm{\xi}}_{i}.$
Similarly, let $a_{i,n}=\lambda_{i}{\bm{\xi}}_{i}^{\top}\big{(}\frac{\mathbf{Z}^{\top}\mathbf{Z}}{n}+{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{Q}{\bm{\Sigma}}^{-\frac{1}{2}}+z\mathbf{I}\big{)}^{-1}{\bm{\xi}}_{i}$ and $b_{i,n}=\lambda_{i}{\bm{\xi}}_{i}^{\top}\big{(}\frac{\mathbf{I}}{1+s_{0}}+{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{Q}{\bm{\Sigma}}^{-\frac{1}{2}}+z\mathbf{I}\big{)}^{-1}{\bm{\xi}}_{i}$, we have that $a_{i,n}/\lambda_{i}-b_{i,n}/\lambda_{i}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0$ and $b_{i,n}/\lambda_{i}$ has the constant lower bound, hence $a_{i,n}/b_{i,n}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1$ for any $i\leq p$. From Lemma C.10, we have $f_{n}(z)/\overline{f}_{n}(z)~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1$. The proof for $f^{\prime}_{n}(z)/\overline{f}^{\prime}_{n}(z)~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1$ is similar to $f_{n}(z)/\overline{f}_{n}(z)~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1$, and we thus omit it. ∎
Appendix D Proof of Theorem 2.7
Let $s_{0}(z)$ be defined in Definition C.3, and set $s_{0}=s_{0}(0)$. We rewrite Theorem 2.7 and extend it as follows:
Theorem D.1.
Suppose Assumptions 2.2, 2.4, 2.5 and 2.6 hold. For any $\mathbf{Q}\in\mathcal{Q}$, a good estimator $\widehat{SR}(\mathbf{Q})$ for $SR(\mathbf{Q})$ which is defined in (2.4), is given as follows.
$\widehat{SR}(\mathbf{Q})=\frac{T_{n,1}(\mathbf{Q})}{\sqrt{\big{|}\widehat{T}_{n,2}(\mathbf{Q})\big{|}}},\quad\text{where}\quad\widehat{T}_{n,2}(\mathbf{Q})=\frac{{\mathrm{tr}}(\widehat{{\bm{\Sigma}}}+\mathbf{Q})^{-1}\widehat{\bm{\Sigma}}(\widehat{{\bm{\Sigma}}}+\mathbf{Q})^{-1}\mathbf{A}}{\big{(}1-\frac{c}{p}{\mathrm{tr}}\widehat{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\big{)}^{2}}.$
The following properties hold:
  1. 1.
    If $\|\mathbf{A}\|_{{\mathrm{tr}}}$ is bounded, it holds that $\widehat{T}_{n,2}(\mathbf{Q})-T_{n,2}(\mathbf{Q})~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0$.
  2. 2.
    If $\mathbf{A}$ is semi-positive definite, it holds that $\widehat{T}_{n,2}(\mathbf{Q})/T_{n,2}(\mathbf{Q})~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1$ and $\widehat{SR}(\mathbf{Q})/SR(\mathbf{Q})~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1$.
  3. 3.
    If $\mathbf{A}$ is semi-positive definite and ${\mathrm{tr}}\Big{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\Big{)}^{-1}\mathbf{A}$ is bounded, it holds that $\widehat{SR}(\mathbf{Q})-SR(\mathbf{Q})~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0$.
We then give the proof of Theorem D.1. In the proof of Theorem D.1, we give the precise asymptotic of $T_{n,1}(\mathbf{Q})$ and $T_{n,2}(\mathbf{Q})$, and then construct statistics only related to $\mathbf{X}$ which share the same asymptotic of $T_{n,1}(\mathbf{Q})$ and $T_{n,2}(\mathbf{Q})$, finally we use these statistics to estimate $SR(\mathbf{Q})$. Define
$g({\bm{\Sigma}})={\mathrm{tr}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\mathbf{A}.$
Recall the definition of $T_{n,1}(\mathbf{Q})$ and $T_{n,2}(\mathbf{Q})$ in (2.2) and (2.3) that
$T_{n,1}(\mathbf{Q})={\mathrm{tr}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\mathbf{A},\quad T_{n,2}(\mathbf{Q})={\mathrm{tr}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\mathbf{A}.$
The proposition below gives us the precise asymptotic of $T_{n,1}(\mathbf{Q})$ and $T_{n,2}(\mathbf{Q})$ and is proved in Appendix D.1.
Proposition D.2.
Suppose that Assumptions in Theorem D.1 hold. Let $T_{n,1}(\mathbf{Q})$ and $T_{n,2}(\mathbf{Q})$ defined in (2.2) and (2.3) respectively, then if $\|\mathbf{A}\|_{{\mathrm{tr}}}$ is bounded, it holds that
$T_{n,1}(\mathbf{Q})-{\mathrm{tr}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\mathbf{A}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0,$
$T_{n,2}(\mathbf{Q})-\frac{(1+s_{0})^{2}-s_{1,{\bm{\Sigma}}}}{(1+s_{0})^{2}}g({\bm{\Sigma}})~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0.$
Here, $s_{0}$ and $s_{1,{\bm{\Sigma}}}$ uniquely solve the equations
$s_{0}=\frac{c}{p}{\mathrm{tr}}~{}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1},\quad s_{1,{\bm{\Sigma}}}=\bigg{(}\frac{s_{1,{\bm{\Sigma}}}}{(1+s_{0})^{2}}-1\bigg{)}\cdot\frac{c}{p}{\mathrm{tr}}~{}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}.$
Moreover, if $\mathbf{A}$ is semi-positive definite, then
$T_{n,1}(\mathbf{Q})/{\mathrm{tr}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\mathbf{A}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1,$
$T_{n,2}(\mathbf{Q})/\bigg{(}\frac{(1+s_{0})^{2}-s_{1,{\bm{\Sigma}}}}{(1+s_{0})^{2}}g({\bm{\Sigma}})\bigg{)}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1.$
The next proposition further indicates us how to construct the statistics only related to $\mathbf{X}$ which share the same asymptotic of $T_{n,1}(\mathbf{Q})$ and $T_{n,2}(\mathbf{Q})$.
Proposition D.3.
Let $s_{0}$ be defined in Proposition D.2. For any matrix $\mathbf{A}$ with $\|\mathbf{A}\|_{{\mathrm{tr}}}$ bounded, define
$\widehat{g}(\widehat{\bm{\Sigma}})=\widehat{T}_{1}-\widehat{T}_{2}={\mathrm{tr}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\widehat{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\mathbf{A},$
where $\widehat{T}_{1}={\mathrm{tr}}(\widehat{{\bm{\Sigma}}}+\mathbf{Q})^{-1}\mathbf{A}$, $\widehat{T}_{2}={\mathrm{tr}}(\widehat{{\bm{\Sigma}}}+\mathbf{Q})^{-1}\mathbf{Q}(\widehat{{\bm{\Sigma}}}+\mathbf{Q})^{-1}\mathbf{A}$ and $\widehat{\bm{\Sigma}}=\mathbf{X}^{\top}\mathbf{X}/n$. Then under the same conditions as Proposition D.2,
$\widehat{g}(\widehat{\bm{\Sigma}})-\frac{(1+s_{0}+s_{1,\mathbf{Q}})g({\bm{\Sigma}})}{(1+s_{0})^{2}}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0,$
where $s_{1,\mathbf{Q}}$ solves the following equation
$s_{1,\mathbf{Q}}=\frac{c}{p}{\mathrm{tr}}~{}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\bigg{(}\frac{s_{1,\mathbf{Q}}}{(1+s_{0})^{2}}{\bm{\Sigma}}-\mathbf{Q}\bigg{)}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}.$
Moreover, if $\mathbf{A}$ is semi-positive definite, then
$\frac{(1+s_{0})^{2}\widehat{g}(\widehat{\bm{\Sigma}})/g({\bm{\Sigma}})}{1+s_{0}+s_{1,\mathbf{Q}}}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1.$
Proposition D.3 constructs the estimation of the key part in the asymptotic of $T_{n,2}(\mathbf{Q})$ given in Proposition D.2, and is proved in Appendix D.2. Now, the only thing remained for us is to construct statistics which give precise estimations of $s_{0}$, $s_{1,{\bm{\Sigma}}}$ and $s_{1,\mathbf{Q}}$ respectively. The next proposition established such precise estimations.
Proposition D.4.
Let $s_{0}$, $s_{1,{\bm{\Sigma}}}$ and $s_{1,\mathbf{Q}}$ be defined in Proposition D.2 and Proposition D.3 respectively. Define
$f_{1}(\widehat{{\bm{\Sigma}}})=\frac{1}{p}{\mathrm{tr}}\mathbf{Q}(\widehat{{\bm{\Sigma}}}+\mathbf{Q})^{-1},\quad f_{2}(\widehat{{\bm{\Sigma}}})=\frac{1}{p}{\mathrm{tr}}\mathbf{Q}(\widehat{{\bm{\Sigma}}}+\mathbf{Q})^{-1}\mathbf{Q}(\widehat{{\bm{\Sigma}}}+\mathbf{Q})^{-1},$
then under the same condition as Proposition D.2,
$\frac{1}{s_{0}}\cdot\frac{c\big{(}1-f_{1}(\widehat{{\bm{\Sigma}}})\big{)}}{1+c\big{(}f_{1}(\widehat{{\bm{\Sigma}}})-1\big{)}}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1,\quad\frac{1}{s_{1,\mathbf{Q}}}\cdot c\cdot\frac{f_{2}(\widehat{{\bm{\Sigma}}})-f_{1}(\widehat{{\bm{\Sigma}}})}{\big{(}1+c(f_{1}(\widehat{{\bm{\Sigma}}})-1)\big{)}^{2}}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1,$
$\frac{1}{s_{1,{\bm{\Sigma}}}}\cdot c\cdot\frac{\big{(}-1+2f_{1}(\widehat{{\bm{\Sigma}}})-f_{2}(\widehat{{\bm{\Sigma}}})+c(f_{1}(\widehat{{\bm{\Sigma}}})-1)^{2}\big{)}}{\big{(}1+c(f_{1}(\widehat{{\bm{\Sigma}}})-1)\big{)}^{4}}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1.$
Moreover, it holds from the conclusion above that
$\frac{1+s_{0}+s_{1,\mathbf{Q}}}{(1+s_{0})^{2}-s_{1,{\bm{\Sigma}}}}/\bigg{(}1-\frac{c}{p}{\mathrm{tr}}\widehat{\bm{\Sigma}}\big{(}\widehat{\bm{\Sigma}}+\mathbf{Q}\big{)}^{-1}\bigg{)}^{2}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}1.$
Proposition D.4 establishes the statistics only related to $\mathbf{X}$ to estimate the constants, and is proved in Appendix D.3. With Proposition D.2-D.4 above, we give the proof of Theorem D.1.
Proof of Theorem D.1.
Below we prove the three conclusions one by one.
The first conclusion. If $\|\mathbf{A}\|_{{\mathrm{tr}}}$ is bounded, from Proposition D.2 and D.3, we conclude that
$T_{n,2}(\mathbf{Q})-\frac{(1+s_{0})^{2}-s_{1,{\bm{\Sigma}}}}{(1+s_{0})^{2}}g({\bm{\Sigma}})~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0,\quad\frac{(1+s_{0})^{2}\widehat{g}(\widehat{\bm{\Sigma}})}{1+s_{0}+s_{1,\mathbf{Q}}}-g({\bm{\Sigma}})~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0.$
Here, $\widehat{g}(\widehat{\bm{\Sigma}})={\mathrm{tr}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\widehat{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\mathbf{A}$. Therefore we have
$T_{n,2}(\mathbf{Q})-\frac{(1+s_{0})^{2}-s_{1,{\bm{\Sigma}}}}{1+s_{0}+s_{1,\mathbf{Q}}}\widehat{g}(\widehat{\bm{\Sigma}})~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0.$ (D.1)
Proposition D.4 further gives us the estimation of $s_{0}$, $s_{1,{\bm{\Sigma}}}$ and $s_{1,\mathbf{Q}}$. Note that $s_{0}=s_{0}(0)$, $s_{1,{\bm{\Sigma}}}=s_{1,{\bm{\Sigma}}}(0)$, where $s_{0}(z)$ and $s_{1,{\bm{\Sigma}}}(z)$ are defined in Lemma C.5. We have $s_{0}>0$, $-s_{0}(1+s_{0})^{2}<s_{1,{\bm{\Sigma}}}<0$ and $1\leq 1+s_{0}+s_{1,\mathbf{Q}}\leq 1+s_{0}$ are all bounded. Easy to conclude that $\frac{(1+s_{0})^{2}-s_{1,{\bm{\Sigma}}}}{(1+s_{0})^{2}}$, $\frac{1+s_{0}+s_{1,\mathbf{Q}}}{(1+s_{0})^{2}}$, $\frac{(1+s_{0})^{2}-s_{1,{\bm{\Sigma}}}}{1+s_{0}+s_{1,\mathbf{Q}}}=\Theta(1)$. From Proposition D.4 we also have
$\frac{1}{(1+c(f_{1}(\widehat{\Sigma})-1))^{2}}/\frac{(1+s_{0})^{2}-s_{1,{\bm{\Sigma}}}}{1+s_{0}+s_{1,\mathbf{Q}}}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1.$ (D.2)
Thus combined (D.1) with (D.2) we have
$T_{n,2}(\mathbf{Q})-\frac{\widehat{g}(\widehat{\bm{\Sigma}})}{(1+c(f_{1}(\widehat{\Sigma})-1))^{2}}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0.$
This directly indicates that
$\widehat{T}_{n,2}(\mathbf{Q})-T_{n,2}(\mathbf{Q})~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0.$
The second conclusion. When $\mathbf{A}$ is semi-positive definite, from Proposition D.2 and D.3, we conclude that
$\frac{(1+s_{0})^{2}-s_{1,{\bm{\Sigma}}}}{(1+s_{0})^{2}}g({\bm{\Sigma}})/T_{n,2}(\mathbf{Q})~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1,\quad\frac{(1+s_{0})^{2}\widehat{g}(\widehat{\bm{\Sigma}})}{1+s_{0}+s_{1,\mathbf{Q}}}/g({\bm{\Sigma}})~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1.$
Hence we have
$\frac{(1+s_{0})^{2}-s_{1,{\bm{\Sigma}}}}{1+s_{0}+s_{1,\mathbf{Q}}}\cdot\frac{\widehat{g}(\widehat{\bm{\Sigma}})}{T_{n,2}(\mathbf{Q})}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1.$ (D.3)
It only requires to prove that
$\frac{1+s_{0}+s_{1,\mathbf{Q}}}{(1+s_{0})^{2}-s_{1,{\bm{\Sigma}}}}\cdot\frac{1}{(1+c(f_{1}(\widehat{\Sigma})-1))^{2}}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1.$ (D.4)
Proposition D.4 shows that (D.4) holds. Combined (D.4) with (D.3), we have
$\frac{1}{(1+c(f_{1}(\widehat{\Sigma})-1))^{2}}\cdot\frac{\widehat{g}(\widehat{\bm{\Sigma}})}{T_{n,2}(\mathbf{Q})}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1,$
which completes the proof of the second conclusion in Theorem D.1.
The third conclusion. As for the last conclusion in Theorem D.1, we both have $\mathbf{A}$ is semi-positive definite and $\mathbf{A}$ is bounded. Therefore by $\widehat{SR}(\mathbf{Q})/SR(\mathbf{Q})~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1$, we can conclude that
$\big{(}\widehat{SR}(\mathbf{Q})-SR(\mathbf{Q})\big{)}/SR(\mathbf{Q})~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0.$
As long as we prove that $SR(\mathbf{Q})$ is almost surely upper bounded, we can easily conclude that $\widehat{SR}(\mathbf{Q})-SR(\mathbf{Q})~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0$. It remains for us to prove that $SR(\mathbf{Q})$ is almost surely upper bounded. By Proposition D.2, we have
$\frac{{\mathrm{tr}}\Big{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\Big{)}^{-1}\mathbf{A}}{SR(\mathbf{Q})\cdot\sqrt{\frac{(1+s_{0})^{2}-s_{1,{\bm{\Sigma}}}}{(1+s_{0})^{2}}\cdot g({\bm{\Sigma}})}}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1.$
Also note that by Lemma C.5, the adjust factor $\frac{(1+s_{0})^{2}-s_{1,{\bm{\Sigma}}}}{(1+s_{0})^{2}}$ has the lower bound $1$ and the upper bound $1+s_{0}$, it only requires us to prove that ${\mathrm{tr}}\big{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\big{)}^{-1}\mathbf{A}/g({\bm{\Sigma}})$ is upper bounded. Recall that $g({\bm{\Sigma}})={\mathrm{tr}}\big{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\big{)}^{-1}{\bm{\Sigma}}\big{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\big{)}^{-1}\mathbf{A}$, we have that
$g({\bm{\Sigma}})$ $={\mathrm{tr}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\mathbf{A}$
$=(1+s_{0}){\mathrm{tr}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\frac{{\bm{\Sigma}}}{1+s_{0}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\mathbf{A}$
$=(1+s_{0})\bigg{[}{\mathrm{tr}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\mathbf{A}-{\mathrm{tr}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\mathbf{Q}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\mathbf{A}\bigg{]}.$ (D.5)
Here, the last equality is by ${\bm{\Sigma}}/(1+s_{0})={\bm{\Sigma}}/(1+s_{0})+\mathbf{Q}-\mathbf{Q}$. By the semi-positive definite of ${\bm{\Sigma}}$, $\mathbf{Q}$ and $\mathbf{A}$, we can easily conclude that
${\mathrm{tr}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}$ $\mathbf{Q}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\mathbf{A}$
$\leq\lambda_{\max}\bigg{(}\mathbf{Q}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\bigg{)}\cdot{\mathrm{tr}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\mathbf{A}$
$\leq\frac{\lambda_{\max}({\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{Q}{\bm{\Sigma}}^{-\frac{1}{2}})}{\lambda_{\max}({\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{Q}{\bm{\Sigma}}^{-\frac{1}{2}}+\mathbf{I}/(1+s_{0}))}\cdot{\mathrm{tr}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\mathbf{A}$
$\leq\frac{C_{\mathbf{Q}}(1+s_{0})}{C_{\mathbf{Q}}(1+s_{0})+1}\cdot{\mathrm{tr}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\mathbf{A}.$ (D.6)
Here, $\lambda_{\max}(\cdot)$ and $\lambda_{\min}(\cdot)$ denote by the largest and smallest eigenvalue of the matrix. The first inequality is by ${\mathrm{tr}}(\mathbf{A}\mathbf{B})\leq\lambda_{\max}(\mathbf{A}){\mathrm{tr}}\mathbf{B}$ for $\mathbf{A},\mathbf{B}\geq 0$, the second inequality is by $\lambda_{\max}\Big{(}\mathbf{Q}\Big{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\Big{)}^{-1}\Big{)}=\lambda_{\max}\Big{(}\Big{(}\frac{\mathbf{I}}{1+s_{0}}+{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{Q}{\bm{\Sigma}}^{-\frac{1}{2}}\Big{)}^{-1}\cdot{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{Q}{\bm{\Sigma}}^{-\frac{1}{2}}\Big{)}$. Therefore, plugging (D.6) into (D.5), we have
$g({\bm{\Sigma}})$ $\geq\frac{1+s_{0}}{C_{\mathbf{Q}}(1+s_{0})+1}\cdot{\mathrm{tr}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\mathbf{A}$
which indicates that
$\frac{{\mathrm{tr}}\Big{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\Big{)}^{-1}\mathbf{A}}{\sqrt{g({\bm{\Sigma}})}}$ $\leq\sqrt{\frac{C_{\mathbf{Q}}(1+s_{0})+1}{1+s_{0}}}\cdot\sqrt{{\mathrm{tr}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\mathbf{A}}$
<+<$\infty$.< + ∞ .
Therefore we conclude that $SR(\mathbf{Q})$ is almost surely bounded, which completes the proof of the third conclusion in Theorem D.1. ∎
D.1 Proof of Proposition D.2
Based on the previous result in Section C, We give the following proposition which covers Proposition D.2 by taking $\lambda=0$ and $\mathbf{B}={\bm{\Sigma}}$.
Proposition D.5.
Define the positive matrix $\mathbf{B}$ taking value on ${\bm{\Sigma}}$ or $\mathbf{Q}$, and let
$T_{n,1}(\lambda,\mathbf{B})$ $={\mathrm{tr}}\bigg{(}\frac{\mathbf{X}^{\top}\mathbf{X}}{n}+\mathbf{Q}+\lambda\mathbf{B}\bigg{)}^{-1}\mathbf{A},$
$T_{n,2}(\lambda,\mathbf{B})$ $={\mathrm{tr}}\bigg{(}\frac{\mathbf{X}^{\top}\mathbf{X}}{n}+\mathbf{Q}+\lambda\mathbf{B}\bigg{)}^{-1}\mathbf{B}\bigg{(}\frac{\mathbf{X}^{\top}\mathbf{X}}{n}+\mathbf{Q}+\lambda\mathbf{B}\bigg{)}^{-1}\mathbf{A}.$
Under the same condition as Proposition D.2, if $\lambda\geq 0$, it holds that
$T_{n,1}(\lambda,\mathbf{B})-{\mathrm{tr}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0,\mathbf{B}}(\lambda)}+\mathbf{Q}+\lambda\mathbf{B}\bigg{)}^{-1}\mathbf{A}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0,$
$T_{n,2}(\lambda,\mathbf{B})-{\mathrm{tr}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0,\mathbf{B}}(\lambda)}+\mathbf{Q}+\lambda\mathbf{B}\bigg{)}^{-1}\bigg{(}\mathbf{B}-\frac{s_{1,\mathbf{B}}(\lambda)}{(1+s_{0,\mathbf{B}}(\lambda))^{2}}{\bm{\Sigma}}\bigg{)}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0,\mathbf{B}}(\lambda)}+\mathbf{Q}+\lambda\mathbf{B}\bigg{)}^{-1}\mathbf{A}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0.$
Here, $s_{0,\mathbf{B}}(\lambda)$ and $s_{1,\mathbf{B}}(\lambda)$ uniquely solve the following equations
$s_{0,\mathbf{B}}(\lambda)=\frac{c}{p}{\mathrm{tr}}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0,\mathbf{B}}(\lambda)}+\mathbf{Q}+\lambda\mathbf{B}\bigg{)}^{-1},$
$s_{1,\mathbf{B}}(\lambda)=\frac{c}{p}{\mathrm{tr}}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0,\mathbf{B}}(\lambda)}+\mathbf{Q}+\lambda\mathbf{B}\bigg{)}^{-1}\bigg{(}\frac{s_{1,\mathbf{B}}(\lambda)}{(1+s_{0,\mathbf{B}}(\lambda))^{2}}{\bm{\Sigma}}-\mathbf{B}\bigg{)}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0,\mathbf{B}}(\lambda)}+\mathbf{Q}+\lambda\mathbf{B}\bigg{)}^{-1}.$
Moreover, if $\mathbf{A}$ is semi-positive definite and $\mathbf{B}={\bm{\Sigma}}$, it holds that
$T_{n,1}(\lambda,\mathbf{B})/{\mathrm{tr}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0,\mathbf{B}}(\lambda)}+\mathbf{Q}+\lambda\mathbf{B}\bigg{)}^{-1}\mathbf{A}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1,$
$T_{n,2}(\lambda,\mathbf{B})/{\mathrm{tr}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0,\mathbf{B}}(\lambda)}+\mathbf{Q}+\lambda\mathbf{B}\bigg{)}^{-1}\bigg{(}\mathbf{B}-\frac{s_{1,\mathbf{B}}(\lambda)}{(1+s_{0,\mathbf{B}}(\lambda))^{2}}{\bm{\Sigma}}\bigg{)}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0,\mathbf{B}}(\lambda)}+\mathbf{Q}+\lambda\mathbf{B}\bigg{)}^{-1}\mathbf{A}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1.$
Proof of Proposition D.5.
It is easy to check that $|T_{n,1}(\lambda,\mathbf{B})|$ is bounded. For the convergence of $T_{n,1}(\lambda,\mathbf{B})$, we can easy to see that
$T_{n,1}(\lambda,\mathbf{B}={\bm{\Sigma}})={\mathrm{tr}}{\bm{\Sigma}}^{-\frac{1}{2}}\bigg{(}\frac{\mathbf{Z}^{\top}\mathbf{Z}}{n}+{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{Q}{\bm{\Sigma}}^{-\frac{1}{2}}+\lambda\mathbf{I}_{p}\bigg{)}^{-1}{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{A},$
$T_{n,1}(\lambda,\mathbf{B}=\mathbf{Q})={\mathrm{tr}}\bigg{(}\frac{\mathbf{X}^{\top}\mathbf{X}}{n}+\mathbf{Q}+\lambda\mathbf{Q}\bigg{)}^{-1}\mathbf{A}.$
Here $\mathbf{X}={\bm{\Sigma}}^{\frac{1}{2}}\mathbf{Z}$. If $\mathbf{B}={\bm{\Sigma}}$, from Proposition C.8 we can see
$T_{n,1}(\lambda,\mathbf{B})-{\mathrm{tr}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0,\mathbf{B}}(\lambda)}+\mathbf{Q}+\lambda\mathbf{B}\bigg{)}^{-1}\mathbf{A}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0,$
$\frac{\partial T_{n,1}(\lambda,\mathbf{B})}{\partial\lambda}-\frac{\partial{\mathrm{tr}}\Big{(}\frac{{\bm{\Sigma}}}{1+s_{0,\mathbf{B}}(\lambda)}+\mathbf{Q}+\lambda\mathbf{B}\Big{)}^{-1}\mathbf{A}}{\partial\lambda}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0$
holds for $\lambda\geq 0$. Here, we have
$\frac{{\mathop{\mathrm{d}}}s_{0,\mathbf{B}}(\lambda)}{{\mathop{\mathrm{d}}}\lambda}=\frac{{\mathop{\mathrm{d}}}\frac{c}{p}{\mathrm{tr}}{\bm{\Sigma}}\Big{(}\frac{{\bm{\Sigma}}}{1+s_{0,\mathbf{B}}(\lambda)}+\mathbf{Q}+\lambda\mathbf{B}\Big{)}^{-1}}{{\mathop{\mathrm{d}}}\lambda}.$
If $\mathbf{B}=\mathbf{Q}$, we can still have that for $\0z\in\mathbb{C}\backslash(-\infty,0]$,
$T_{n,1}(z,\mathbf{B})-{\mathrm{tr}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0,\mathbf{B}}(z)}+\mathbf{Q}+z\mathbf{B}\bigg{)}^{-1}\mathbf{A}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0,$
and the convergence can be extended to $z=0$. The analytic function is internally uniform convergence, therefore the derivatives also converge. We conclude that when $\lambda\geq 0$,
$T_{n,1}(\lambda,\mathbf{B})-{\mathrm{tr}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0,\mathbf{B}}(\lambda)}+\mathbf{Q}+\lambda\mathbf{B}\bigg{)}^{-1}\mathbf{A}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0,$
$\frac{\partial T_{n,1}(\lambda,\mathbf{B})}{\partial\lambda}-\frac{\partial{\mathrm{tr}}\Big{(}\frac{{\bm{\Sigma}}}{1+s_{0,\mathbf{B}}(\lambda)}+\mathbf{Q}+\lambda\mathbf{B}\Big{)}^{-1}\mathbf{A}}{\partial\lambda}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0.$
Direct calculation gives us
$\frac{\partial T_{n,1}(\lambda,\mathbf{B})}{\partial\lambda}=-T_{n,2}(\lambda,\mathbf{B}),$
$\frac{\partial{\mathrm{tr}}\Big{(}\frac{{\bm{\Sigma}}}{1+s_{0,\mathbf{B}}(\lambda)}+\mathbf{Q}+\lambda\mathbf{B}\Big{)}^{-1}\mathbf{A}}{\partial\lambda}$
$\qquad\qquad=-{\mathrm{tr}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0,\mathbf{B}}(\lambda)}+\mathbf{Q}+\lambda\mathbf{B}\bigg{)}^{-1}\bigg{(}\mathbf{B}-\frac{s^{\prime}_{0,\mathbf{B}}(\lambda)}{(1+s_{0,\mathbf{B}}(\lambda))^{2}}{\bm{\Sigma}}\bigg{)}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0,\mathbf{B}}(\lambda)}+\mathbf{Q}+\lambda\mathbf{B}\bigg{)}^{-1}\mathbf{A},$
$s^{\prime}_{0,\mathbf{B}}(\lambda)=\frac{c}{p}{\mathrm{tr}}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0,\mathbf{B}}(\lambda)}+\mathbf{Q}+\lambda\mathbf{B}\bigg{)}^{-1}\bigg{(}\frac{s^{\prime}_{0,\mathbf{B}}(\lambda)}{(1+s_{0,\mathbf{B}}(\lambda))^{2}}{\bm{\Sigma}}-\mathbf{B}\bigg{)}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0,\mathbf{B}}(\lambda)}+\mathbf{Q}+\lambda\mathbf{B}\bigg{)}^{-1}.$
By the definition of $s_{1,\mathbf{B}}(\lambda)$ in Proposition D.5 we have $s^{\prime}_{0,\mathbf{B}}(\lambda)=s_{1,\mathbf{B}}(\lambda)$, therefore the convergence of $T_{n,2}(\lambda,\mathbf{B})$ is proved by
$\bigg{|}\frac{\partial T_{n,1}(\lambda,\mathbf{B})}{\partial\lambda}-\frac{\partial{\mathrm{tr}}\Big{(}\frac{{\bm{\Sigma}}}{1+s_{0,\mathbf{B}}(\lambda)}+\mathbf{Q}+\lambda\mathbf{B}\Big{)}^{-1}\mathbf{A}}{\partial\lambda}\bigg{|}$
$\quad=\bigg{|}T_{n,2}(\lambda)-{\mathrm{tr}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0,\mathbf{B}}(\lambda)}+\mathbf{Q}+\lambda\mathbf{B}\bigg{)}^{-1}\bigg{(}\mathbf{B}-\frac{s^{\prime}_{0}(\lambda)}{(1+s_{0,\mathbf{B}}(\lambda))^{2}}{\bm{\Sigma}}\bigg{)}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0,\mathbf{B}}(\lambda)}+\mathbf{Q}+\lambda\mathbf{B}\bigg{)}^{-1}\mathbf{A}\bigg{|}$
$\quad~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0.$
This completes the first conclusion of Proposition D.5. When $\mathbf{A}$ is semi-positive definite and $\mathbf{B}={\bm{\Sigma}}$, we have $T_{n,1}(\lambda,\mathbf{B})=f_{n}(\lambda)$ and $T_{n,2}(\lambda,\mathbf{B})=f^{\prime}_{n}(\lambda)$. The conclusion directly comes from Lemma C.9. This completes the proof of Proposition D.5. ∎
By taking $\mathbf{B}={\bm{\Sigma}}$, Proposition D.2 holds directly from Proposition D.5.
D.2 Proof of Proposition D.3
It is easy to get that the constants $s_{0}$ defined in Proposition D.2 and $s_{1,\mathbf{B}}$ ($\mathbf{B}={\bm{\Sigma}}$ or $\mathbf{B}=\mathbf{Q}$ defined in Proposition D.2 and D.3) are defined by $s_{0}=s_{0,\mathbf{B}}(0)$ and $s_{1,\mathbf{B}}=s_{1,\mathbf{B}}(0)$ in Proposition D.5. We remind the readers that $s_{1,\mathbf{B}}$ solves the following equation:
$s_{1,\mathbf{B}}=\frac{c}{p}{\mathrm{tr}}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0,\mathbf{B}}(\lambda)}+\mathbf{Q}\bigg{)}^{-1}\bigg{(}\frac{s_{1,\mathbf{B}}}{(1+s_{0})^{2}}{\bm{\Sigma}}-\mathbf{B}\bigg{)}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}.$
Moreover, from Proposition D.5, if we let $\mathbf{B}=\mathbf{Q}$ and $\lambda=0$, then
${\mathrm{tr}}\big{(}\widehat{{\bm{\Sigma}}}+\mathbf{Q}\big{)}^{-1}\mathbf{Q}\big{(}\widehat{{\bm{\Sigma}}}+\mathbf{Q}\big{)}^{-1}\mathbf{A}-{\mathrm{tr}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\bigg{(}\mathbf{Q}-\frac{s_{1,\mathbf{Q}}}{(1+s_{0})^{2}}{\bm{\Sigma}}\bigg{)}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\mathbf{A}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0,$
by the definition $g({\bm{\Sigma}})={\mathrm{tr}}\Big{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\Big{)}^{-1}{\bm{\Sigma}}\Big{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\Big{)}^{-1}\mathbf{A}$ and $\widehat{T}_{2}={\mathrm{tr}}\big{(}\widehat{{\bm{\Sigma}}}+\mathbf{Q}\big{)}^{-1}\mathbf{Q}\big{(}\widehat{{\bm{\Sigma}}}+\mathbf{Q}\big{)}^{-1}\mathbf{A}$, the equation above can be rewrited as
$\widehat{T}_{2}-{\mathrm{tr}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\mathbf{Q}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\mathbf{A}+\frac{s_{1,\mathbf{Q}}}{(1+s_{0})^{2}}g({\bm{\Sigma}})~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0.$ (D.7)
(D.7) gives an equation to estimate $g({\bm{\Sigma}})$. However, another equation is required to eliminate ${\mathrm{tr}}\big{(}{\bm{\Sigma}}/(1+s_{0})+\mathbf{Q}\big{)}^{-1}\mathbf{Q}\big{(}{\bm{\Sigma}}/(1+s_{0})+\mathbf{Q}\big{)}^{-1}\mathbf{A}$ in (D.7) and get the desired estimation of $g({\bm{\Sigma}})$. To get the other equation, from the definition of $g({\bm{\Sigma}})$ we have
$g({\bm{\Sigma}})$ $={\mathrm{tr}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\mathbf{A}$
$=(1+s_{0}){\mathrm{tr}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\frac{{\bm{\Sigma}}}{1+s_{0}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\mathbf{A}$
$=(1+s_{0})\bigg{[}{\mathrm{tr}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\mathbf{A}-{\mathrm{tr}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\mathbf{Q}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\mathbf{A}\bigg{]}.$ (D.8)
Combing (D.8) with ${\mathrm{tr}}\big{(}\widehat{{\bm{\Sigma}}}+\mathbf{Q}\big{)}^{-1}\mathbf{A}-{\mathrm{tr}}\big{(}{\bm{\Sigma}}/(1+s_{0})+\mathbf{Q}\big{)}^{-1}\mathbf{A}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0$ in Proposition D.5, we successfully obtain another estimation of $g({\bm{\Sigma}})$:
$g({\bm{\Sigma}})-(1+s_{0})\bigg{[}{\mathrm{tr}}\bigg{(}\widehat{{\bm{\Sigma}}}+\mathbf{Q}\bigg{)}^{-1}\mathbf{A}-{\mathrm{tr}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\mathbf{Q}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\mathbf{A}\bigg{]}$
$\qquad\qquad=g({\bm{\Sigma}})-(1+s_{0})\bigg{[}\widehat{T}_{1}-{\mathrm{tr}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\mathbf{Q}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\mathbf{A}\bigg{]}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0.$ (D.9)
Here, we remind the readers that $\widehat{T}_{1}={\mathrm{tr}}\big{(}\widehat{{\bm{\Sigma}}}+\mathbf{Q}\big{)}^{-1}\mathbf{A}$. From (D.7) and (D.9), we could successfully eliminate ${\mathrm{tr}}\big{(}{\bm{\Sigma}}/(1+s_{0})+\mathbf{Q}\big{)}^{-1}\mathbf{Q}\big{(}{\bm{\Sigma}}/(1+s_{0})+\mathbf{Q}\big{)}^{-1}\mathbf{A}$ and achieve that
[…]
Therefore ${\mathrm{tr}}\big{(}{\bm{\Sigma}}/(1+s_{0})+\mathbf{Q}\big{)}^{-1}\mathbf{Q}\big{(}{\bm{\Sigma}}/(1+s_{0})+\mathbf{Q}\big{)}^{-1}\mathbf{A}$ is eliminated and we conclude that $\frac{(1+s_{0})^{2}(\widehat{T}_{1}-\widehat{T}_{2})}{1+s_{0}+s_{1,\mathbf{Q}}}-g({\bm{\Sigma}})~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0$. Similar to the proof of Lemma C.9, when $\mathbf{A}$ is semi-positive definite, we have $\frac{(1+s_{0})^{2}(\widehat{T}_{1}-\widehat{T}_{2})/g({\bm{\Sigma}})}{1+s_{0}+s_{1,\mathbf{Q}}}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1$.
D.3 Proof of Proposition D.4
We first recall that $f_{1}(\widehat{{\bm{\Sigma}}})=\frac{1}{p}{\mathrm{tr}}\mathbf{Q}(\widehat{{\bm{\Sigma}}}+\mathbf{Q})^{-1},f_{2}(\widehat{{\bm{\Sigma}}})=\frac{1}{p}{\mathrm{tr}}(\widehat{{\bm{\Sigma}}}+\mathbf{Q})^{-1}\mathbf{Q}(\widehat{{\bm{\Sigma}}}+\mathbf{Q})^{-1}\mathbf{Q}$, and the definitions of $s_{0}$, $s_{1,{\bm{\Sigma}}}$ and $s_{1,\mathbf{Q}}$ are given by the solutions of the following equations:
$s_{0}=\frac{c}{p}{\mathrm{tr}}~{}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1},\quad s_{1,{\bm{\Sigma}}}=\bigg{(}\frac{s_{1,{\bm{\Sigma}}}}{(1+s_{0})^{2}}-1\bigg{)}\cdot\frac{c}{p}{\mathrm{tr}}~{}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1},$
$s_{1,\mathbf{Q}}=\frac{c}{p}{\mathrm{tr}}~{}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\bigg{(}\frac{s_{1,\mathbf{Q}}}{(1+s_{0})^{2}}{\bm{\Sigma}}-\mathbf{Q}\bigg{)}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}.$
Then similar to the proof of Proposition D.5, for $\mathbf{A}$ semi-positive definite we have
${\mathrm{tr}}\bigg{(}\frac{\mathbf{X}^{\top}\mathbf{X}}{n}+\mathbf{Q}\bigg{)}^{-1}\mathbf{A}/{\mathrm{tr}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\mathbf{A}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1.$
Let $\mathbf{A}=\mathbf{Q}/p$, we have the following equation:
$f_{1}(\widehat{{\bm{\Sigma}}})/\bigg{(}\frac{1}{p}{\mathrm{tr}}~{}\mathbf{Q}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\bigg{)}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1,$ (D.10)
Moreover, if we let $\mathbf{A}=\mathbf{Q}/p$ in the Proposition D.3, we can also have
$\frac{(1+s_{0})^{2}\big{[}f_{1}(\widehat{{\bm{\Sigma}}})-f_{2}(\widehat{{\bm{\Sigma}}})\big{]}}{1+s_{0}+s_{1,\mathbf{Q}}}/\bigg{(}\frac{1}{p}{\mathrm{tr}}~{}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\mathbf{Q}\bigg{)}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1.$ (D.11)
With equations (D.10) and (D.11), we are now able to prove the conclusions in Proposition D.4.
The estimation of $s_{0}$. For $s_{0}=\frac{c}{p}{\mathrm{tr}}~{}{\bm{\Sigma}}\Big{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\Big{)}^{-1}$, we could further simplify the equations as
$s_{0}$ $=\frac{c}{p}{\mathrm{tr}}~{}{\bm{\Sigma}}\Big{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\Big{)}^{-1}=c(1+s_{0})\cdot\bigg{[}1-\frac{1}{p}{\mathrm{tr}}~{}\mathbf{Q}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\bigg{]}.$
Combined the equation above with (D.10), we have
$\bigg{(}1-\frac{s_{0}}{c(1+s_{0})}\bigg{)}/f_{1}(\widehat{{\bm{\Sigma}}})~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1,\quad\text{which indicates}\quad\frac{c\big{(}1-f_{1}(\widehat{{\bm{\Sigma}}})\big{)}/s_{0}}{1-c\big{(}1-f_{1}(\widehat{{\bm{\Sigma}}})\big{)}}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1.$
The estimation of $s_{1,\mathbf{Q}}$. For $s_{1,\mathbf{Q}}=\frac{c}{p}{\mathrm{tr}}~{}{\bm{\Sigma}}\Big{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\Big{)}^{-1}\Big{(}\frac{s_{1,\mathbf{Q}}}{(1+s_{0})^{2}}{\bm{\Sigma}}-\mathbf{Q}\Big{)}\Big{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\Big{)}^{-1}$, we could further simplify the equations as
$s_{1,\mathbf{Q}}$ $=\frac{cs_{1,\mathbf{Q}}}{p}{\mathrm{tr}}~{}\frac{{\bm{\Sigma}}}{1+s_{0}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\frac{{\bm{\Sigma}}}{1+s_{0}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}$
$\qquad-\frac{c}{p}{\mathrm{tr}}~{}\mathbf{Q}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}$
$=\frac{cs_{1,\mathbf{Q}}}{p}{\mathrm{tr}}~{}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}-\mathbf{Q}\bigg{)}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\frac{{\bm{\Sigma}}}{1+s_{0}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}$
$\qquad-\frac{c}{p}{\mathrm{tr}}~{}\mathbf{Q}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}$
$=\frac{cs_{1,\mathbf{Q}}}{p}{\mathrm{tr}}~{}\frac{{\bm{\Sigma}}}{1+s_{0}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}-\frac{cs_{1,\mathbf{Q}}}{p}{\mathrm{tr}}~{}\mathbf{Q}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\frac{{\bm{\Sigma}}}{1+s_{0}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}$
$\qquad-\frac{c}{p}{\mathrm{tr}}~{}\mathbf{Q}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}$
$=\frac{cs_{1,\mathbf{Q}}}{p}{\mathrm{tr}}~{}\frac{{\bm{\Sigma}}}{1+s_{0}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}$
$\qquad-\bigg{(}\frac{cs_{1,\mathbf{Q}}}{p(1+s_{0})}+\frac{c}{p}\bigg{)}\cdot{\mathrm{tr}}~{}\mathbf{Q}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}$
$=cs_{1,\mathbf{Q}}\bigg{[}1-\frac{1}{p}{\mathrm{tr}}~{}\mathbf{Q}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\bigg{]}$
$\qquad-\bigg{(}\frac{cs_{1,\mathbf{Q}}}{p(1+s_{0})}+\frac{c}{p}\bigg{)}\cdot{\mathrm{tr}}~{}\mathbf{Q}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}$
$=\frac{s_{0}s_{1,\mathbf{Q}}}{1+s_{0}}-\bigg{(}\frac{cs_{1,\mathbf{Q}}}{p(1+s_{0})}+\frac{c}{p}\bigg{)}\cdot{\mathrm{tr}}~{}\mathbf{Q}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}.$
Here, the last equality is by the definition of $s_{0}$. From the last equality above we have
$s_{1,\mathbf{Q}}=-(1+s_{0}+s_{1,\mathbf{Q}})\cdot\frac{c}{p}{\mathrm{tr}}~{}\mathbf{Q}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}.$ (D.12)
Combined (D.12) with (D.11), we conclude that
$\frac{-c(1+s_{0})^{2}(f_{1}(\widehat{{\bm{\Sigma}}})-f_{2}(\widehat{{\bm{\Sigma}}}))}{s_{1,\mathbf{Q}}}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1.$ (D.13)
Note also that $\bigg{(}1+\frac{c(1-f_{1}(\widehat{{\bm{\Sigma}}}))}{1-c\big{(}1-f_{1}(\widehat{{\bm{\Sigma}}})\big{)}}\bigg{)}/(1+s_{0})~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1$ holds, replacing $(1+s_{0})$ with $\frac{1}{1-c\big{(}1-f_{1}(\widehat{{\bm{\Sigma}}})\big{)}}$ into (D.13) gives
$c\cdot\frac{(f_{2}(\widehat{{\bm{\Sigma}}})-f_{1}(\widehat{{\bm{\Sigma}}}))/s_{1,\mathbf{Q}}}{\big{(}c\big{(}1-f_{1}(\widehat{{\bm{\Sigma}}})\big{)}-1\big{)}^{2}}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1.$
The estimation of $s_{1,{\bm{\Sigma}}}$. For $s_{1,{\bm{\Sigma}}}=\Big{(}\frac{s_{1,{\bm{\Sigma}}}}{(1+s_{0})^{2}}-1\Big{)}\cdot\frac{c}{p}{\mathrm{tr}}~{}{\bm{\Sigma}}\Big{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\Big{)}^{-1}{\bm{\Sigma}}\Big{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\Big{)}^{-1}$, we could further simplify the equations as
$s_{1,{\bm{\Sigma}}}$ $=\bigg{(}\frac{s_{1,{\bm{\Sigma}}}}{(1+s_{0})^{2}}-1\bigg{)}\cdot\frac{c}{p}{\mathrm{tr}}~{}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}$
$=\frac{c(s_{1,{\bm{\Sigma}}}-(1+s_{0})^{2})}{p}{\mathrm{tr}}~{}\frac{{\bm{\Sigma}}}{1+s_{0}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\frac{{\bm{\Sigma}}}{1+s_{0}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}$
$=c(s_{1,{\bm{\Sigma}}}-(1+s_{0})^{2})\bigg{(}1-\frac{1}{p}{\mathrm{tr}}~{}\mathbf{Q}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\bigg{)}$
$\qquad-\frac{c(s_{1,{\bm{\Sigma}}}-(1+s_{0})^{2})}{p(1+s_{0})}\cdot{\mathrm{tr}}~{}\mathbf{Q}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}$
$=\frac{c(s_{1,{\bm{\Sigma}}}-(1+s_{0})^{2})}{1+s_{0}}\bigg{(}\frac{s_{0}}{c}-\frac{1}{p}{\mathrm{tr}}~{}\mathbf{Q}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\bigg{)}.$
Here, the last equality uses the fact $1-\frac{1}{p}{\mathrm{tr}}~{}\mathbf{Q}\Big{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\Big{)}^{-1}=\frac{s_{0}}{c(1+s_{0})}$. Combined the results above with (D.10) and (D.11), we conclude that
$c[s_{1,{\bm{\Sigma}}}-(1+s_{0})^{2}]\cdot\bigg{(}\frac{s_{0}}{c(1+s_{0})}-(1+s_{0})\cdot\frac{f_{1}(\widehat{{\bm{\Sigma}}})-f_{2}(\widehat{{\bm{\Sigma}}})}{1+s_{0}+s_{1,\mathbf{Q}}}\bigg{)}/s_{1,{\bm{\Sigma}}}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1,$
Furthermore, by (D.13) and replacing $(1+s_{0})$ with $\frac{1}{1-c\big{(}1-f_{1}(\widehat{{\bm{\Sigma}}})\big{)}}$ we have
$c\bigg{[}s_{1,{\bm{\Sigma}}}-\frac{1}{\big{(}1-c\big{(}1-f_{1}(\widehat{{\bm{\Sigma}}})\big{)}\big{)}^{2}}\bigg{]}\cdot\Bigg{(}1-f_{1}(\widehat{{\bm{\Sigma}}})-\frac{f_{1}(\widehat{{\bm{\Sigma}}})-f_{2}(\widehat{{\bm{\Sigma}}})}{1+\frac{c(f_{1}(\widehat{{\bm{\Sigma}}})-f_{2}(\widehat{{\bm{\Sigma}}}))}{c(1-f_{1}(\widehat{{\bm{\Sigma}}}))-1}}\Bigg{)}/s_{1,{\bm{\Sigma}}}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1.$
Solve this linear equation with $s_{1,{\bm{\Sigma}}}$ and we conclude that
$c\cdot\frac{-1+2f_{1}(\widehat{{\bm{\Sigma}}})-f_{2}(\widehat{{\bm{\Sigma}}})+c(f_{1}(\widehat{{\bm{\Sigma}}})-1)^{2}}{\big{(}1+c(f_{1}(\widehat{{\bm{\Sigma}}})-1)\big{)}^{4}\cdot s_{1,{\bm{\Sigma}}}}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1.$
Note that $s_{0},|s_{1,{\bm{\Sigma}}}|$ and $|s_{1,\mathbf{Q}}|$ are bounded by Lemma C.5, direct calculation gives
$\frac{1+s_{0}+s_{1,\mathbf{Q}}}{(1+s_{0})^{2}-s_{1,{\bm{\Sigma}}}}/\bigg{(}1-\frac{c}{p}{\mathrm{tr}}\widehat{\bm{\Sigma}}\big{(}\widehat{\bm{\Sigma}}+\mathbf{Q}\big{)}^{-1}\bigg{)}^{2}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}1.$
We complete the proof of Proposition D.4.
Appendix E Proof of Theorem 3.2
In this section, we give the proof of Theorem 3.2. Before the proof, we give the following lemmas. The following lemma summarizes some results from Section C and Section D.
Lemma E.1.
Under the condition of Theorem 3.2, suppose that $\mathbf{A}$ is semi-positive, it holds that
$\frac{(1+s_{0})^{2}}{(1+s_{0})^{2}-s_{1,{\bm{\Sigma}}}}\cdot\frac{{\mathrm{tr}}(\widehat{{\bm{\Sigma}}}+\mathbf{Q})^{-1}{\bm{\Sigma}}(\widehat{{\bm{\Sigma}}}+\mathbf{Q})^{-1}\mathbf{A}}{{\mathrm{tr}}\Big{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\Big{)}^{-1}{\bm{\Sigma}}\Big{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\Big{)}^{-1}\mathbf{A}}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1.$
$\frac{(1+s_{0})^{2}}{1+s_{0}+s_{1,\mathbf{Q}}}\cdot\frac{{\mathrm{tr}}(\widehat{{\bm{\Sigma}}}+\mathbf{Q})^{-1}\widehat{\bm{\Sigma}}(\widehat{{\bm{\Sigma}}}+\mathbf{Q})^{-1}\mathbf{A}}{{\mathrm{tr}}\Big{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\Big{)}^{-1}{\bm{\Sigma}}\Big{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\Big{)}^{-1}\mathbf{A}}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1.$
Proof of Lemma E.1.
The first result comes from Proposition D.2, and the second result comes from Proposition D.3. ∎
Lemma C.5 and E.1 jointly show that ${\mathrm{tr}}(\widehat{{\bm{\Sigma}}}+\mathbf{Q})^{-1}{\bm{\Sigma}}(\widehat{{\bm{\Sigma}}}+\mathbf{Q})^{-1}\mathbf{A}$, ${\mathrm{tr}}(\widehat{{\bm{\Sigma}}}+\mathbf{Q})^{-1}\widehat{\bm{\Sigma}}(\widehat{{\bm{\Sigma}}}+\mathbf{Q})^{-1}\mathbf{A}$ and ${\mathrm{tr}}\Big{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\Big{)}^{-1}{\bm{\Sigma}}\Big{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\Big{)}^{-1}\mathbf{A}$ are in the same order.The following lemma indicates that ${\mathrm{tr}}\Big{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\Big{)}^{-1}{\bm{\Sigma}}\Big{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\Big{)}^{-1}\mathbf{A}=\Theta\Big{(}{\mathrm{tr}}\Big{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\Big{)}^{-1}\mathbf{A}\Big{)}$. This equality will further simplify our analysis in the proof of Theorem 3.2.
Lemma E.2.
Under the condition of Theorem 3.2, suppose that $\mathbf{A}$ is semi-positive, it holds that
${\mathrm{tr}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\mathbf{A}=\Theta\bigg{(}{\mathrm{tr}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\mathbf{A}\bigg{)}.$
Moreover, if $\mathbf{A}={\bm{\xi}}{\bm{\xi}}^{\top}$, then
${\mathrm{tr}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\mathbf{A}=\Theta\big{(}\|{\bm{\Sigma}}^{-\frac{1}{2}}{\bm{\xi}}\|_{2}^{2}\big{)}.$
The proof of Lemma E.2 can be found in Section E.1. We give an additional lemma, which is applied to handle the dependence of $\mathbf{A}$ ($\mathbf{A}$ is rank-one and semi-positive in Theorem 3.2) and $\widehat{\bm{\Sigma}}$ in the proof of Theorem 3.2.
Lemma E.3.
Recall that
$\mathcal{A}_{rr}=\mathbf{r}^{\top}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}\mathbf{r},\quad\mathcal{A}_{r1}=\mathbf{r}^{\top}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}{\bm{1}},\quad\mathcal{A}_{11}={\bm{1}}^{\top}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}{\bm{1}}.$
Define
$\alpha_{n}=\mathbf{r}^{\top}\big{(}\widehat{\bm{\Sigma}}+\mathbf{Q}\big{)}^{-1}\mathbf{r}-\mu_{0}\mathbf{r}^{\top}\big{(}\widehat{\bm{\Sigma}}+\mathbf{Q}\big{)}^{-1}{\bm{1}},\quad\alpha_{0}=\mathcal{A}_{rr}-\mu_{0}\mathcal{A}_{r1},$
$\beta_{n}=\mu_{0}{\bm{1}}^{\top}\big{(}\widehat{\bm{\Sigma}}+\mathbf{Q}\big{)}^{-1}{\bm{1}}-\mathbf{r}^{\top}\big{(}\widehat{\bm{\Sigma}}+\mathbf{Q}\big{)}^{-1}{\bm{1}},\quad\beta_{0}=\mu_{0}\mathcal{A}_{11}-\mathcal{A}_{r1}.$
Given some large enough constant $C$, then under the condition of Theorem 3.2, the following properties hold:
  1. 1.
    If $\mu_{0}\leq C\sqrt{\mathcal{A}_{rr}/\mathcal{A}_{11}}$, it holds that $\frac{\alpha_{n}-\alpha_{0}}{\mathcal{A}_{rr}}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0,\quad\frac{\beta_{n}-\beta_{0}}{\sqrt{\mathcal{A}_{rr}\cdot\mathcal{A}_{11}}}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0$.
  2. 2.
    If $\mu_{0}\geq C\sqrt{\mathcal{A}_{rr}/\mathcal{A}_{11}}$, it holds that $\frac{\alpha_{n}-\alpha_{0}}{\mu_{0}\sqrt{\mathcal{A}_{rr}\mathcal{A}_{11}}}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0,\quad\frac{\beta_{n}-\beta_{0}}{\mu_{0}\mathcal{A}_{11}}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0$.
An additional conclusion is $D=\Theta(\mathcal{A}_{rr}\mathcal{A}_{11})$ where $D$ is defined in (3.5).
The proof of Lemma E.3 can be found in Section E.2. With the lemmas above, we are now ready to prove Theorem 3.2.
Proof of Theorem 3.2.
We first give the proof of ratio consistency. Recall that
$\hfil~{}&\mathbf{g}=D^{-1}\big{[}B(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{1}}-A(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\mathbf{r}\big{]},~{}\mathbf{h}=D^{-1}\big{[}C(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\mathbf{r}-A(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{1}}\big{]},\hfil~{}\\ \hfil~{}&A=\mathbf{r}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{1}},~{}B=\mathbf{r}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\mathbf{r},~{}C={\bm{1}}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{1}},~{}D=BC-A^{2},\hfil~{}\end{array}$ (E.3)
and
$\begin{split}&\sigma_{0}^{2}=\mathbf{w}^{*\top}{\bm{\Sigma}}\mathbf{w}^{*}=(\mathbf{g}+\mu_{0}\cdot\mathbf{h})^{\top}{\bm{\Sigma}}(\mathbf{g}+\mu_{0}\cdot\mathbf{h}),\\ &\widehat{\sigma}^{2}=(\mathbf{g}+\mu_{0}\cdot\mathbf{h})^{\top}\widehat{\bm{\Sigma}}(\mathbf{g}+\mu_{0}\cdot\mathbf{h})/(1-c/p\cdot{\mathrm{tr}}\widehat{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1})^{2}.\end{split}$ (E.4)
The first conclusion in Theorem 3.2 states that $\widehat{\sigma}^{2}/\sigma_{0}^{2}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1$. By direct calculation of (E.4), we observe that
$\begin{split}&\sigma_{0}^{2}=(\alpha_{n}{\bm{1}}+\beta_{n}\mathbf{r})^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}(\alpha_{n}{\bm{1}}+\beta_{n}\mathbf{r})/D^{2},\\ &\widehat{\sigma}^{2}=\frac{(\alpha_{n}{\bm{1}}+\beta_{n}\mathbf{r})^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\widehat{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}(\alpha_{n}{\bm{1}}+\beta_{n}\mathbf{r})}{D^{2}(1-c/p\cdot{\mathrm{tr}}\widehat{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1})^{2}}.\end{split}$ (E.5)
Here, we have
$\begin{split}&\alpha_{n}=\mathbf{r}^{\top}\big{(}\widehat{\bm{\Sigma}}+\mathbf{Q}\big{)}^{-1}\mathbf{r}-\mu_{0}\mathbf{r}^{\top}\big{(}\widehat{\bm{\Sigma}}+\mathbf{Q}\big{)}^{-1}{\bm{1}},\\ &\beta_{n}=\mu_{0}{\bm{1}}^{\top}\big{(}\widehat{\bm{\Sigma}}+\mathbf{Q}\big{)}^{-1}{\bm{1}}-\mathbf{r}^{\top}\big{(}\widehat{\bm{\Sigma}}+\mathbf{Q}\big{)}^{-1}{\bm{1}}.\end{split}$ (E.6)
It is clear that $\alpha_{n}{\bm{1}}+\beta_{n}\mathbf{r}$ is dependent with $\widehat{\bm{\Sigma}}$ and is not deterministic, hence we can not directly apply the results from Theorem 2.7. Fortunately, the dependence of $\alpha_{n}{\bm{1}}+\beta_{n}\mathbf{r}$ and $\widehat{\bm{\Sigma}}$ is only due to the two numbers $\alpha_{n}$ and $\beta_{n}$. We further define
$\alpha_{0}=\mathcal{A}_{rr}-\mu_{0}\mathcal{A}_{r1},\quad\beta_{0}=\mu_{0}\mathcal{A}_{11}-\mathcal{A}_{r1}.$ (E.7)
In the following steps of the proof of the ratio consistency, we mainly focus on the proof of the two key results below:
$\frac{(1+s_{0})^{2}}{(1+s_{0})^{2}-s_{1,{\bm{\Sigma}}}}\cdot\frac{(\alpha_{n}{\bm{1}}+\beta_{n}\mathbf{r})^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}(\alpha_{n}{\bm{1}}+\beta_{n}\mathbf{r})}{(\alpha_{0}{\bm{1}}+\beta_{0}\mathbf{r})^{\top}({\bm{\Sigma}}/(1+s_{0})+\mathbf{Q})^{-1}{\bm{\Sigma}}({\bm{\Sigma}}/(1+s_{0})+\mathbf{Q})^{-1}(\alpha_{0}{\bm{1}}+\beta_{0}\mathbf{r})}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1,$ (E.8)
and
$\frac{(1+s_{0})^{2}}{1+s_{0}+s_{1,\mathbf{Q}}}\cdot\frac{(\alpha_{n}{\bm{1}}+\beta_{n}\mathbf{r})^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\widehat{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}(\alpha_{n}{\bm{1}}+\beta_{n}\mathbf{r})}{(\alpha_{0}{\bm{1}}+\beta_{0}\mathbf{r})^{\top}({\bm{\Sigma}}/(1+s_{0})+\mathbf{Q})^{-1}{\bm{\Sigma}}({\bm{\Sigma}}/(1+s_{0})+\mathbf{Q})^{-1}(\alpha_{0}{\bm{1}}+\beta_{0}\mathbf{r})}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1.$ (E.9)
Proof of (E.8). To prove (E.8), we first give the order of the dominator $(\alpha_{0}{\bm{1}}+\beta_{0}\mathbf{r})^{\top}({\bm{\Sigma}}/(1+s_{0})+\mathbf{Q})^{-1}{\bm{\Sigma}}({\bm{\Sigma}}/(1+s_{0})+\mathbf{Q})^{-1}(\alpha_{0}{\bm{1}}+\beta_{0}\mathbf{r})$. From Lemma E.2, it holds that
$(\alpha_{0}{\bm{1}}+\beta_{0}\mathbf{r})^{\top}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}(\alpha_{0}{\bm{1}}+\beta_{0}\mathbf{r})$
$\quad=\Theta\bigg{(}(\alpha_{0}{\bm{1}}+\beta_{0}\mathbf{r})^{\top}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}(\alpha_{0}{\bm{1}}+\beta_{0}\mathbf{r})\bigg{)}.$ (E.10)
Recall the definition of $\alpha_{0}$ and $\beta_{0}$ in (E.7), we have
$(\alpha_{0}{\bm{1}}+\beta_{0}\mathbf{r})^{\top}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}(\alpha_{0}{\bm{1}}+\beta_{0}\mathbf{r})=(\mathcal{A}_{11}\mathcal{A}_{rr}-\mathcal{A}_{r1}^{2})\cdot\big{(}\mu_{0}(\mathcal{A}_{11}\mu_{0}-2\mathcal{A}_{r1})+\mathcal{A}_{rr}\big{)}.$ (E.11)
By Assumption 3.1, we can easily get that $\mathcal{A}_{11}\mathcal{A}_{rr}-\mathcal{A}_{r1}^{2}=\Theta(\mathcal{A}_{11}\mathcal{A}_{rr})$. We separate the proof of (E.8) into two cases: $\mu_{0}\leq C\sqrt{\mathcal{A}_{rr}/\mathcal{A}_{11}}$ and $\mu_{0}\geq C\sqrt{\mathcal{A}_{rr}/\mathcal{A}_{11}}$ for some sufficiently large constant $C>0$. Before the proof of (E.8) and (E.9), we define several terms.
$\begin{split}\mathcal{J}_{0}:&=(\alpha_{0}{\bm{1}}+\beta_{0}\mathbf{r})^{\top}({\bm{\Sigma}}/(1+s_{0})+\mathbf{Q})^{-1}{\bm{\Sigma}}({\bm{\Sigma}}/(1+s_{0})+\mathbf{Q})^{-1}(\alpha_{0}{\bm{1}}+\beta_{0}\mathbf{r}),\\ \mathcal{J}_{1}:&=(\alpha_{n}{\bm{1}}+\beta_{n}\mathbf{r})^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}(\alpha_{n}{\bm{1}}+\beta_{n}\mathbf{r}),\\ \mathcal{J}_{2}:&=(\alpha_{0}{\bm{1}}+\beta_{0}\mathbf{r})^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}(\alpha_{0}{\bm{1}}+\beta_{0}\mathbf{r}).\end{split}$ (E.12)
It is clear that $\mathcal{J}_{1}/D^{2}=\sigma_{0}^{2}$. Also from (E.10) and (E.11), we have
$\mathcal{J}_{0}=\Theta(\mathcal{A}_{11}\mathcal{A}_{rr})\cdot(\mu_{0}(\mathcal{A}_{11}\mu_{0}-2\mathcal{A}_{r1})+\mathcal{A}_{rr}).$ (E.13)
Case 1: If $\mu_{0}\leq C\sqrt{\mathcal{A}_{rr}/\mathcal{A}_{11}}$ holds, we have
$\mu_{0}(\mathcal{A}_{11}\mu_{0}-2\mathcal{A}_{r1})+\mathcal{A}_{rr}\leq C^{2}(\mathcal{A}_{rr}+\mathcal{A}_{r1}\cdot\sqrt{\mathcal{A}_{rr}/\mathcal{A}_{11}})=O(\mathcal{A}_{rr}).$
Here, the last equality is by $\mathcal{A}_{r1}^{2}/(\mathcal{A}_{rr}\cdot\mathcal{A}_{11})\leq 1$. For the minimum value of $\mu_{0}(\mathcal{A}_{11}\mu_{0}-2\mathcal{A}_{r1})+\mathcal{A}_{rr}$, it takes value when $\mu_{0}=\mathcal{A}_{r1}/\mathcal{A}_{11}$, therefore we have
$\mu_{0}(\mathcal{A}_{11}\mu_{0}-2\mathcal{A}_{r1})+\mathcal{A}_{rr}\geq\frac{\mathcal{A}_{11}\mathcal{A}_{rr}-\mathcal{A}_{r1}^{2}}{\mathcal{A}_{11}}\geq\Omega(\mathcal{A}_{rr}).$
Hence we conclude that when $\mu_{0}=O(\sqrt{\mathcal{A}_{rr}/\mathcal{A}_{11}})$, $\mu_{0}(\mathcal{A}_{11}\mu_{0}-2\mathcal{A}_{r1})+\mathcal{A}_{rr}=\Theta(\mathcal{A}_{rr})$. We have $\big{(}\mu_{0}(\mathcal{A}_{11}\mu_{0}-2\mathcal{A}_{r1})+\mathcal{A}_{rr}\big{)}=\Theta(\mathcal{A}_{rr})$. Combined this with (E.12) and (E.13) we have
$\frac{\mathcal{J}_{0}}{\mathcal{A}_{rr}^{2}\mathcal{A}_{11}}=\Theta(1).$ (E.14)
We claim that the following facts hold.
$1.~{}\frac{(1+s_{0})^{2}}{(1+s_{0})^{2}-s_{1,{\bm{\Sigma}}}}=\Theta(1),\quad 2.~{}\frac{(1+s_{0})^{2}}{(1+s_{0})^{2}-s_{1,{\bm{\Sigma}}}}\cdot\frac{\mathcal{J}_{2}}{\mathcal{J}_{0}}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1,$
$0\displaystyle 3.~{}\frac{\mathcal{J}_{0}}{\mathcal{A}_{rr}^{2}\mathcal{A}_{11}}=\Theta(1),\quad 4.~{}\frac{\mathcal{J}_{1}-\mathcal{J}_{2}}{\mathcal{A}_{rr}^{2}\mathcal{A}_{11}}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}$ 0 .
Then Equation (E.8), which is equivalent to $\frac{(1+s_{0})^{2}}{(1+s_{0})^{2}-s_{1,{\bm{\Sigma}}}}\cdot\mathcal{J}_{1}/\mathcal{J}_{0}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1$ can be directly concluded from the above facts. Moreover, by $\mathcal{J}_{1}/D^{2}=\sigma_{0}^{2}$ and $D=\Theta(\mathcal{A}_{11}\mathcal{A}_{rr})$ in Lemma E.3, we have
$\sigma_{0}^{2}=\Theta(1/\mathcal{A}_{11}).$
It is easy to verify that the first fact comes from Lemma C.5, the second fact comes from Lemma E.1 and the third fact comes from (E.14). Therefore, to prove (E.8), it only remains to prove the last fact, that is $(\mathcal{J}_{1}-\mathcal{J}_{2})/\mathcal{A}_{rr}^{2}\mathcal{A}_{11}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0$. By Lemma E.3, it is not hard to see that
$\frac{\alpha_{n}}{\mathcal{A}_{rr}}-\frac{\alpha_{0}}{\mathcal{A}_{rr}}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0,\quad\frac{\beta_{n}}{\sqrt{\mathcal{A}_{rr}\cdot\mathcal{A}_{11}}}-\frac{\beta_{0}}{\sqrt{\mathcal{A}_{rr}\cdot\mathcal{A}_{11}}}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0.$
We apply perturbation analysis to prove $(\mathcal{J}_{1}-\mathcal{J}_{2})/\mathcal{A}_{rr}^{2}\mathcal{A}_{11}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0$. Define $\mathcal{J}^{\circ}=(\alpha_{0}{\bm{1}}+\beta_{n}\mathbf{r})^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}(\alpha_{0}{\bm{1}}+\beta_{n}\mathbf{r})$, $\mathcal{J}_{1}-\mathcal{J}_{2}$ can be separated into $\mathcal{J}_{1}-\mathcal{J}^{\circ}$ and $\mathcal{J}^{\circ}-\mathcal{J}_{2}$. We only need to prove that the two terms $(\mathcal{J}_{1}-\mathcal{J}^{\circ})/(\mathcal{A}_{rr}^{2}\mathcal{A}_{11})$ and $(\mathcal{J}^{\circ}-\mathcal{J}_{2})/(\mathcal{A}_{rr}^{2}\mathcal{A}_{11})$ both converge to 00. We need the following equations. Similar to the proof of Lemma E.2, we have
${\bm{1}}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{1}}=\Theta(\mathcal{A}_{11}),$
$\big{|}\mathbf{r}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{1}}\big{|}=O(|\mathcal{A}_{r1}|),$
$\mathbf{r}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\mathbf{r}=\Theta(\mathcal{A}_{rr}),$
$\frac{|\alpha_{0}|}{\mathcal{A}_{rr}}=\frac{|\mathcal{A}_{rr}-\mu_{0}\mathcal{A}_{r1}|}{\mathcal{A}_{rr}}\leq 1+\mu_{0}\frac{|\mathcal{A}_{r1}|}{\mathcal{A}_{rr}}\leq 1+C,$
$\frac{|\beta_{0}|}{\sqrt{\mathcal{A}_{rr}\cdot\mathcal{A}_{11}}}=\frac{|\mu_{0}\mathcal{A}_{11}-\mathcal{A}_{r1}|}{\sqrt{\mathcal{A}_{rr}\cdot\mathcal{A}_{11}}}\leq 1+\mu_{0}\frac{|\mathcal{A}_{11}|}{\sqrt{\mathcal{A}_{rr}\cdot\mathcal{A}_{11}}}\leq 1+C.$
Combining the conclusions above with Lemma E.3, we can see
$\frac{\mathcal{J}_{1}-\mathcal{J}^{\circ}}{\mathcal{A}_{rr}^{2}\mathcal{A}_{11}}$ $=O\bigg{(}\frac{\alpha_{n}^{2}-\alpha_{0}^{2}}{\mathcal{A}_{rr}^{2}}\bigg{)}+O\bigg{(}\frac{\alpha_{n}-\alpha_{0}}{\mathcal{A}_{rr}}\bigg{)}=o(1),$
$\frac{\mathcal{J}^{\circ}-\mathcal{J}_{2}}{\mathcal{A}_{rr}^{2}\mathcal{A}_{11}}$ $=O\bigg{(}\frac{\beta_{n}^{2}-\beta_{0}^{2}}{\mathcal{A}_{rr}\mathcal{A}_{11}}\bigg{)}+O\bigg{(}\frac{\beta_{n}-\beta_{0}}{\sqrt{\mathcal{A}_{rr}\mathcal{A}_{11}}}\bigg{)}=o(1).$
Here, the convergence also utilize the fact that $|\alpha_{n}+\alpha_{0}|/\mathcal{A}_{rr}\leq 3(1+C)$, $|\beta_{n}+\beta_{0}|/\sqrt{\mathcal{A}_{rr}\cdot\mathcal{A}_{11}}\leq 3(1+C)$. We conclude that $\mathcal{J}_{1}-\mathcal{J}_{2}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0$, hence we have $\frac{(1+s_{0})^{2}}{(1+s_{0})^{2}-s_{1,{\bm{\Sigma}}}}\cdot\frac{\mathcal{J}_{1}}{\mathcal{J}_{0}}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1$ which completes the proof of (E.8). Similarly we can prove (E.9). We hence complete the proof of (E.8) when $\mu_{0}\leq C\sqrt{\mathcal{A}_{rr}/\mathcal{A}_{11}}$.
Case 2: If $\mu_{0}\geq C\sqrt{\mathcal{A}_{rr}/\mathcal{A}_{11}}$ holds, we have
$\mathcal{A}_{11}\mu_{0}-2\mathcal{A}_{r1}$ $=\mathcal{A}_{11}\mu_{0}/2+\mathcal{A}_{11}\mu_{0}/2-2\mathcal{A}_{r1}$
$\geq\mathcal{A}_{11}\mu_{0}/2+\frac{C}{2}\sqrt{\mathcal{A}_{11}\mathcal{A}_{rr}}-\sqrt{\mathcal{A}_{11}\mathcal{A}_{rr}}\geq\mathcal{A}_{11}\mu_{0}/2,$
where the first inequality is by condition $\mu_{0}\geq C\sqrt{\mathcal{A}_{rr}/\mathcal{A}_{11}}$ and the second inequality is by $C$ large enough. Similarly, we can conclude that $\mathcal{A}_{11}\mu_{0}-2\mathcal{A}_{r1}\leq 2\mu_{0}\mathcal{A}_{11}$. Hence it holds that
$\mathcal{A}_{11}\mu_{0}^{2}/2\leq\mu_{0}(\mathcal{A}_{11}\mu_{0}-2\mathcal{A}_{r1})+\mathcal{A}_{rr}\leq 3\mathcal{A}_{11}\mu_{0}^{2}.$
Here, we use the fact $\mathcal{A}_{rr}\leq\mathcal{A}_{11}\mu_{0}^{2}$. Combined the inequalities above with (E.11), we have
$(\alpha_{0}{\bm{1}}+\beta_{0}\mathbf{r})^{\top}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}(\alpha_{0}{\bm{1}}+\beta_{0}\mathbf{r})=\Theta\big{(}\mu_{0}^{2}\mathcal{A}_{rr}\mathcal{A}_{11}^{2}\big{)}.$ (E.15)
Combined this with (E.12) and (E.13) we have
$\frac{\mathcal{J}_{0}}{\mu_{0}^{2}\mathcal{A}_{rr}\mathcal{A}_{11}^{2}}=\Theta(1).$ (E.16)
Here, the second equality uses (E.10) and (E.15). Similar to the case when $\mu_{0}\leq C\sqrt{\mathcal{A}_{rr}/\mathcal{A}_{11}}$, we claim that the following facts hold.
$1.~{}\frac{(1+s_{0})^{2}}{(1+s_{0})^{2}-s_{1,{\bm{\Sigma}}}}=\Theta(1),\quad 2.~{}\frac{(1+s_{0})^{2}}{(1+s_{0})^{2}-s_{1,{\bm{\Sigma}}}}\cdot\frac{\mathcal{J}_{2}}{\mathcal{J}_{0}}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1,$
$0\displaystyle 3.~{}\frac{\mathcal{J}_{0}}{\mu_{0}^{2}\mathcal{A}_{rr}\mathcal{A}_{11}^{2}}=\Theta(1),\quad 4.~{}\frac{\mathcal{J}_{1}-\mathcal{J}_{2}}{\mu_{0}^{2}\mathcal{A}_{rr}\mathcal{A}_{11}^{2}}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}$ 0 .
Then Equation (E.8), which is equivalent to $\frac{(1+s_{0})^{2}}{(1+s_{0})^{2}-s_{1,{\bm{\Sigma}}}}\cdot\mathcal{J}_{1}/\mathcal{J}_{0}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1$, can be directly concluded from the above facts. Moreover, by $\mathcal{J}_{1}/D^{2}=\sigma_{0}^{2}$ and $D=\Theta(\mathcal{A}_{11}\mathcal{A}_{rr})$ in Lemma E.3, we have
$\sigma_{0}^{2}=\Theta(\mu_{0}^{2}/\mathcal{A}_{rr}).$
It is easy to verify that the first fact comes from Lemma C.5, the second fact comes from Lemma E.1 and the third fact comes from (E.16). Therefore, to prove (E.8), it only remains to prove that $(\mathcal{J}_{1}-\mathcal{J}_{2})/(\mu_{0}^{2}\mathcal{A}_{rr}\mathcal{A}_{11}^{2})~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0$. By Lemma E.3, it is easy to see that
$\frac{\alpha_{n}}{\mu_{0}\sqrt{\mathcal{A}_{rr}\mathcal{A}_{11}}}-\frac{\alpha_{0}}{\mu_{0}\sqrt{\mathcal{A}_{rr}\mathcal{A}_{11}}}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0,\quad\frac{\beta_{n}}{\mu_{0}\mathcal{A}_{11}}-\frac{\beta_{0}}{\mu_{0}\mathcal{A}_{11}}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0.$
We apply perturbation analysis to prove $(\mathcal{J}_{1}-\mathcal{J}_{2})/(\mu_{0}^{2}\mathcal{A}_{rr}\mathcal{A}_{11}^{2})~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0$. Define $\mathcal{J}^{\circ}=(\alpha_{0}{\bm{1}}+\beta_{n}\mathbf{r})^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}(\alpha_{0}{\bm{1}}+\beta_{n}\mathbf{r})$, $\mathcal{J}_{1}-\mathcal{J}_{2}$ can be separated into $\mathcal{J}_{1}-\mathcal{J}^{\circ}$ and $\mathcal{J}^{\circ}-\mathcal{J}_{2}$. We only need to prove that the two terms $(\mathcal{J}_{1}-\mathcal{J}^{\circ})/(\mu_{0}^{2}\mathcal{A}_{rr}\mathcal{A}_{11}^{2})$ and $(\mathcal{J}^{\circ}-\mathcal{J}_{2})/(\mu_{0}^{2}\mathcal{A}_{rr}\mathcal{A}_{11}^{2})$ both converge to 00. We have the following equations.
${\bm{1}}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{1}}=\Theta(\mathcal{A}_{11}),$
$\big{|}\mathbf{r}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{1}}\big{|}=O(|\mathcal{A}_{r1}|),$
$\mathbf{r}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\mathbf{r}=\Theta(\mathcal{A}_{rr}),$
$\frac{|\alpha_{0}|}{\mu_{0}\sqrt{\mathcal{A}_{rr}\mathcal{A}_{11}}}=\frac{|\mathcal{A}_{rr}-\mu_{0}\mathcal{A}_{r1}|}{\mu_{0}\sqrt{\mathcal{A}_{rr}\mathcal{A}_{11}}}\leq\frac{1}{C}+\mu_{0}\frac{|\mathcal{A}_{r1}|}{\mu_{0}\sqrt{\mathcal{A}_{rr}\mathcal{A}_{11}}}\leq\frac{1}{C}+1,$
$\frac{|\beta_{0}|}{\mu_{0}\mathcal{A}_{11}}=\frac{|\mu_{0}\mathcal{A}_{11}-\mathcal{A}_{r1}|}{\mu_{0}\mathcal{A}_{11}}\leq\frac{1}{C}+\mu_{0}\frac{|\mathcal{A}_{11}|}{\mu_{0}\mathcal{A}_{11}}\leq\frac{1}{C}+1.$
Combined the conclusions above with Lemma E.3, we can see
$\frac{\mathcal{J}_{1}-\mathcal{J}^{\circ}}{\mu_{0}^{2}\mathcal{A}_{rr}\mathcal{A}_{11}^{2}}$ $=O\bigg{(}\frac{\alpha_{n}^{2}-\alpha_{0}^{2}}{\mu_{0}^{2}\mathcal{A}_{rr}\mathcal{A}_{11}}\bigg{)}+O\bigg{(}\frac{\alpha_{n}-\alpha_{0}}{\mu_{0}\sqrt{\mathcal{A}_{rr}\mathcal{A}_{11}}}\bigg{)}=o(1),$
$\frac{\mathcal{J}^{\circ}-\mathcal{J}_{2}}{\mu_{0}^{2}\mathcal{A}_{rr}\mathcal{A}_{11}^{2}}$ $=O\bigg{(}\frac{\beta_{n}^{2}-\beta_{0}^{2}}{\mu_{0}^{2}\mathcal{A}_{11}^{2}}\bigg{)}+O\bigg{(}\frac{\beta_{n}-\beta_{0}}{\mu_{0}\mathcal{A}_{11}}\bigg{)}=o(1).$
We conclude that $(\mathcal{J}_{1}-\mathcal{J}_{2})/(\mu_{0}^{2}\mathcal{A}_{rr}\mathcal{A}_{11}^{2})~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0$, hence we have $\frac{(1+s_{0})^{2}}{(1+s_{0})^{2}-s_{1,{\bm{\Sigma}}}}\cdot\frac{\mathcal{J}_{1}}{\mathcal{J}_{0}}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1$, which completes the proof of (E.8). We hence complete the proof of (E.8) when $\mu_{0}\geq C\sqrt{\mathcal{A}_{rr}/\mathcal{A}_{11}}$.
Proof of (E.9). We can prove (E.9) similar to the proof of (E.8), with a few small differences. We need to prove that the constant $(1+s_{0})^{2}/(1+s_{0}+s_{1,\mathbf{Q}})=\Theta(1)$, which can be easily concluded from Lemma C.5, and we simply replace the matrix ${\bm{\Sigma}}$ by $\widehat{\bm{\Sigma}}$ in the expressions of $\mathcal{J}_{1}$ and $\mathcal{J}_{2}$.
Proof of ratio consistency. From (E.8) and (E.9), we can conclude
$\frac{1+s_{0}+s_{1,\mathbf{Q}}}{(1+s_{0})^{2}-s_{1,{\bm{\Sigma}}}}\cdot\frac{(\alpha_{n}{\bm{1}}+\beta_{n}\mathbf{r})^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}(\alpha_{n}{\bm{1}}+\beta_{n}\mathbf{r})}{(\alpha_{n}{\bm{1}}+\beta_{n}\mathbf{r})^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\widehat{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}(\alpha_{n}{\bm{1}}+\beta_{n}\mathbf{r})}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1.$ (E.17)
Here the convergence comes from (E.8)/(E.9). Moreover, Proposition D.4 indicates that
$\frac{1+s_{0}+s_{1,\mathbf{Q}}}{(1+s_{0})^{2}-s_{1,{\bm{\Sigma}}}}/\bigg{(}1-\frac{c}{p}{\mathrm{tr}}\widehat{\bm{\Sigma}}\big{(}\widehat{\bm{\Sigma}}+\mathbf{Q}\big{)}^{-1}\bigg{)}^{2}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}1.$
Combined this with (E.17) we have
$\frac{\sigma_{0}^{2}}{\widehat{\sigma}^{2}}=\bigg{(}1-\frac{c}{p}{\mathrm{tr}}\widehat{\bm{\Sigma}}\big{(}\widehat{\bm{\Sigma}}+\mathbf{Q}\big{)}^{-1}\bigg{)}^{2}\cdot\frac{(\alpha_{n}{\bm{1}}+\beta_{n}\mathbf{r})^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}(\alpha_{n}{\bm{1}}+\beta_{n}\mathbf{r})}{(\alpha_{n}{\bm{1}}+\beta_{n}\mathbf{r})^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\widehat{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}(\alpha_{n}{\bm{1}}+\beta_{n}\mathbf{r})}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1$
This completes the proof of ratio consistency.
Proof of difference consistency. As for the remaining conclusions on the consistency of Sharpe ratio or variance estimation, from the analysis above we have
  1. 1.
    $\sigma_{0}^{2}=\Theta(1/\mathcal{A}_{11})$ when $\mu_{0}\leq C\sqrt{\mathcal{A}_{rr}/\mathcal{A}_{11}}$;
  2. 2.
    $\sigma_{0}^{2}=\Theta(\mu_{0}^{2}/{\mathcal{A}_{rr}})$ when $\mu_{0}\geq C\sqrt{\mathcal{A}_{rr}/\mathcal{A}_{11}}$.
When $\mathcal{A}_{rr}$ is bounded, if $\mu_{0}\leq C\sqrt{\mathcal{A}_{rr}/\mathcal{A}_{11}}$ we can see that $\mu_{0}/\sigma_{0}=\Theta(\mu_{0}\cdot\sqrt{\mathcal{A}_{11}})=O(\sqrt{\mathcal{A}_{rr}})=O(1)$; if $\mu_{0}\geq C\sqrt{\mathcal{A}_{rr}/\mathcal{A}_{11}}$, we can see that $\mu_{0}/\sigma_{0}=\Theta(\mu_{0}/(\mu_{0}/\sqrt{\mathcal{A}_{rr}}))=O(1)$. Therefore by ratio consistency we have $\frac{\mu_{0}}{\sigma_{0}}-\frac{\mu_{0}}{\widehat{\sigma}}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0$. When $r_{0}=O(\mu_{0})$, we also have $\frac{r_{0}}{\sigma_{0}}-\frac{r_{0}}{\widehat{\sigma}}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0$, this completes the proof of the consistency of Sharpe ratio.
To prove $\sigma_{0}^{2}-\widehat{\sigma}^{2}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0$. By $\|{\bm{\Sigma}}/p\|_{{\mathrm{tr}}}\leq C$, we have $\lambda_{\min}\big{(}{\bm{\Sigma}}/(1+s_{0})+\mathbf{Q}\big{)}^{-1}\geq\Theta(1/p)$, which indicates $\mathcal{A}_{11}\geq\Theta(1)$. Therefore, if $\mu_{0}\leq C\sqrt{\mathcal{A}_{rr}/\mathcal{A}_{11}}$ we have $\sigma_{0}^{2}=\Theta(1/\mathcal{A}_{11})=O(1)$; if $\mu_{0}\geq C\sqrt{\mathcal{A}_{rr}/\mathcal{A}_{11}}$ we have $\sigma_{0}^{2}=\Theta(\mu_{0}^{2}/{\mathcal{A}_{rr}})$. The condition $\mu_{0}\leq C\sqrt{\mathcal{A}_{rr}}$ then indicates that $\sigma_{0}^{2}=O(1)$. By ratio consistency we have $\sigma_{0}^{2}-\widehat{\sigma}^{2}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0$. We complete the proof of Theorem 3.2. ∎
E.1 Proof of Lemma E.2
Note that $\mathbf{A}$ is semi-positive defined. It is equal for us to prove that for any sequence of ${\bm{\xi}}\in\mathbb{R}^{p}$,
${\bm{\xi}}^{\top}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}{\bm{\xi}}=\Theta\bigg{(}{\bm{\xi}}^{\top}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}{\bm{\xi}}\bigg{)}.$
We have
${\bm{\xi}}^{\top}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}{\bm{\xi}}$
$\quad={\bm{\xi}}^{\top}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-\frac{1}{2}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-\frac{1}{2}}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-\frac{1}{2}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-\frac{1}{2}}{\bm{\xi}}.$
The first conclusion in Lemma E.2 holds from the fact that
$\lambda_{\max}\bigg{(}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-\frac{1}{2}}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-\frac{1}{2}}\bigg{)}\leq 1+s_{0},$
$\lambda_{\min}\bigg{(}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-\frac{1}{2}}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-\frac{1}{2}}\bigg{)}$
$\qquad=\lambda_{\min}\bigg{(}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\bigg{)}^{-1}{\bm{\Sigma}}\bigg{)}\geq\frac{1+s_{0}}{C_{\mathbf{Q}}(1+s_{0})+1}.$
To prove ${\bm{\xi}}^{\top}\big{(}{\bm{\Sigma}}/(1+s_{0})+\mathbf{Q}\big{)}^{-1}{\bm{\xi}}=\Theta(\|{\bm{\Sigma}}^{-\frac{1}{2}}{\bm{\xi}}\|_{2}^{2})$, we have
${\bm{\xi}}^{\top}\big{(}{\bm{\Sigma}}/(1+s_{0})+\mathbf{Q}\big{)}^{-1}{\bm{\xi}}=({\bm{\Sigma}}^{-\frac{1}{2}}{\bm{\xi}})^{\top}\big{(}\mathbf{I}/(1+s_{0})+{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{Q}{\bm{\Sigma}}^{-\frac{1}{2}}\big{)}^{-1}{\bm{\Sigma}}^{-\frac{1}{2}}{\bm{\xi}},$
the conclusion directly holds from the fact that the maximum and minimum eigenvalue of $\big{(}\mathbf{I}/(1+s_{0})+{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{Q}{\bm{\Sigma}}^{-\frac{1}{2}}\big{)}^{-1}$ can be upper and lower bounded by a constant, respectively. This completes the proof of Lemma E.2.
E.2 Proof of Lemma E.3
We first prove for the case of $\mu_{0}\leq C\sqrt{\mathcal{A}_{rr}/\mathcal{A}_{11}}$. From the results of Proposition D.2, we have that
$\frac{\mathbf{r}^{\top}\big{(}\widehat{\bm{\Sigma}}+\mathbf{Q}\big{)}^{-1}\mathbf{r}}{\mathcal{A}_{rr}}-1~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0.$
For the term $\mu_{0}\mathbf{r}^{\top}\big{(}\widehat{\bm{\Sigma}}+\mathbf{Q}\big{)}^{-1}{\bm{1}}/\mathcal{A}_{rr}$, we have
$\frac{\mu_{0}\mathbf{r}^{\top}\big{(}\widehat{\bm{\Sigma}}+\mathbf{Q}\big{)}^{-1}{\bm{1}}}{\mathcal{A}_{rr}}=\frac{\mu_{0}({\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{r})^{\top}\big{(}\mathbf{Z}^{\top}\mathbf{Z}/n+{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{Q}{\bm{\Sigma}}^{-\frac{1}{2}}\big{)}^{-1}{\bm{\Sigma}}^{-\frac{1}{2}}{\bm{1}}}{\mathcal{A}_{rr}}.$
We next prove that
$\mu_{0}\cdot\bigg{\|}\frac{{\bm{\Sigma}}^{-\frac{1}{2}}{\bm{1}}({\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{r})^{\top}}{\mathcal{A}_{rr}}\bigg{\|}_{{\mathrm{tr}}}$
is bounded. If the results above hold, by applying the results in Proposition D.2 we will have
$\frac{\mu_{0}({\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{r})^{\top}\big{(}\mathbf{Z}^{\top}\mathbf{Z}/n+{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{Q}{\bm{\Sigma}}^{-\frac{1}{2}}\big{)}^{-1}{\bm{\Sigma}}^{-\frac{1}{2}}{\bm{1}}}{\mathcal{A}_{rr}}-\frac{\mu_{0}({\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{r})^{\top}\big{(}\mathbf{I}/(1+s_{0})+{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{Q}{\bm{\Sigma}}^{-\frac{1}{2}}\big{)}^{-1}{\bm{\Sigma}}^{-\frac{1}{2}}{\bm{1}}}{\mathcal{A}_{rr}}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0,$
which directly implies
$\frac{\mu_{0}\mathbf{r}^{\top}\big{(}\widehat{\bm{\Sigma}}+\mathbf{Q}\big{)}^{-1}{\bm{1}}}{\mathcal{A}_{rr}}-\frac{\mu_{0}\mathbf{r}^{\top}\big{(}{\bm{\Sigma}}/(1+s_{0})+\mathbf{Q}\big{)}^{-1}{\bm{1}}}{\mathcal{A}_{rr}}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0.$
Then the first conclusion in Lemma E.3 will hold.
It remains only to prove that $\mu_{0}\cdot\Big{\|}\frac{{\bm{\Sigma}}^{-\frac{1}{2}}{\bm{1}}({\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{r})^{\top}}{\mathcal{A}_{rr}}\Big{\|}_{{\mathrm{tr}}}$ is bounded. Note that $\mu_{0}=O(\sqrt{\mathcal{A}_{rr}/\mathcal{A}_{11}})$, therefore it is equal to prove that
$\bigg{\|}\frac{{\bm{\Sigma}}^{-\frac{1}{2}}{\bm{1}}({\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{r})^{\top}}{\sqrt{\mathcal{A}_{rr}\cdot\mathcal{A}_{11}}}\bigg{\|}_{{\mathrm{tr}}}$
is bounded. From Lemma E.2, we can see that $\sqrt{\mathcal{A}_{rr}\cdot\mathcal{A}_{11}}=\Theta(\|{\bm{\Sigma}}^{-\frac{1}{2}}{\bm{1}}\|_{2}\cdot\|{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{r}\|_{2})$, we conclude that
$\mu_{0}\cdot\bigg{\|}\frac{{\bm{\Sigma}}^{-\frac{1}{2}}{\bm{1}}({\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{r})^{\top}}{\mathcal{A}_{rr}}\bigg{\|}_{{\mathrm{tr}}}$
is bounded, We complete the proof of the first conclusion when $\mu_{0}\leq C\sqrt{\mathcal{A}_{rr}/\mathcal{A}_{11}}$. The proof of the second conclusion when $\mu_{0}\leq C\sqrt{\mathcal{A}_{rr}/\mathcal{A}_{11}}$ follows the same procedure and thus we omit it.
We then prove the case of $\mu_{0}\geq C\sqrt{\mathcal{A}_{rr}/\mathcal{A}_{11}}$. For the first equation, we have
$\frac{\mathbf{r}^{\top}\big{(}\widehat{\bm{\Sigma}}+\mathbf{Q}\big{)}^{-1}\mathbf{r}}{\mathcal{A}_{rr}}-\frac{\mathcal{A}_{rr}}{\mathcal{A}_{rr}}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0.$
Note that $\mu_{0}\sqrt{\mathcal{A}_{11}/\mathcal{A}_{rr}}\geq C>0$, we divide $\mu_{0}\sqrt{\mathcal{A}_{11}/\mathcal{A}_{rr}}$ into the equation above and get that
$\frac{\mathbf{r}^{\top}\big{(}\widehat{\bm{\Sigma}}+\mathbf{Q}\big{)}^{-1}\mathbf{r}}{\mu_{0}\sqrt{\mathcal{A}_{rr}\mathcal{A}_{11}}}-\frac{\mathcal{A}_{rr}}{\mu_{0}\sqrt{\mathcal{A}_{rr}\mathcal{A}_{11}}}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0.$ (E.18)
We also know that
$\frac{\mathbf{r}^{\top}\big{(}\widehat{\bm{\Sigma}}+\mathbf{Q}\big{)}^{-1}{\bm{1}}}{\sqrt{\mathcal{A}_{rr}\mathcal{A}_{11}}}-\frac{\mathcal{A}_{r1}}{\sqrt{\mathcal{A}_{rr}\mathcal{A}_{11}}}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0$ (E.19)
since $\Big{\|}\frac{{\bm{\Sigma}}^{-\frac{1}{2}}{\bm{1}}({\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{r})^{\top}}{\sqrt{\mathcal{A}_{rr}\cdot\mathcal{A}_{11}}}\Big{\|}_{{\mathrm{tr}}}$ is bounded, therefore (E.18) and (E.19) show that
$\frac{\alpha_{n}-\alpha_{0}}{\mu_{0}\sqrt{\mathcal{A}_{rr}\mathcal{A}_{11}}}=\frac{\mathbf{r}^{\top}\big{(}\widehat{\bm{\Sigma}}+\mathbf{Q}\big{)}^{-1}\mathbf{r}-\mu_{0}\mathbf{r}^{\top}\big{(}\widehat{\bm{\Sigma}}+\mathbf{Q}\big{)}^{-1}{\bm{1}}}{\mu_{0}\sqrt{\mathcal{A}_{rr}\mathcal{A}_{11}}}-\frac{\mathcal{A}_{rr}-\mu_{0}\mathcal{A}_{r1}}{\mu_{0}\sqrt{\mathcal{A}_{rr}\mathcal{A}_{11}}}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0.$
We can similarly show that
$\frac{\mu_{0}{\bm{1}}^{\top}\big{(}\widehat{\bm{\Sigma}}+\mathbf{Q}\big{)}^{-1}{\bm{1}}-\mathbf{r}^{\top}\big{(}\widehat{\bm{\Sigma}}+\mathbf{Q}\big{)}^{-1}{\bm{1}}}{\mu_{0}\mathcal{A}_{11}}-\frac{\mu_{0}\mathcal{A}_{11}-\mathcal{A}_{r1}}{\mu_{0}\mathcal{A}_{11}}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0.$
For the order of $D=BC-A^{2}$, it is clear from (E.19) that
$\frac{A^{2}-\mathcal{A}_{r1}}{\mathcal{A}_{rr}\mathcal{A}_{11}}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0,\quad\frac{BC}{\mathcal{A}_{rr}\mathcal{A}_{11}}-1~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0.$
Therefore we have $(BC-A^{2})/(\mathcal{A}_{rr}\mathcal{A}_{11})-(\mathcal{A}_{rr}\mathcal{A}_{11}-\mathcal{A}_{r1}^{2})/(\mathcal{A}_{rr}\mathcal{A}_{11})~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0$. By Assumption 3.1 we have $(\mathcal{A}_{rr}\mathcal{A}_{11}-\mathcal{A}_{r1}^{2})/(\mathcal{A}_{rr}\mathcal{A}_{11})=\Theta(1)$, therefore $D/(\mathcal{A}_{rr}\mathcal{A}_{11})=\Theta(1)$. This completes the proof of Lemma E.3.
Appendix F Relationship Between Ridge Regularization and Maximum Sharpe Ratio
According to the theory proposed by Markowitz (1952), the maximum population Sharpe ratio can be expressed as:
$SR_{\max}=\sqrt{{\bm{\mu}}^{\top}{\bm{\Sigma}}^{-1}{\bm{\mu}}}.$
In this section, we discuss the relationship between $SR_{\max}$ and the optimal Sharpe ratio $SR(\mathbf{Q})$ we can achieve by choosing $\mathbf{Q}$. The following proposition demonstrates that there exists $\widetilde{\mathbf{Q}}$ such that $SR(\widetilde{\mathbf{Q}})$ can approximate $SR_{\max}$ very closely.
Proposition F.1.
Suppose that Assumptions 2.2, 2.4, 2.5 and 2.6 hold, and $\|{\bm{\Sigma}}^{-1}\|_{{\mathop{\mathrm{op}}}}$ is bounded. Then for any given $\varepsilon>0$, there exists deterministic sequences of matrices $\widetilde{\mathbf{Q}}\in\mathbb{R}^{p\times p}$ such that with probability 1,
$1-\varepsilon\leq\lim_{n\to+\infty}SR(\widetilde{\mathbf{Q}})/SR_{\max}\leq 1$
The proof of Proposition F.1 is given in Section F.1. This proposition tells us that a well-designed regularization matrix $\widetilde{\mathbf{Q}}$ allows $SR(\widetilde{\mathbf{Q}})$ to closely approximate the maximum population Sharpe ratio. Furthermore, we could use the criteria established in Theorem 2.7 to select $\mathbf{Q}$ that approximately maximizes $SR(\mathbf{Q})$.
Another interesting case arises when ${\bm{\Sigma}}$ exhibits a specific factor structure, defined as ${\bm{\Sigma}}=\mathbf{B}\mathbf{B}^{\top}+\mathbf{D}$. Here, $\mathbf{B}=(\bm{b}_{1},\dots,\bm{b}_{K})\in\mathbb{R}^{p\times K}$ with each $\|\bm{b}_{j}\|_{2}=\Theta(\sqrt{p})$ for $j\in[K]$, $\operatorname{rank}(\mathbf{B})=K$ and $\mathbf{D}$ is a positive matrix with eigenvalues bounded away from 00 and +$\infty$+ ∞. Under this model, the following proposition provides further insight into the role of $\mathbf{Q}$.
Proposition F.2.
Assume that Assumptions 2.2, 2.4, 2.5 and 2.6 are satisfied, and ${\bm{\Sigma}}=\mathbf{B}\mathbf{B}^{\top}+\mathbf{D}$ as specified. Let $K$ be fixed and suppose $\bm{b}_{j}^{\top}\mathbf{D}^{-1}{\bm{\mu}}=O(\|{\bm{\mu}}\|_{2})$. Then, for any given $\varepsilon>0$, there exists deterministic sequences of matrices $\widetilde{\mathbf{Q}}$ with bounded operator norm such that with probablity 1,
$1-\varepsilon\leq\lim_{n\to+\infty}SR(\widetilde{\mathbf{Q}})/SR_{\max}\leq 1$
The proof of Proposition F.2 is given in Section F.2. This proposition demonstrates that even with the presence of factors, i.e. spiked eigenvalues, it is still possible to identify a regularization matrix $\widetilde{\mathbf{Q}}$ with a bounded operator norm that can approximate the maximum population Sharpe ratio. Again, we can utilize Theorem 2.7 to evaluate the effectiveness of the design of $\mathbf{Q}$.
F.1 Proof of Proposition F.1
Recall the portfolio allocation vector $\mathbf{w}\propto(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\mu}}$, where $\widehat{\bm{\Sigma}}$ represents the sample covariance matrix and $\mathbf{Q}$ is a Ridge regularization term. The Sharpe ratio $SR(\mathbf{Q})$ can be written as:
$SR(\mathbf{Q})=\frac{{\bm{\mu}}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\mu}}}{\sqrt{{\bm{\mu}}^{\top}{\mathrm{tr}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\mu}}}}.$
From Proposition D.2, it can be seen that:
$\frac{1}{\sqrt{1-\frac{s_{1,{\bm{\Sigma}}}}{(1+s_{0})^{2}}}}\cdot\frac{{\bm{\mu}}^{\top}({\bm{\Sigma}}/(1+s_{0})+\mathbf{Q})^{-1}{\bm{\mu}}}{\sqrt{{\bm{\mu}}^{\top}({\bm{\Sigma}}/(1+s_{0})+\mathbf{Q})^{-1}{\bm{\Sigma}}({\bm{\Sigma}}/(1+s_{0})+\mathbf{Q})^{-1}{\bm{\mu}}}}\cdot\frac{1}{SR(\mathbf{Q})}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1,$
where $s_{0}$ and $s_{1,{\bm{\Sigma}}}$ are defined in Lemma C.5 by letting $z=0$. If $\mathbf{Q}$ is carefully designed such that $\mathbf{Q}=q\cdot{\bm{\Sigma}}$, we have
$\frac{1}{\sqrt{1-\frac{s_{1,{\bm{\Sigma}}}}{(1+s_{0})^{2}}}}\cdot\frac{{\bm{\mu}}^{\top}({\bm{\Sigma}}/(1+s_{0})+q{\bm{\Sigma}})^{-1}{\bm{\mu}}}{\sqrt{{\bm{\mu}}^{\top}({\bm{\Sigma}}/(1+s_{0})+q{\bm{\Sigma}})^{-1}{\bm{\Sigma}}({\bm{\Sigma}}/(1+s_{0})+q{\bm{\Sigma}})^{-1}{\bm{\mu}}}}\cdot\frac{1}{SR(\mathbf{Q})}$
$\quad=\frac{1}{\sqrt{1-\frac{s_{1,{\bm{\Sigma}}}}{(1+s_{0})^{2}}}}\cdot\sqrt{{\bm{\mu}}^{\top}{\bm{\Sigma}}^{-1}{\bm{\mu}}}/SR(\mathbf{Q})~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1.$
Here, $s_{0}=\frac{c}{p}{\mathrm{tr}}{\bm{\Sigma}}({\bm{\Sigma}}/(1+s_{0})+q{\bm{\Sigma}})^{-1}=\frac{c(1+s_{0})}{q(1+s_{0})+1}$. It should be noted that $s_{0}$ is uniformly bounded, and for any $\varepsilon>0$, there exists $M>0$ such that for any given $q>M$ it holds that $30<s_{0}<\varepsilon/$ 3. We have $30<-s_{1,{\bm{\Sigma}}}/(1+s_{0})^{2}\leq s_{0}\leq\varepsilon/$ 3, where the first inequality comes from Lemma C.5. By taking $\widetilde{\mathbf{Q}}=2M{\bm{\Sigma}}$, we conclude that with probability 1,
$1\geq\lim_{n\to+\infty}SR(\widetilde{\mathbf{Q}})/SR_{\max}\geq\lim_{n\to+\infty}\frac{1}{\sqrt{1-\frac{s_{1,{\bm{\Sigma}}}}{(1+s_{0})^{2}}}}\geq\lim_{n\to+\infty}\frac{1}{\sqrt{1+\frac{\varepsilon}{3}}}\geq 1-\varepsilon,$
which completes the proof of Proposition F.1.
F.2 Proof of Proposition F.2
From Proposition D.2, it can be seen that:
$\frac{1}{\sqrt{1-\frac{s_{1,{\bm{\Sigma}}}}{(1+s_{0})^{2}}}}\cdot\frac{{\bm{\mu}}^{\top}({\bm{\Sigma}}/(1+s_{0})+\mathbf{Q})^{-1}}{\sqrt{{\bm{\mu}}^{\top}({\bm{\Sigma}}/(1+s_{0})+\mathbf{Q})^{-1}{\bm{\Sigma}}({\bm{\Sigma}}/(1+s_{0})+\mathbf{Q})^{-1}{\bm{\mu}}}}\cdot\frac{1}{bmu/SR(\mathbf{Q})}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1,$
where $s_{0}$ and $s_{1,{\bm{\Sigma}}}$ are defined in Lemma C.5 by letting $z=0$. We first consider the expression of ${\bm{\mu}}^{\top}({\bm{\Sigma}}/(1+s_{0})+\mathbf{Q})^{-1}{\bm{\mu}}$ and ${\bm{\mu}}^{\top}({\bm{\Sigma}}/(1+s_{0})+\mathbf{Q})^{-1}{\bm{\Sigma}}({\bm{\Sigma}}/(1+s_{0})+\mathbf{Q})^{-1}{\bm{\mu}}$ by specifically choosing $\mathbf{Q}=q\cdot D$. Letting $\lambda=1+(1+s_{0})q$, we have
$\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+q\mathbf{D}\bigg{)}^{-1}/(1+s_{0})$ $=\big{(}\mathbf{B}\mathbf{B}^{\top}+\lambda\mathbf{D}\big{)}^{-1}$
$=\frac{1}{\lambda}\cdot\big{(}\mathbf{D}^{-1}-\mathbf{D}^{-1}\mathbf{B}(\lambda\mathbf{I}_{K}+\mathbf{B}^{\top}\mathbf{D}^{-1}\mathbf{B})^{-1}\mathbf{B}^{\top}\mathbf{D}^{-1}\big{)}$
$=\frac{1}{\lambda}\cdot\big{(}\mathbf{D}^{-1}-\mathbf{D}^{-1}\mathbf{B}{\bm{\Xi}}\mathbf{B}^{\top}\mathbf{D}^{-1}\big{)},$ (F.1)
where ${\bm{\Xi}}$ is defined as ${\bm{\Xi}}=(\lambda\mathbf{I}_{K}+\mathbf{B}^{\top}\mathbf{D}^{-1}\mathbf{B})^{-1}\in\mathbb{R}^{K\times K}$. Here, the second equality is by Woodbury formula. We can easily show that $\|{\bm{\Xi}}\|_{{\mathop{\mathrm{op}}}}=O(1/p)$. From (F.1), we have
$\frac{{\bm{\mu}}^{\top}({\bm{\Sigma}}/(1+s_{0})+\mathbf{Q})^{-1}{\bm{\mu}}}{\sqrt{{\bm{\mu}}^{\top}({\bm{\Sigma}}/(1+s_{0})+\mathbf{Q})^{-1}{\bm{\Sigma}}({\bm{\Sigma}}/(1+s_{0})+\mathbf{Q})^{-1}{\bm{\mu}}}}$
$\qquad=\frac{{\bm{\mu}}^{\top}\big{(}\mathbf{D}^{-1}-\mathbf{D}^{-1}\mathbf{B}{\bm{\Xi}}\mathbf{B}^{\top}\mathbf{D}^{-1}\big{)}{\bm{\mu}}}{\sqrt{{\bm{\mu}}^{\top}\big{(}\mathbf{D}^{-1}-\mathbf{D}^{-1}\mathbf{B}{\bm{\Xi}}\mathbf{B}^{\top}\mathbf{D}^{-1}\big{)}{\bm{\Sigma}}\big{(}\mathbf{D}^{-1}-\mathbf{D}^{-1}\mathbf{B}{\bm{\Xi}}\mathbf{B}^{\top}\mathbf{D}^{-1}\big{)}{\bm{\mu}}}}.$
By the condition that $|\bm{b}_{j}^{\top}\mathbf{D}^{-1}{\bm{\mu}}|=O(\|{\bm{\mu}}\|_{2})$, we can see
${\bm{\mu}}^{\top}\big{(}\mathbf{D}^{-1}-\mathbf{D}^{-1}\mathbf{B}{\bm{\Xi}}\mathbf{B}^{\top}\mathbf{D}^{-1}\big{)}{\bm{\mu}}={\bm{\mu}}^{\top}\mathbf{D}^{-1}{\bm{\mu}}+O(\|{\bm{\mu}}\|_{2}^{2}/p),$
${\bm{\mu}}^{\top}\big{(}\mathbf{D}^{-1}-\mathbf{D}^{-1}\mathbf{B}{\bm{\Xi}}\mathbf{B}^{\top}\mathbf{D}^{-1}\big{)}\mathbf{B}=\lambda{\bm{\mu}}^{\top}\mathbf{D}^{-1}\mathbf{B}{\bm{\Xi}}=O(\|{\bm{\mu}}\|_{2}^{2}/p),$
${\bm{\mu}}^{\top}\big{(}\mathbf{D}^{-1}-\mathbf{D}^{-1}\mathbf{B}{\bm{\Xi}}\mathbf{B}^{\top}\mathbf{D}^{-1}\big{)}\mathbf{D}\big{(}\mathbf{D}^{-1}-\mathbf{D}^{-1}\mathbf{B}{\bm{\Xi}}\mathbf{B}^{\top}\mathbf{D}^{-1}\big{)}{\bm{\mu}}$
$\qquad={\bm{\mu}}^{\top}\mathbf{D}^{-1}{\bm{\mu}}-2{\bm{\mu}}^{\top}\mathbf{D}^{-1}\mathbf{B}{\bm{\Xi}}\mathbf{B}\mathbf{D}^{-1}{\bm{\mu}}+{\bm{\mu}}^{\top}\mathbf{D}^{-1}\mathbf{B}{\bm{\Xi}}\mathbf{B}^{\top}\mathbf{D}^{-1}\mathbf{B}{\bm{\Xi}}\mathbf{B}\mathbf{D}^{-1}{\bm{\mu}}$
$\qquad={\bm{\mu}}^{\top}\mathbf{D}^{-1}{\bm{\mu}}+O(\|{\bm{\mu}}\|_{2}^{2}/p).$
Here, we use the fact $\mathbf{B}^{\top}\mathbf{D}^{-1}\mathbf{B}{\bm{\Xi}}=\mathbf{I}_{K}-\lambda{\bm{\Xi}}$ and $\|{\bm{\Xi}}\|_{{\mathop{\mathrm{op}}}}=O(1/p)$. Hence, we conclude that by letting $\mathbf{Q}=q\cdot\mathbf{D}$, it holds
$\frac{{\bm{\mu}}^{\top}({\bm{\Sigma}}/(1+s_{0})+\mathbf{Q})^{-1}{\bm{\mu}}}{\sqrt{{\bm{\mu}}^{\top}({\bm{\Sigma}}/(1+s_{0})+\mathbf{Q})^{-1}{\bm{\Sigma}}({\bm{\Sigma}}/(1+s_{0})+\mathbf{Q})^{-1}{\bm{\mu}}}}$ $=\frac{{\bm{\mu}}^{\top}\mathbf{D}^{-1}{\bm{\mu}}+O(\|{\bm{\mu}}\|_{2}^{2}/p)}{\sqrt{{\bm{\mu}}^{\top}\mathbf{D}^{-1}{\bm{\mu}}+O(\|{\bm{\mu}}\|_{2}^{2}/p)}}$
$=\sqrt{{\bm{\mu}}^{\top}\mathbf{D}^{-1}{\bm{\mu}}}+O(\|{\bm{\mu}}\|_{2}/p).$
Note also that
$SR_{\max}^{2}={\bm{\mu}}^{\top}{\bm{\Sigma}}^{-1}{\bm{\mu}}={\bm{\mu}}^{\top}\mathbf{D}^{-1}{\bm{\mu}}-{\bm{\mu}}^{\top}\mathbf{D}^{-1}\mathbf{B}(\mathbf{I}_{K}+\mathbf{B}^{\top}\mathbf{D}^{-1}\mathbf{B})^{-1}\mathbf{B}^{\top}{\bm{\mu}}={\bm{\mu}}^{\top}\mathbf{D}^{-1}{\bm{\mu}}+O(\|{\bm{\mu}}\|_{2}^{2}/p),$
we have
$\lim_{n\to+\infty}\frac{{\bm{\mu}}^{\top}({\bm{\Sigma}}/(1+s_{0})+\mathbf{Q})^{-1}{\bm{\mu}}}{\sqrt{{\bm{\mu}}^{\top}({\bm{\Sigma}}/(1+s_{0})+\mathbf{Q})^{-1}{\bm{\Sigma}}({\bm{\Sigma}}/(1+s_{0})+\mathbf{Q})^{-1}{\bm{\mu}}}}\cdot\frac{1}{SR_{\max}}=1.$
Therefore, by letting $\mathbf{Q}=q\cdot\mathbf{D}$, with probability 1,
$1\geq\lim_{n\to+\infty}SR(\mathbf{Q})/SR_{\max}=\lim_{n\to+\infty}\frac{1}{\sqrt{1-\frac{s_{1,{\bm{\Sigma}}}}{(1+s_{0})^{2}}}}.$
Here, $s_{0}$ and $s_{1,{\bm{\Sigma}}}$ are defined in Lemma C.5 by letting $z=0$ and $\mathbf{Q}=q\cdot\mathbf{D}$. The solution of $s_{0}$ is given by
$s_{0}=\frac{c}{p}{\mathrm{tr}}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+q\mathbf{D}\bigg{)}^{-1},$
rewrite the formula above we can see
$s_{0}=\frac{c(1+s_{0})}{p}{\mathrm{tr}}(\mathbf{B}\mathbf{B}^{\top}+\mathbf{D})\big{(}\mathbf{B}\mathbf{B}^{\top}+\lambda\mathbf{D}\big{)}^{-1}.$
Here, $\lambda=1+(1+s_{0})q$. Easy to see that $\|\mathbf{B}^{\top}\big{(}\mathbf{B}\mathbf{B}^{\top}+\lambda\mathbf{D}\big{)}^{-1}\mathbf{B}\|_{{\mathop{\mathrm{op}}}}=O(1)$, hence we have
$s_{0}$ $=\frac{c(1+s_{0})}{p}{\mathrm{tr}}\mathbf{D}\big{(}\mathbf{B}\mathbf{B}^{\top}+\lambda\mathbf{D}\big{)}^{-1}+O(1/p)$
$=\frac{c}{p}{\mathrm{tr}}\bigg{(}\frac{\mathbf{D}^{-\frac{1}{2}}\mathbf{B}\mathbf{B}^{\top}\mathbf{D}^{-\frac{1}{2}}+\mathbf{I}}{1+s_{0}}+q\mathbf{I}\bigg{)}^{-1}+O(1/p)$
$\leq c/q+O(1/p)$
Here, we use the fact that $s_{0}$ is uniformly bounded. For any $\varepsilon>0$, there exists constant $M>0$ such that for any $q>M$, $s_{0}\leq\varepsilon/6+O(1/p)$, hence by letting $\widetilde{\mathbf{Q}}=2M\mathbf{D}$, we have $\lim_{n\to+\infty}s_{0}\leq\varepsilon/6$. Thus when $\mathbf{Q}=\widetilde{\mathbf{Q}}=2M\mathbf{D}$, with probability 1 it holds that $60<\lim_{n\to+\infty}-s_{1,{\bm{\Sigma}}}/(1+s_{0})^{2}\leq\lim_{n\to+\infty}s_{0}\leq\varepsilon/$ 6, which indicates that
$\lim_{n\to+\infty}SR(\widetilde{\mathbf{Q}})/SR_{\max}=\lim_{n\to+\infty}\frac{1}{\sqrt{1-\frac{s_{1,{\bm{\Sigma}}}}{(1+s_{0})^{2}}}}\geq\frac{1}{\sqrt{1+\varepsilon/6}}\geq 1-\varepsilon$
with probability 1. This completes the proof.